### v4

# On the quantitative lower bounds for solutions to the Boltzmann equation in nonconvex domains

Abstract.
In this article, we study the continuous mild solutions to the Boltzmann equation in a bounded spatial domain, under either angular cutoff assumption or non-cutoff assumption. Without assuming convexity of the spatial domain, we establish a Maxwellian lower bound in the cutoff case, and a weaker-than-Maxwellian lower bound for the non-cutoff case. This extends the results of [1, 2], where the convexity of the domain was required.
1. Introduction
In this paper, we investigate the lower bounds of the mild solutions to the Boltzmann equation on a spatial domain $\Omega\subset\mathbb{R}^{3}$ satisfying
Assumption A.
$\Omega$ is a connected, bounded open set with $C^{2}$ boundary.
The Boltzmann equation under consideration reads
(1.1) $\partial_{t}f(t,x,v)+v\cdot\nabla f(t,x,v)=Q[f,f](t,x,v),\ \forall(t,x,v)\in[0,T)\times\Omega\times\mathbb{R}^{3},$
(1.2) $f(0,x,v)=f_{0}(x,v),\ \forall(x,v)\in\Omega\times\mathbb{R}^{3},$
where $T>0$. We consider the mixed boundary condition: for any $(x,v)\in\Gamma^{-}$, we have
(1.3) $f(x,v)=\alpha f(t,x,R(x,v))+(1-\alpha)\left(\int_{w\cdot n(x)>0}f(t,x,w)(w\cdot n(x))dw\right)\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}$
for some $T_{B}>0$ and $\alpha\in[0,1]$. Here, $\Gamma^{-}:=\{(x,v)\in\partial\Omega\times\mathbb{R}^{3}|n(x)\cdot v<0\}$, $R(x,v):=v-2(v\cdot n(x))n(x)$, where $n(x)$ is the outward unit normal vector at the boundary point $x\in\partial\Omega$.
The collision operator $Q$ is defined as
(1.4) $Q[h_{1},h_{2}](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})-h_{2}(v)h_{1}(v_{*})]\,dv_{*}\,d\sigma,$
where $^{\prime}, vv_{*}$, $v^{\prime}_{*}$ are defined as:
$v^{\prime}:=\frac{v+v_{*}}{2}+\frac{|v-v_{*}|}{2}\sigma,\ v^{\prime}_{*}:=\frac{v+v_{*}}{2}-\frac{|v-v_{*}|}{2}\sigma$
with $\cos{\theta}:=\left\langle\frac{v-v_{*}}{|v-v_{*}|},\sigma\right\rangle$. We assume that the collision kernel $B\geq 0$ and satisfies:
Assumption B.
(1.5) $B(|v-v_{*}|,\cos{\theta})=\Phi(|v-v_{*}|)b(\cos{\theta}),$
where $\Phi:=\Phi(|v-v_{*}|)$ is a function defined on $0,\infty)$ satisfying
(1.6) $c_{\Phi}r^{\gamma}\leq\Phi(r)\leq C_{\Phi}r^{\gamma},\ \forall r\in[0,\infty),$
or
(1.7) $\begin{cases}c_{\Phi}r^{\gamma}\leq\Phi(r)\leq C_{\Phi}r^{\gamma},\ &\forall r\in[1,\infty),\\ c_{\Phi}\leq\Phi(r)\leq C_{\Phi},\ &\forall r\in[0,1]\end{cases}$
for some positive constants $c_{\Phi}$, $C_{\Phi}$ and $\gamma\in(-3,1]$. Here, we assume that the function $b:\theta\mapsto b(\cos{\theta})$ is continuous on $(0,\pi]$, positive near $\frac{\pi}{2}$, and satisfies
(1.8) $\lim\limits_{\theta\rightarrow 0^{+}}\frac{b(\cos{\theta})\sin{\theta}}{\theta^{-(1+\nu)}}=b_{0},$
for some $b_{0}>0$ and $\nu\in(-\infty,2)$.
The problems of quantifying the positivity of the solutions to (1.1)-(1.3) have attracted considerable attention from many authors. Beyond interest in physics, the problem also plays a crucial role in kinetic theory. For example, the exponential lower bound is crucial in the study of the behavior of the solution to the related Laudau equation in [4, 5]. In [6], the Maxwellian lower bound is applied to derive the uniqueness of the solution to the Boltzmann equation in $L_{v}^{1}L_{x}^{\infty}(1+|v|^{2+0})$. In 1933, Carleman [3] showed the existence of exponential lower bounds on the radially symmetric solution of the spatially homogeneous Boltzmann equation with angular cutoff hard potentials. Since then, many results have been derived. For example, in [12] Pulvirenti and Wennberg improved the result by proving the Maxwellian lower bound for non-radially symmetric solutions of the spatially homogeneous Boltzmann equation with angular cutoff hard potentials. In [11], Mouhot removed the homogeneous assumption and obtained the Maxwellian lower bound for the torus domain. This result was extended to the bounded convex domain with $C^{2}$ boundary by Briant in [1, 2]. We note that in the articles [1, 2, 11], they also derived the ”less than Maxwellian” lower bounds for the non-cutoff case. Later, this result was improved to Maxwellian lower bounds in the case of hard and moderately soft potentials with the assumption that $\Omega$ is a torus by Imbert, Mouhot, Silvestre in [9] and in the case where $\Omega$ is $\mathbb{R}^{3}$ by Henderson, Snelson, and Tarfulea in [8].
However, the problem of Maxwellian lower bounds on a non-convex spatial domain remains an open problem in both the cutoff and non-cutoff cases. In this article, we derive a Maxwellian lower bound on an open bounded connected but not necessarily convex domain with a $C^{2}$ boundary for the cutoff case. For the non-cutoff case, we also derive a weaker than Maxwellian lower bound.
In [1, 2, 11], the following spreading property of the collision operator
(1.9) $Q^{+}\left[\mathbf{1}_{B(v_{0},\delta)},\mathbf{1}_{B(v_{0},\delta)}\right](v)\geq C\xi^{\frac{1}{2}}\mathbf{1}_{B(v_{0},\delta\sqrt{2}(1-\xi))}(v),$
for some $C>0$ depending on the cross section, and any $\xi\in(0,1)$, is a key ingredient to deriving a lower bound around one point, which could then be extended to any other point using the convexity of the domain.
In this article, we focus on the geometry of a $C^{2}$ connected domain and show that for any two points of $\Omega$ there always exists a suitable zigzag between these two points, which allows the spreading of the lower bound from one point to the other point like in [1, 2]. One major difficulty arises from the requirement that these zigzags in the domain must remain sufficiently distant from the boundary of $\Omega$. In fact, the shape structure affects the amount of gas molecules passing through the boundary. In [1, 2], the lower bounds depend on: (1) hydrodynamic quantities. (2) the modulus of continuity of $f_{0}$. In our result, the lower bounds also depend on $\operatorname{Conn}_{d}(\Omega)$, the maximum number of ”good” zigzag segments needed to connect any two points in the spatial domain. We put the definition of $\operatorname{Conn}_{d}(\Omega)$ in (2.37) and address the detailed analysis concerning the geometry in Section 2. Our main contribution lies in showing that the Maxwellian lower bound can be propagated throughout a non-convex domain by exploiting geometric connectivity properties, quantified through $\operatorname{Conn}_{d}(\Omega)$.
Throughout Sections 1–5, we assume that $\nu<0$ (cutoff case), where $\nu$ is defined in (1.8). To classify the type of solutions to (1.1)–(1.3), we first decompose the collision operator $Q$ as follows:
(1.10) $\begin{split}&Q[h_{1},h_{2}](v)\\ =&\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})-h_{2}(v)h_{1}(v_{*})]\,dv_{*}\,d\sigma\\ =&\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})]\,dv_{*}\,d\sigma\\ &-h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})h_{1}(v_{*})\,dv_{*}\,d\sigma\\ =&:Q^{+}[h_{1},h_{2}](v)-h_{2}(v)L[h_{1}](v),\end{split}$
where
$Q^{+}[h_{1},h_{2}](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})]\,dv_{*}\,d\sigma,$
$L[h_{1}](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})h_{1}(v_{*})\,dv_{*}\,d\sigma.$
Next, we define the following notations:
$t_{b}(x,v):=\sup\left\{\{0\}\cup\{t>0\mid x-sv\in\Omega,\,\forall\,0<s<t\}\right\},$
$x_{b}(x,v):=x-t_{b}(x,v)v,$
$\Gamma^{0}:=\{(x,v)\in\partial\Omega\times\mathbb{R}^{3}\mid n(x)\cdot v=0\},$
$\Gamma^{+}:=\{(x,v)\in\partial\Omega\times\mathbb{R}^{3}\mid n(x)\cdot v>0\},$
$\Gamma^{-}:=\{(x,v)\in\partial\Omega\times\mathbb{R}^{3}\mid n(x)\cdot v<0\},$
$\Gamma_{-}^{0}:=\{(x,v)\in\Gamma^{0}\mid t_{b}(x,v)=0,t_{b}(x,-v)>0,\,\exists\,\delta>0\ni\,x-s^{\prime}v\in(\overline{\Omega})^{c}\,\forall\,s^{\prime}\in(0,\delta)\},$
(1.11) $\begin{split}&\Gamma_{conti}:=\left\{\{0\}\times\overline{\Omega}\times\mathbb{R}^{3}\right\}\cup\left\{(0,\infty)\times(\Gamma^{-}\cup\Gamma^{0}_{-})\right\}\\ &\cup\left\{(t,x,v)\in(0,T)\times\{\Omega\times\mathbb{R}^{3}\cup\Gamma^{+}\}\ |t<t_{b}(x,v)\,\mathrm{or}\,(x_{b}(x,v),v)\in\Gamma^{-}\cup\Gamma^{0}_{-})\right\}.\end{split}$
Using the above notations and decomposition, we are ready to introduce the continuous mild solutions of the Boltzmann equation:
Definition 1.1.
We assume that the domain $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and the collision kernel $B$ satisfies Assumption B with $\nu<0$, where $\nu$ is introduced in (1.8). Given a nonnegative, continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$, we call a nonnegative function $f$ defined on $0,T)\times(\overline{\Omega}\times\mathbb{R}^{3})$ with $|f(t,x,v)|\leq C(1+|v|)^{-r}$ for some constant $C>0$ and $r>3$ for any $0<t\leq T$, $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$ a ”continuous mild” solution to (1.1)–(1.3) with initial data $f_{0}$ if $f$ is continuous on $\Gamma_{conti}$ and for $(t,x,v)\in[0,T)\times\Omega\times\mathbb{R}^{3}$ and the function $f(t,x,v)$ satisfies the following integral identities:
(1.12) $\begin{split}f(t,x,v)&=f_{0}(X_{0,t}(x,v),v)\exp\left(-\int_{0}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right)\\ &+\int_{0}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &\hskip 28.45274ptQ^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\end{split}$
when $t\leq t_{\partial}(x,v):=\sup\{t\geq 0|x-vs\in\overline{\Omega},\ \forall s\in[0,t]\}$, and
(1.13) $\begin{split}f(t,x,v)&=\alpha f(t-t_{\partial}(x,v),X_{t-t_{\partial}(x,v),t}(x,v),R(X_{t-t_{\partial}(x,v),t}(x,v),v))\\ &\hskip 14.22636pt\exp\left(-\int_{t-t_{\partial}(x,v)}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right)\\ &+(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v))>0}f(t,X_{t-t_{\partial}(x,v),t}(x,v),w)(w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v)))dw\right)\\ &\hskip 51.21504pt\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}\exp\left(-\int_{t-t_{\partial}(x,v)}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right)\\ &+\int_{t-t_{\partial}(x,v)}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &\hskip 39.83368ptQ^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\end{split}$
when $t\geq t_{\partial}(x,v)$. Here, a detailed definition of the characteristic line $X_{s,t}(x,v)$ is provided in the Appendix.
Remark 1.1.
In the case of specular reflection ($\alpha=1$), the continuous mild solution of (1.1)–(1.3) can also be written as the continuous function on $\Gamma_{conti}$, $f$ such that for all $(t,x,v)\in[0,T]\times\overline{\Omega}\times\mathbb{R}^{3}$,
(1.14) $\begin{split}f(t,x,v)&=f_{0}(X_{0,t}(x,v),V_{0,t}(x,v))\exp\left(-\int_{0}^{t}L[f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds\right)\\ &+\int_{0}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](V_{s^{\prime},t}(x,v))\,ds^{\prime}\right)\\ &\hskip 28.45274ptQ^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds,\end{split}$
where $V_{s,t}$ is defined in the Appendix.
Remark 1.2.
The continuity assumption of the solution $f(t,x,v)$ is technical, as the proof only involves the continuity of the initial condition. Nonetheless, in the diffusive reflection boundary condition ($\alpha=0$), the existence of a solution of bounded variation is established in [7]. However, as shown in [10], unlike in the convex case, the solution loses continuity, with singularities propagating along the grazing set. Motivated by [10], we therefore consider the set $\Gamma_{conti}$. In the same work, the authors also showed that the continuity of solutions on $\Gamma_{conti}$ is ensured under diffuse reflection boundary conditions, with continuity of the initial data and addition conditions.
Before stating our main result, we introduce the following hydrodynamic quantities:
(1.15) $\varrho_{f}(t.x):=\int_{v\in\mathbb{R}^{3}}f^{2}(t,x,v)\,dv$
(1.16) $e_{f}(t,x):=\int_{v\in\mathbb{R}^{3}}|v|^{2}f(t,x,v)\,dv,\,E_{f}:=\sup\limits_{[0,T)\times\Omega}(e_{f}(t,x)+\mathcal{\varrho}_{f}(t.x)),$
(1.17) $l_{f,p}(t,x):=\left(\int_{v\in\mathbb{R}^{3}}f^{p}(t,x,v)\,dv\right)^{\frac{1}{p}},\,L_{f,p}:=\sup\limits_{[0,T)\times\Omega}l_{f,p}(t,x),$
It is also known that the quantity (total mass)
(1.18) $M:=\int_{\Omega\times\mathbb{R}^{3}}f(t,x,v)\,dx\,dv$
does not depend on $t$, where $f$ is a (continuous mild) solution of (1.1)–(1.3).
We can now state our main result as follows:
Theorem 1.1.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the kernel $B$ satisfies Assumption B with $\nu<0$. We consider a non-negative continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) on $0,T)\times\overline{\Omega}\times\mathbb{R}^{3}$, with initial condition $f_{0}$ for some $T>0$, and let $\alpha\in[0,1]$. We assume that $f(t,x,v)$ satisfies the following properties:
  1. (1)
    $M>0$;
  2. (2)
    $E_{f}<\infty$ if $\gamma\geq 0$ and $\max\{E_{f},L_{f,{p_{\gamma}}}\}<\infty$, where $p_{\gamma}>\frac{3}{3+\gamma}>0$, if $3<\gamma<0$.
Then, the following lower bound holds: There exists $0<\tau_{0}\leq T$ such that for any $\tau\in(0,\tau_{0})$, there exist $\Delta_{\tau_{0}}>0$, $\rho>0$ and $\theta>0$ depending on $\tau_{0}$, $M$, $C_{\Phi},c_{\Phi},\gamma,b_{0},\nu,E_{f}$ (and $L_{f,p}$ if $\gamma<0$ ), $\tau$, $\alpha$, $\Omega$ and on the modulus of continuity of $f_{0}$, such that
(1.19) $f(t,x,v)\geq\frac{\rho}{(2\pi\theta)^{\frac{3}{2}}}e^{-\frac{|v|^{2}}{2\theta}},\ \forall\,t\in[\tau,\Delta_{\tau_{0}}),\ \forall\,(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}.$
The next theorem can be derived from Theorem 1.1:
Theorem 1.2.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the kernel $B$ satisfies Assumption B with $\nu<0$. We consider a non-negative continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) on $0,T)\times\overline{\Omega}\times\mathbb{R}^{3}$, with initial condition $f_{0}$ for some $T>0$, and let $\alpha\in[0,1]$. We assume that $f(t,x,v)$ satisfies the following properties:
  1. (1)
    $M>0$;
  2. (2)
    $E_{f}<\infty$ if $\gamma\geq 0$ and $\max\{E_{f},L_{f,{p_{\gamma}}}\}<\infty$, where $p_{\gamma}>\frac{3}{3+\gamma}>0$, if $3<\gamma<0$.
Then, the following lower bound holds: For any $\tau\in(0,T)$, there exist $\rho>0$ and $\theta>0$ depending on $M$, $C_{\Phi},c_{\Phi},\gamma,b_{0},\nu,E_{f}$ (and $L_{f,p}$ if $\gamma<0$ ), $\tau$, $\alpha$, $\Omega$ and on the modulus of continuity of $f_{0}$, such that
(1.20) $f(t,x,v)\geq\frac{\rho}{(2\pi\theta)^{\frac{3}{2}}}e^{-\frac{|v|^{2}}{2\theta}},\ \forall\,t\in[\tau,T),\ \forall\,(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}.$
For the non-cutoff case, the result is stated in Chapter 6. We note that the numbers $\rho$ and $\theta$ are computable thanks to the special zigzag which will be introduced in Proposition 2.5, through which we can guarantee that a certain amount and range of the lower bound can be generated through the propagation in Proposition 3.4 and derive a computable uniform lower bound (which depends on $\Omega$ since the construction of the zigzag is affected by the geometric properties of $\Omega$) as in Proposition 3.5.
The remainder of the article is organized as follows. In Section 2, we analyze the boundary of non-convex domain to derive some key properties. In Section 3, we focus on the derivation of a series of diluting initial lower bounds and the propagation effect of lower bounds. In Section 4 and Section 5, we provide a detailed proof of the Maxwellian lower bound on mild solution to the Boltzmann equation with non-fully specular ($0\leq\alpha<1$) and fully specular reflection boundary condition ($\alpha=1$), respectively. In Section 6, a weaker exponential lower bound of solutions to the Boltzmann equation for non-cutoff case is provided.
2. Geometric properties near the boundary
In this section, we introduce some useful lemmas.
Lemma 2.1.
Given $\Omega$ which satisfies Assumption A, there exists $\delta:=\delta(\Omega)>0$, such that for any $0<d<\min\{1,\delta\}$, one can find $m_{1}\in\mathbb{N}$, $\{x^{0}_{i}\}_{i=1}^{m_{1}}\in\partial\Omega$ depending on $d$ such that
(2.1) $\partial\Omega\subset\underset{1\leq i\leq m_{1}}{\bigcup}B\left(x_{i}^{0},\frac{d}{8}\right)$
so that, for each $1\leq i\leq m_{1}$, after orthogonal transformation of variables, the boundary near $x_{i}^{0}$ can be represented as a graph of a $C^{2}$ scalar function on $\mathbb{R}^{2}$. More precisely, for each $1\leq i\leq m_{1}$, there exists a orthonormal basis $\{e_{i}^{1},e_{i}^{2},-n(x_{i}^{0})\}$ and a $C^{2}$ function $\phi_{i}:\mathbb{R}^{2}\rightarrow\mathbb{R}$ such that
$\phi_{i}(0,0)=0,\nabla\phi_{i}(0,0)=0,$
(2.2) $\begin{split}&\partial\Omega\cap B(x_{i}^{0},3d)=\{y\in B(x_{i}^{0},3d)\mid y=x_{i}^{0}+u_{1}e_{i}^{1}+u_{2}e_{i}^{2}-\phi_{i}(u_{1},u_{2})n(x^{0}_{i})\},\\ &\partial\Omega\cap\overline{B(x_{i}^{0},3d)}=\{y\in\overline{B(x_{i}^{0},3d)}\mid y=x_{i}^{0}+u_{1}e_{i}^{1}+u_{2}e_{i}^{2}-\phi_{i}(u_{1},u_{2})n(x^{0}_{i})\},\\ &\Omega\cap B(x_{i}^{0},3d)=\{y\in B(x_{i}^{0},3d)\mid\\ &\hskip 99.58464pty=x_{i}^{0}+u_{1}e_{i}^{1}+u_{2}e_{i}^{2}-u_{3}n(x^{0}_{i}),u_{3}>\phi_{i}(u_{1},u_{2})\}.\end{split}$
Furthermore, we have $|\nabla\phi_{i}|<\frac{1}{100}$, $|\nabla^{2}\phi_{i}|<\widetilde{C}(\Omega)$ for some $\widetilde{C}(\Omega)>\max\left\{\frac{4}{d},\frac{1}{2d_{r}}\right\}$, and for any $1\leq i\leq m_{1}$,
(2.3) $B\left(x_{i}^{0}-\frac{d}{2}n(x_{i}^{0}),\frac{d}{2}\right)\subset B\left(x_{i}^{0},d\right)\cap\Omega.$
Here, $d_{r}$ is introduced in Remark 2.1.
Proof.
This is a direct consequence of the application of the Heine-Borel theorem. ∎
Remark 2.1.
It is known that any bounded domain with $C^{2}$ boundary satisfies the uniform interior sphere condition. That is, there exists $d_{r}$ such that for any $x\in\partial\Omega$, one can find a ball with radius $d_{r}>0$ in $\Omega$ that intersects with $\partial\Omega$ only on $x$. In this article, we choose $d$ from Lemma 2.1 to be smaller than $d_{r}$.
Corollary 2.1.
Let $\Omega$ satisfy Assumption A with $\delta:=\delta(\Omega)>0$. For any $0<d<\min\{1,\delta\}$, which is given in Lemma 2.1, given $1\leq i\leq m_{1}$, if $x\in B(x^{0}_{i},\frac{3d}{2})$ with coordinate representation
$x=x_{i}^{0}+x_{1}e^{1}_{i}+x_{2}e^{2}_{i}-x_{3}n(x_{i}^{0}),$
then
(2.4) $d(x,\partial\Omega\cap B(x^{0}_{i},3d))\geq\frac{1}{2}(x_{3}-\phi_{i}(x_{1},x_{2})).$
Proof.
This is a direct consequence of the fact that $|\nabla\phi_{i}|<\frac{1}{100}$ and the mean value theorem. Indeed, given $x=x_{i}^{0}+x_{1}e^{1}_{i}+x_{2}e^{2}_{i}-x_{3}n(x_{i}^{0})$, if there exists $x^{\prime}=x_{i}^{0}+x^{\prime}_{1}e^{1}_{i}+x^{\prime}_{2}e^{2}_{i}-x^{\prime}_{3}n(x_{i}^{0})\in B(x,\frac{1}{2}(x_{3}-\phi_{i}(x_{1},x_{2})))$ such that $x^{\prime}\in\partial\Omega\cap B(x^{0}_{i},3d))$, we have by Lemma 2.1 that
$\phi_{i}(x^{\prime}_{1},x^{\prime}_{2})=x^{\prime}_{3}.$
By the mean value theorem, there exists $\widetilde{x}_{1}$ between $x_{1}$ and $x^{\prime}_{1}$, $\widetilde{x}_{2}$ between $x_{2}$ and $x^{\prime}_{2}$ such that
(2.5) $|\nabla\phi_{i}(\widetilde{x}_{1},\widetilde{x}_{2})|=\frac{|\phi_{i}(x_{1},x_{2})-\phi_{i}(x^{\prime}_{1},x^{\prime}_{2})|}{|(x_{1},x_{2})-(x^{\prime}_{1},x^{\prime}_{2})|}\geq\frac{\frac{1}{2}(x_{3}-\phi_{i}(x_{1},x_{2}))}{\frac{1}{2}(x_{3}-\phi_{i}(x_{1},x_{2}))}=1,$
from which we deduce a contradiction to the fact that $|\nabla\phi_{i}(\widetilde{x}_{1},\widetilde{x}_{2})|<\frac{1}{100}$. ∎
Remark 2.2.
Notice that by the triangle inequality we have $\Omega_{\frac{d}{8}}\subset\bigcup\limits_{i=1}^{m_{1}}B(x_{i}^{0},\frac{d}{4})$, where $\Omega_{\epsilon}:=\{x\in\Omega|\ d(x,\partial\Omega)<\epsilon\}$.
Next, we define $y_{i}^{0}:=x_{i}^{0}-\frac{3d}{4}n(x_{i}^{0})$. Observe that $y_{i}^{0}\in\Omega-\Omega_{\frac{d}{8}}$, since $B(y_{i}^{0},\frac{d}{8})\subset B\left(x_{i}^{0}-\frac{d}{2}n(x_{i}^{0}),\frac{d}{2}\right)\subset B\left(x_{i}^{0},d\right)\cap\Omega$.
Lemma 2.2.
Given $\Omega$ which satisfies Assumption A. Let $0<d<\min(1,\delta)$ and $y_{i}^{0}\in\Omega$ as defined in Lemma 2.1.
Then, there exist $m_{2}\in\mathbb{N}$, $\{y_{i}^{0}\}_{i=m_{1}+1}^{m_{1}+m_{2}}\subset\Omega-\Omega_{\frac{d}{8}}$ such that
  1. (1)
    For any $x\in\Omega_{\frac{d}{8}}$, there exists $1\leq i\leq m_{1}$ such that $\overline{xy_{i}^{0}}\subset\Omega$.
  2. (2)
    For any $x\in\Omega-\Omega_{\frac{d}{8}}$, there exists $m_{1}+1\leq i\leq m_{1}+m_{2}$ with the property that $x\in B(y_{i}^{0},\frac{d}{16})$ and $\overline{xy_{i}^{0}}\subset\Omega$.
Here, $\overline{xy}:=\{sx+(1-s)y\mid 0\leq s\leq 1\}$.
Proof.
By Lemma 2.1 and Remark 2.2, there exist $m_{1}$ and $\{x_{i}^{0}\}_{i=1}^{m_{1}}$ such that
(2.6) $\Omega_{\frac{d}{8}}\subset\bigcup_{i=1}^{m_{1}}B\left(x_{i}^{0},\frac{d}{4}\right).$
Because of the representation of $\partial\Omega\cap B(x^{0}_{i},d)$ in Lemma 2.1, we see that
(2.7) $\overline{xy_{i}^{0}}\in\Omega,\ \forall x\in B\left(y_{i}^{0},\frac{d}{4}\right).$
Next, we cover $\Omega-\Omega_{\frac{d}{8}}$ by $\{B(x,\frac{d}{16})\}_{x\in\Omega-\Omega_{\frac{d}{8}}}$. Since $\Omega-\Omega_{\frac{d}{8}}$ is compact. The Heine-Borel covering theorem ensures the existence of a finite sub-cover of this collection (Here, we choose the sub-cover with the minimum number of covers). We then denote the corresponding centers by $\{y_{i}^{0}\}_{i=m_{1}+1}^{m_{1}+m_{2}}$. Finally, we conclude the proof by observing that
$B\left(y_{i}^{0},\frac{d}{16}\right)\subset\Omega-\Omega_{\frac{d}{16}}\subset\Omega,\ \forall\,m_{1}+1\leq i\leq m_{1}+m_{2}.$
Lemma 2.3.
Suppose $\Omega$ which satisfies Assumption A, with $\delta$ and $d$ as given in Lemma 2.1. For any $1\leq i\leq m_{1}$, given
$x=x^{0}_{i}+x_{1}e^{1}_{i}+x_{2}e^{2}_{i}-x_{3}n(x^{0}_{i}),$
$y=x^{0}_{i}+y_{1}e^{1}_{i}+y_{2}e^{2}_{i}-y_{3}n(x^{0}_{i})$
with $x\in\overline{B(x^{0}_{i},\frac{d}{4})},\,y\in\overline{B(x^{0}_{i},\frac{3d}{4})}$, $(y_{1}^{2}+y_{2}^{2})^{\frac{1}{2}}\leq\frac{d}{4}$, and $\frac{x-y}{|x-y|}\cdot n(x^{0}_{i})\geq\frac{1}{2}$,
we have
(2.8) $d(\overline{xy},\partial\Omega\cap\overline{B(x^{0}_{i},d)})=d(x,\partial\Omega\cap\overline{B(x^{0}_{i},d)}).$
Proof.
Note that it suffices to show that for any $0\leq q_{1}<1$, there exists $\epsilon^{\prime}(q_{1})>0$ such that for any $q_{1}<q_{2}<q_{1}+\epsilon^{\prime}(q_{1})$, we have $d(q_{1}x+(1-q_{1})y,\partial\Omega\cap\overline{B(x^{0}_{i},d)})>d(q_{2}x+(1-q_{2})y,\partial\Omega\cap\overline{B(x^{0}_{i},d)})$.
Figure 2.1. A figure illustrating how the points in the proof of Lemma 2.3 is located.
To show that, we first notice that by non-emptiness and compactness of $\partial\Omega\cap\overline{B(x^{0}_{i},d)}$, there exists a point $z^{\prime\prime}\in\partial\Omega\cap\overline{B(x^{0}_{i},d)}$ such that
$|z^{\prime\prime}-q_{1}x-(1-q_{1})y|=d(q_{1}x+(1-q_{1})y,\partial\Omega\cap\overline{B(x^{0}_{i},d)}).$
We parameterize $x$, $y$, $z^{\prime\prime}$ as follows:
$x=x^{0}_{i}+x_{1}e^{1}_{i}+x_{2}e^{2}_{i}-x_{3}n(x^{0}_{i}),$
$y=x^{0}_{i}+y_{1}e^{1}_{i}+y_{2}e^{2}_{i}-y_{3}n(x^{0}_{i}),$
$z^{\prime\prime}=x^{0}_{i}+z^{\prime\prime}_{1}e^{1}_{i}+z^{\prime\prime}_{2}e^{2}_{i}-z^{\prime\prime}_{3}n(x^{0}_{i}).$
We define the following quantities:
(2.9) $Z_{1}:=\frac{(z^{\prime\prime}-q_{1}x-(1-q_{1})y)}{|z^{\prime\prime}-q_{1}x-(1-q_{1})y|}=\frac{z^{\prime\prime}-z_{q_{1}}}{|z^{\prime\prime}-z_{q_{1}}|},$
(2.10) $Z_{2}:=\frac{x-q_{1}x-(1-q_{1})y}{|x-q_{1}x-(1-q_{1})y|}=\frac{x-y}{|x-y|}.$
Before we proceed, we also define $z_{q_{1}}:=q_{1}x+(1-q_{1})y$, and we notice that
$Z_{1}=n(z^{\prime\prime}).$
By Lemma 2.1, we have (we recall $Z_{1}$ from (2.9))
$Z_{1}\cdot n(x^{0}_{i})$
$=$ $n(z^{\prime\prime})\cdot n(x^{0}_{i})$
$=$ $\frac{1}{\sqrt{|\partial_{1}\phi_{i}(z^{\prime\prime}_{1},z^{\prime\prime}_{2})|^{2}+|\partial_{2}\phi_{i}(z^{\prime\prime}_{1},z^{\prime\prime}_{2})|^{2}+1}}$
$>$ $\sqrt{\frac{10000}{10001}}.$
Notice that we have from the assumption of Lemma 2.3 that (we recall $Z_{2}$ from (2.10))
$n(x^{0}_{i})\cdot Z_{2}\geq\frac{1}{2}.$
Hence, we deduce that
(2.11) $\begin{split}&Z_{1}\cdot Z_{2}\\ =&n(x^{0}_{i})\cdot Z_{2}+(Z_{1}-n(x^{0}_{i}))\cdot Z_{2}\\ \geq&\frac{1}{2}-|Z_{1}-n(x^{0}_{i})||Z_{2}|\\ =&\frac{1}{2}-\sqrt{|Z_{1}|^{2}+|n(x^{0}_{i})|^{2}-2Z_{1}\cdot n(x^{0}_{i})}\\ >&\frac{1}{2}-\sqrt{2\left(1-\sqrt{\frac{10000}{10001}}\right)}\\ >&\frac{\sqrt{2}}{3}.\end{split}$
Now, we define
$W_{1,//}:=((z^{\prime\prime}-z_{q_{1}})\cdot Z_{2})Z_{2}=|z^{\prime\prime}-z_{q_{1}}|(Z_{1}\cdot Z_{2})Z_{2},$
$W_{1,\perp}:=(z^{\prime\prime}-z_{q_{1}})-W_{1,//}.$
Clearly, we have $z^{\prime\prime}-z_{q_{1}}=W_{1,//}+W_{1,\perp}$, $W_{1,//}\mathrel{\!/\mkern-5.0mu/\!}Z_{2}\mathrel{\!/\mkern-5.0mu/\!}x-y$, and $W_{1,\perp}\perp Z_{2}$.
Finally, by taking $\epsilon^{\prime}(q_{1}):=\min\left\{\frac{\sqrt{2}|z^{\prime\prime}-z_{q_{1}}|}{6|x-y|},1-q_{1}\right\}$, we have
$z^{\prime\prime}-q_{2}x-(1-q_{2})y\in\overline{xy},$
and
$0$ $<|x-y|(q_{2}-q_{1})<|z^{\prime\prime}-q_{1}x-(1-q_{1})y|\frac{\sqrt{2}}{3}$
$<|z^{\prime\prime}-z_{q_{1}}|Z_{1}\cdot Z_{2}$
$=|W_{1,//}|,$
for any $q_{1}<q_{2}<q_{1}+\epsilon^{\prime}(q_{1})$.
Hence, we have
$|W_{1,//}-|x-y|(q_{2}-q_{1})Z_{2}|<|W_{1,//}|,$
and we deduce that
$d\bigg(q_{2}x+(1-q_{2})y,\partial\Omega\cap\overline{B(x^{0}_{i},d)}\bigg)<d\left(q_{1}x+(1-q_{1})y,\partial\Omega\cap\overline{B(x^{0}_{i},d)}\right)$
for any $q_{1}<q_{2}<q_{1}+\epsilon^{\prime}(q_{1})$. ∎
Now, we introduce one more property regarding the geometry of $\Omega$:
Lemma 2.4.
Suppose that $\Omega$ satisfies Assumption A with $\delta$, $d$ as given in Lemma 2.1. Then, $\Omega-\Omega_{\frac{d}{8}}$ is pathwise connected; that is, given $x^{\prime}_{1},x^{\prime}_{2}\in\Omega-\Omega_{\frac{d}{8}}$, there exists a continuous curve in $\Omega-\Omega_{\frac{d}{8}}$ joining $x^{\prime}_{1}$ and $x^{\prime}_{2}$.
Proof.
Given $x^{\prime}_{1},x^{\prime}_{2}\in\Omega-\Omega_{\frac{d}{8}}$. Since the domain $\Omega$ is connected, $\Omega$ is pathwise connected. Hence, there exist $l\in\mathbb{N},\{\widetilde{x}_{i}\}_{i=1}^{l}\in\Omega$ such that the following curve
$\phi(s)=\begin{cases}[1-(l+1)s]x^{\prime}_{1}+(l+1)s\widetilde{x}_{1},\ &\mathrm{for}\ 0\leq s\leq\frac{1}{l+1},\\ [1-(l+1)s+j]\widetilde{x}_{j}+[(l+1)s-j]\widetilde{x}_{j+1},\ &\mathrm{for}\ \frac{j}{l+1}\leq s\leq\frac{j+1}{l+1},\,1\leq j\leq l-1,\\ [1-(l+1)s+l]\widetilde{x}_{l}+[(l+1)s-l]x^{\prime}_{2},\ &\mathrm{for}\ \frac{l}{l+1}\leq s\leq 1,\end{cases}$
lies within $\Omega$.
As Remark 2.2 shows, we have
$\{\phi(s)\mid 0\leq s\leq 1\}\cap\Omega_{\frac{d}{8}}\subset\bigcup_{i=1}^{m_{1}}B(x^{0}_{i},\frac{d}{4}).$
Without loss of generality, we can assume that there is a function $k:\mathbb{N}\cap[1,l]\longrightarrow\mathbb{N}\cap[1,m_{1}]$ such that k(l1)=k(l)k(l-1)=k(l),
$\left\{\begin{aligned} \widetilde{x}_{1}&\in\partial B(x^{0}_{k(1)},\frac{d}{4}),\\ \overline{\widetilde{x}_{j}\widetilde{x}_{j+1}}&\subset\overline{B(x^{0}_{k(j)},\frac{d}{4})},\,|\widetilde{x}_{j}-\widetilde{x}_{j+1}|\leq\frac{d}{10},\quad\forall\,j=1,...,l-1,\\ \widetilde{x}_{l}&\in\partial B(x^{0}_{k(l)},\frac{d}{4}),\end{aligned}\right$.
and that
$\overline{x^{\prime}_{1}\widetilde{x}_{1}}\subset\Omega-\Omega_{\frac{d}{8}},\ \overline{\widetilde{x}_{l}x^{\prime}_{2}}\subset\Omega-\Omega_{\frac{d}{8}}.$
Indeed, given $0\leq j\leq l$ a segment $\overline{\widetilde{x}_{j}\widetilde{x}_{j+1}}$ (here we consider $\widetilde{x}_{0}:=x^{\prime}_{1}$, $\widetilde{x}_{l+1}:=x^{\prime}_{2}$), and a number $1\leq k\leq m_{1}$, the set $\overline{\widetilde{x}_{j}\widetilde{x}_{j+1}}\cap\partial B(x_{k}^{0},\frac{d}{4})$ is finite. In fact, after adding finite subsections of the line, we can assume that either the entire line segment $\overline{\widetilde{x}_{j}\widetilde{x}_{j+1}}$ lies within a ball $\overline{B(x_{k(j)}^{0},\frac{d}{4})}$ for some $1\leq k(j)\leq m_{1}$ or belongs to $\cap_{k=1}^{m_{1}}B^{c}(x_{k}^{0},\frac{d}{4})\subset\Omega-\Omega_{\frac{d}{8}}$.
Figure $2.$ Example of construction of the re-indexation of k
Now, we define
$\widetilde{\phi}(s)=\begin{cases}[1-2(l+1)s]x^{\prime}_{1}+2(l+1)s\widetilde{x}_{1},\ &\mathrm{for}\ 0\leq s\leq\frac{1}{2(l+1)},\\ [2-2(l+1)s]\widetilde{x}_{1}+[2(l+1)s-1]\widetilde{x}^{\prime}_{1},\ &\mathrm{for}\ \frac{1}{2(l+1)}\leq s\leq\frac{1}{l+1},\\ [1-(l+1)s+j]\widetilde{x}^{\prime}_{j}+[(l+1)s-j]\widetilde{x}^{\prime}_{j+1},\ &\mathrm{for}\ \frac{j}{l+1}\leq s\leq\frac{j+1}{l+1},\ 1\leq j\leq l-1,\\ [1-2(l+1)s+2l]\widetilde{x}^{\prime}_{l}+[2(l+1)s-2l]\widetilde{x}_{l},\ &\mathrm{for}\ \frac{l}{l+1}\leq s\leq\frac{2l+1}{2(l+1)},\\ [2-2(l+1)s+2l]\widetilde{x}_{l}+[2(l+1)s-2l-1]x^{\prime}_{2},\ &\mathrm{for}\ \frac{2l+1}{2(l+1)}\leq s\leq 1,\end{cases}$
where
(2.12) $\widetilde{x}^{\prime}_{j}:=\widetilde{x}_{j}-\frac{d}{2}n(x^{0}_{k(j)})\ ,\ 1\leq j\leq l.$
Now, we show that this new curve belongs to $\Omega-\Omega_{\frac{d}{8}}$. To do this, it suffices to show that for $1\leq j\leq l-1$,
$\overline{\widetilde{x}^{\prime}_{j}\widetilde{x}^{\prime}_{j+1}},\overline{\widetilde{x}_{1}\widetilde{x}^{\prime}_{1}},\overline{\widetilde{x}_{l}\widetilde{x}^{\prime}_{l}}\subset\Omega-\Omega_{\frac{d}{8}}.$
To show that $\overline{\widetilde{x}^{\prime}_{j}\widetilde{x}^{\prime}_{j+1}}\subset\Omega-\Omega_{\frac{d}{8}}$ for $1\leq j\leq l-1$, we re-parametrize $\widetilde{x}_{j}$ as follows:
[…]
We deduce that
[…]
$d((1-s)\widetilde{x}^{\prime}_{j}+s\widetilde{x}^{\prime}_{j+1},\partial\Omega)$
$\geq$ $d((1-s)\widetilde{x}^{\prime}_{j}+s\widetilde{x}^{\prime}_{j+1},\partial(\Omega\cap B(x^{0}_{k(j)},3d)))$
$\geq$ $\min\{d((1-s)\widetilde{x}^{\prime}_{j}+s\widetilde{x}^{\prime}_{j+1},\overline{\Omega}\cap\partial B(x^{0}_{k(j)},3d)),d((1-s)\widetilde{x}^{\prime}_{j}+s\widetilde{x}^{\prime}_{j+1},\partial\Omega\cap\overline{B(x^{0}_{k(j)},3d)})\}.$
Notice that
$d((1-s)\widetilde{x}^{\prime}_{j}+s\widetilde{x}^{\prime}_{j+1},\overline{\Omega}\cap\partial B(x^{0}_{k(j)},3d))\geq d(B(x^{0}_{k(j)},d),\partial B(x^{0}_{k(j)},3d))=2d.$
Since $|\widetilde{x}^{\prime}_{j}-x^{0}_{k(j)}|\leq|\widetilde{x}^{\prime}_{j}-\widetilde{x}_{j}|+|\widetilde{x}_{j}-x^{0}_{k(j)}|\leq\frac{3d}{4}$, we have $\widetilde{x}^{\prime}_{j}\in B(x^{0}_{k(j)},\frac{3d}{2})$. By Corollary 2.1 we have
(2.13) $d(\widetilde{x}^{\prime}_{j},\partial\Omega\cap\overline{B(x^{0}_{k(j)},3d)})\geq\frac{d}{4}$
We also notice that, using the same argument as (2.11), we have $|n(x^{0}_{k(j)})-n(x^{0}_{k(j+1)})|<\frac{1}{50}$ for $1\leq j\leq l-1$, from which we deduce that
(2.14) $|\widetilde{x}^{\prime}_{j}-\widetilde{x}^{\prime}_{j+1}|\leq|\widetilde{x}_{j}-\widetilde{x}_{j+1}|+\frac{d}{2}|n(x^{0}_{k(j)})-n(x^{0}_{k(j+1)})|\leq\frac{d}{10}+\frac{d}{100}<\frac{d}{9}.$
Hence, for any $0\leq s\leq 1$
$d((1-s)\widetilde{x}^{\prime}_{j}+s\widetilde{x}^{\prime}_{j+1},\partial\Omega\cap\overline{B(x^{0}_{k(j)},3d)})\geq\frac{d}{4}-\frac{sd}{9}\geq\frac{d}{8}.$
Therefore, we conclude that for any $0\leq s\leq 1$
$d((1-s)\widetilde{x}^{\prime}_{j}+s\widetilde{x}^{\prime}_{j+1},\partial\Omega)\geq\frac{d}{8}.$
We conclude that
(2.15) $\overline{\widetilde{x}^{\prime}_{j}\widetilde{x}^{\prime}_{j+1}}\in\Omega-\Omega_{\frac{d}{8}},$
for $2\leq j\leq l-2$.
For the case $j=1$ or $l$, we use Lemma 2.3. Notice that for $j=1$ or $l$, we have $\widetilde{x}_{j}\in\overline{B(x^{0}_{k(j)},\frac{d}{4})},\,\widetilde{x}^{\prime}_{j}\in\overline{B(x^{0}_{k(j)},\frac{3d}{4})}$. Since for $i=1$ or $l$
$\sqrt{\widetilde{x}^{2}_{j,1}+\widetilde{x}^{2}_{j,2}}\leq\frac{d}{4},$
and
$\frac{\widetilde{x}_{j}-\widetilde{x}^{\prime}_{j}}{|\widetilde{x}_{j}-\widetilde{x}^{\prime}_{j}|}\cdot n(x^{0}_{k(j)})=1,$
all the assumptions of Lemma 2.3 are satisfied. Hence, we deduce that
$d(\overline{\widetilde{x}_{j}\widetilde{x}^{\prime}_{j}},\partial\Omega)=d(\overline{\widetilde{x}_{j}\widetilde{x}^{\prime}_{j}},\partial\Omega\cap\overline{B(x^{0}_{k(j)},d)})=d(\widetilde{x}_{j},\partial\Omega\cap\overline{B(x^{0}_{k(j)},d)})\geq\frac{d}{8},$
for $j=1$ or $l$. We conclude that the curve $\widetilde{\phi}(s)$ is the desired curve between $x_{1}^{\prime}$ and $x_{2}^{\prime}$ which belongs to $\Omega-\Omega_{\frac{d}{8}}$.
Before we proceed, for given $\delta$, $d$, we define $\mathbb{Y}:=\mathbb{Y}_{1}\cup\mathbb{Y}_{2}$, where
$\mathbb{Y}_{1}:=\{y_{i}^{0}\mid 1\leq i\leq m_{1}\},\,\mathbb{Y}_{2}:=\{y_{i}^{0}\mid m_{1}+1\leq i\leq m_{1}+m_{2}\}.$
Now, we introduce a useful property of $\mathbb{Y}$:
Lemma 2.5.
Suppose that $\Omega$ satisfies Assumption A with $\delta$, $d$ as given in Lemma 2.1. Given $x^{\prime}_{1},x^{\prime}_{2}\in\Omega$. Then, there exist $\{y_{i}\}_{i=1}^{N(x^{\prime}_{1},x^{\prime}_{2})}\in\mathbb{Y}$ such that $\{x^{\prime}_{1},y_{1},y_{2},...,y_{N(x^{\prime}_{1},x^{\prime}_{2})},x^{\prime}_{2}\}$ forms a zigzag in $\Omega$, that is,
$\overline{x^{\prime}_{1}y_{1}}\in\Omega,\ \overline{y_{N(x^{\prime}_{1},x^{\prime}_{2})}x^{\prime}_{2}}\in\Omega,\ \overline{y_{i}y_{i+1}}\in\Omega,\ \forall 1\leq i\leq N(x^{\prime}_{1},x^{\prime}_{2})-1.$
Moreover, we have
$d(\overline{x^{\prime}_{1}y_{1}},\partial\Omega)\geq\min\left\{d(x^{\prime}_{1},\partial\Omega),\frac{d}{10}\right\},$
$d(\overline{x^{\prime}_{2}y_{N(x^{\prime}_{1},x^{\prime}_{2})}},\partial\Omega)\geq\min\left\{d(x^{\prime}_{2},\partial\Omega),\frac{d}{10}\right\},$
$d(\overline{y_{i},y_{i+1}},\partial\Omega)>\frac{d}{10},$
for $1\leq i\leq N(x^{\prime}_{1},x^{\prime}_{2})-1$. We call such a zigzag a “good zigzag” from $x^{\prime}_{1}$ to $x^{\prime}_{2}$ with number of segments $N(x^{\prime}_{1},x^{\prime}_{2})+1$.
Proof.
Given $x^{\prime}_{1},x^{\prime}_{2}\in\Omega$, we shall show that there exist $y^{\prime}_{1},y^{\prime}_{2}\in\mathbb{Y}$ such that
(2.16) $d(\overline{x^{\prime}_{i}y^{\prime}_{i}},\partial\Omega)\geq\min\left\{d(x^{\prime}_{i},\partial\Omega),\frac{d}{10}\right\},$
for $i=1,2$. Indeed, if $x^{\prime}_{i}\in\Omega_{\frac{d}{8}}$, then by Lemma 2.1, there exists $1\leq l(i)\leq m_{1}$ such that $x^{\prime}_{i}\in B(x_{l(i)}^{0},\frac{d}{4})$ and we can parameterize the boundary of $\Omega$ near $x^{0}_{l(i)}$ as follows:
$\partial\Omega\cap B(x_{l(i)}^{0},d)$
[…]
$\hskip 5.69046ptx_{l(i)}^{0}+u_{1}e_{l(i)}^{1}+u_{2}e_{l(i)}^{2}-\phi_{l(i)}(u_{1},u_{2})n(x^{0}_{l(i)})\in B(x_{l(i)}^{0},d)\}$
with $|\nabla\phi_{l(i)}|<\frac{1}{100}$. We shall show that
(2.17) $\overline{x^{\prime}_{i}y^{0}_{l(i)}}\subset B(x_{l(i)}^{0},d)\cap\Omega.$
Observe that, $\overline{x^{\prime}_{i}y^{0}_{l(i)}}\subset B(x_{l(i)}^{0},d)$ by the convexity of $B(x_{l(i)}^{0},d)$, so it suffices to show that $\overline{x^{\prime}_{i}y^{0}_{l(i)}}\subset\Omega$. To show this, we define $s^{\prime}:=\inf{\{s\mid sx^{\prime}_{i}+(1-s)y^{0}_{l(i)}\notin\Omega\}}$. If $\overline{x^{\prime}_{i}y^{0}_{l(i)}}\not\subset\Omega$, we have $0\leq s^{\prime}<1$. Notice that $0<s^{\prime}$ since $x^{\prime}_{i}\in\Omega$ and $\Omega$ are open. Next, we observe that $z:=s^{\prime}x^{\prime}_{i}+(1-s^{\prime})y^{0}_{l(i)}\in\partial\Omega$. Now, we recall the Lemma 2.1. We can re-parametrize $y^{0}_{l(i)}$, $z$ and $x^{\prime}_{i}$ as follows:
[…]
Since $x^{\prime}_{i}\in B(x_{l(i)}^{0},\frac{d}{4})$, we have
$|x^{\prime}_{i,1}|^{2}+|x^{\prime}_{i,2}|^{2}\leq|x^{\prime}_{i,1}|^{2}+|x^{\prime}_{i,2}|^{2}+|x^{\prime}_{i,3}|^{2}=|x^{\prime}_{i}-x^{0}_{l(i)}|^{2}\leq\frac{d^{2}}{16}.$
Now, we consider the following function $\psi:[0,\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}]\rightarrow\mathbb{R}$:
$\psi(\zeta):=\phi_{l(i)}(z_{1}+\frac{x^{\prime}_{i,1}-z_{1}}{\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}}\zeta,z_{2}+\frac{x^{\prime}_{i,2}-z_{2}}{\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}}\zeta).$
By the mean value theorem, there exists $\zeta^{\prime}$ such that
$\psi^{\prime}(\zeta^{\prime})$ $=\frac{\psi\left(\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}\right)-\psi(0)}{\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}-0}$
$\leq\frac{x^{\prime}_{i,3}-\phi_{l(i)}(z_{1},z_{2})}{\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}}.$
Here we used the fact that $x^{\prime}_{i,3}\geq\phi_{l(i)}(x^{\prime}_{i,1},x^{\prime}_{i,2})$ from the fact that $x^{\prime}_{i}\in\Omega$ and (2.2).
Note that $y^{0}_{l(i)}$, $z$ and $x^{\prime}_{i}$ belong to a straight line, so we have
$\frac{x^{\prime}_{i,3}-\phi_{l(i)}(z_{1},z_{2})}{\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}}$ $=\frac{\phi_{l(i)}(z_{1},z_{2})-\frac{3d}{4}}{\sqrt{|z_{1}|^{2}+|z_{2}|^{2}}}$
$\leq\frac{1}{100}-\frac{3}{4}=-\frac{74}{100}.$
Here we use the fact that $z\in\overline{B(x^{0}_{l(i)},d}).$
Finally, we get a contradiction to $\overline{x^{\prime}_{i}y^{0}_{l(i)}}\not\subset\Omega$ by computing:
$\nabla\phi_{l(i)}\left(z_{1}+\frac{x^{\prime}_{i,1}-z_{1}}{\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}}\zeta^{\prime},z_{2}+\frac{x^{\prime}_{i,2}-z_{2}}{\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}}\zeta^{\prime}\right)$
$\hskip 21.33955pt\cdot\left(\frac{x^{\prime}_{i,1}-z_{1}}{\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}},\frac{x^{\prime}_{i,2}-z_{2}}{\sqrt{|z_{1}-x^{\prime}_{i,1}|^{2}+|z_{2}-x^{\prime}_{i,2}|^{2}}}\right)$
$=$ $\psi^{\prime}(\zeta^{\prime})\leq-\frac{74}{100},$
which implies that $|\nabla\phi_{l(i)}|\geq\frac{74}{100}>\frac{1}{100}$. Hence, we deduce that (2.17) holds.
Hence, by (2.17), we have
$d(\overline{x^{\prime}_{i}y^{0}_{l(i)}},\partial\Omega)\geq d(\overline{x^{\prime}_{i}y^{0}_{l(i)}},\partial(\Omega\cap B(x^{0}_{l(i)},d))).$
Next, we show that
$d(\overline{x^{\prime}_{i}y^{0}_{l(i)}},\partial(\Omega\cap B(x^{0}_{l(i)},d)))\geq\min\{d(x^{\prime}_{i},\partial\Omega),\frac{d}{10}\}.$
Since $\partial(\Omega\cap B(x^{0}_{l(i)},d))\subset[\partial B(x^{0}_{l(i)},d)\cap\overline{\Omega}]\cup[\partial\Omega\cap\overline{B(x^{0}_{l(i)},d)})]$, we have
$\begin{split}&d(\overline{x^{\prime}_{i}y^{0}_{l(i)}},\partial(\Omega\cap B(x^{0}_{l(i)},d)))\\ \geq&d(\overline{x^{\prime}_{i}y^{0}_{l(i)}},[\partial B(x^{0}_{l(i)},d)\cap\overline{\Omega}]\cup[\partial\Omega\cap\overline{B(x^{0}_{l(i)},d)})])\\ =&\min\{d(\overline{x^{\prime}_{i}y^{0}_{l(i)}},\partial B(x^{0}_{l(i)},d)\cap\overline{\Omega}),d(\overline{x^{\prime}_{i}y^{0}_{l(i)}},\partial\Omega\cap\overline{B(x^{0}_{l(i)},d)})\}.\end{split}$
By the convexity of $B(x^{0}_{l(i)},d)$, we have
$d(\overline{x^{\prime}_{i}y^{0}_{l(i)}},\partial B(x^{0}_{l(i)},d)\cap\overline{\Omega})\geq\frac{d}{4}\geq\min\{d(x^{\prime}_{i},\partial\Omega),\frac{d}{10}\}.$
Next, we notice that
(2.18) $\frac{x^{\prime}_{i}-y^{0}_{l(i)}}{|x^{\prime}_{i}-y^{0}_{l(i)}|}\cdot n(x^{0}_{l(i)})\geq\frac{2\sqrt{2}}{3}.$
Now, we use Lemma 2.3. To show that all the assumption of Lemma 2.3 is satisfied, we notice that we have $x^{\prime}_{i}\in B(x^{0}_{l(i)},\frac{d}{4})$, $y^{0}_{l(i)}=x^{0}_{l(i)}-\frac{3d}{4}n(x^{0}_{l(i)})\in\overline{B(x^{0}_{l(i)},\frac{3d}{4})}$, and (2.18). As a result, we deduce that
(2.19) $\begin{split}d(\overline{x^{\prime}_{i}y^{0}_{l(i)}},\partial\Omega\cap\overline{B(x^{0}_{l(i)},d)})=d(x^{\prime}_{i},\partial\Omega\cap\overline{B(x^{0}_{l(i)},d)})\geq\min\{d(x^{\prime}_{i},\partial\Omega),\frac{d}{10}\}.\end{split}$
We showed that (2.16) holds when $x^{\prime}_{i}\in\Omega_{\frac{d}{8}}$.
In the other case, namely when $x^{\prime}_{i}\notin\Omega_{\frac{d}{8}}$, by Lemma 2.2, there exists $y\in\mathbb{Y}_{2}$ s.t. $x^{\prime}_{i}\in B(y,\frac{d}{16})$. By using the fact that $B\left(x^{\prime}_{i},\frac{d}{8}\right)\cup B\left(y,\frac{d}{8}\right)\subset\Omega$, and that
$\overline{x^{\prime}_{i}y}\subset B\left(x^{\prime}_{i},\frac{d}{8}\right)\cup B\left(y,\frac{d}{8}\right),$
we have
$d(\overline{x^{\prime}_{i}y},\partial\Omega)$ $\geq d\left(\overline{x^{\prime}_{i}y},\partial\left[B\left(x^{\prime}_{i},\frac{d}{8}\right)\cup B\left(y,\frac{d}{8}\right)\right]\right)$
$\geq\sqrt{(\frac{d}{8})^{2}-(\frac{d}{32})^{2}}>\frac{d}{10}.$
Next, we show that for any $w_{1},w_{2}\in\mathbb{Y}$, there exists a zigzag which satisfies the condition mentioned in Lemma 2.5.
We consider $w_{1},w_{2}\in\mathbb{Y}$ and define
(2.20) $\begin{split}&\mathbb{O}(w_{1}):=\bigg\{y\in\mathbb{Y}\mid\mathrm{There\ exists\ a\ zigzag}\,\ \{w_{1},u_{1},...,u_{s},y\}\ni\,u_{i}\in\mathbb{Y}\,\forall\,i=1,...,s,\\ &d(\overline{w_{1}u_{1}},\partial\Omega)>\frac{d}{10},d(\overline{u_{i}u_{i+1}},\partial\Omega)>\frac{d}{10}\forall\,1\leq i\leq s-1,d(\overline{u_{s}y},\partial\Omega)\>\frac{d}{10}\bigg\}.\end{split}$
If $B(\mathbb{O}(w_{1}),\frac{d}{16}):=\{\widetilde{x}\mid\exists\,\widetilde{y}\in\mathbb{O}(w_{1})\ \ni\ |\widetilde{x}-\widetilde{y}|<\frac{d}{16}\}$ covers $\Omega-\Omega_{\frac{d}{8}}$, then $\mathbb{O}(w_{1})$ covers $\mathbb{Y}$. We notice that $\mathbb{Y}\subset\Omega-\Omega_{\frac{d}{8}}$. For a given point $y\in\mathbb{Y}$, there exists a point $x\in\mathbb{O}(w_{1})\subset\mathbb{Y}$ with $|x-y|<\frac{d}{16}$ and a zigzag $\{w_{1},u_{1},...,u_{s},x\}$ such that
$d(\overline{w_{1}u_{1}},\partial\Omega)>\frac{d}{10},\ d(\overline{u_{i}u_{i+1}},\partial\Omega)>\frac{d}{10},\ \forall 1\leq i\leq s-1,\ d(\overline{u_{s}x},\partial\Omega)>\frac{d}{10}.$
We notice that
$d(\overline{xy},\partial\Omega)\geq d\left(\overline{xy},\partial\left[B\left(x,\frac{d}{8}\right)\cup B\left(y,\frac{d}{8}\right)\right]\right)\geq\sqrt{(\frac{d}{8})^{2}-(\frac{d}{32})^{2}}>\frac{d}{10}.$
Thus, the zigzag $\{w_{1},u_{1},...,u_{s},x,y\}$ satisfies the condition mentioned in Lemma 2.5, and
$y\in\mathbb{O}(w_{1}).$
On the other hand, suppose that $\Omega-\Omega_{\frac{d}{8}}\nsubseteq B\left(\mathbb{O}(w_{1}),\frac{d}{16}\right)$. We consider a point $\overline{w}_{3}\in\Omega-\Omega_{\frac{d}{8}}$ such that $d(\mathbb{O}(w_{1}),\overline{w}_{3})\geq d/16$. By Lemma 2.4, there exists a continuous function $\overline{\phi}:[0,1]\longrightarrow\Omega-\Omega_{\frac{d}{8}}$ such that $\overline{\phi}(0)=w_{1}$, $\overline{\phi}(1)=\overline{w}_{3}$.
Notice that $d(\mathbb{O}(w_{1}),\overline{\phi}(\cdot))$ is a continuous function from $[0,1]$ to $\mathbb{R}$ and that $d(\mathbb{O}(w_{1}),\overline{\phi}(0))=0,\ d(\mathbb{O}(w_{1}),\overline{\phi}(1))\geq\frac{d}{16}$. We define
$\overline{\xi}:=\sup\left\{0\leq s\leq 1\mid d(\mathbb{O}(w_{1}),\overline{\phi}(s))\leq\frac{d}{15}\right\}.$
Now, we define $w_{3}:=\overline{\phi}(\overline{\xi})$. Notice that $\frac{d}{16}\leq d(\mathbb{O}(w_{1}),w_{3})\leq\frac{d}{15}$. Since $\mathbb{O}(w_{1})$ is finite, there exists a point $z$ in $\mathbb{O}(w_{1})$ such that $\frac{d}{16}\leq d(z,w_{3})\leq\frac{d}{15}$.
By Lemma 2.2, since $w_{3}\in\Omega-\Omega_{\frac{d}{8}}$, there exists $w_{4}\in\mathbb{Y}_{2}$ such that $d(w_{3},w_{4})<\frac{d}{16}$, which implies $d(z,w_{4})\leq d(w_{3},w_{4})+d(z,w_{3})<\frac{31}{240}d$. Hence, we deduce that
$d(\overline{zw_{4}},\partial\Omega)\geq d\left(\overline{zw_{4}},\partial\left(B\left(z,\frac{d}{8}\right)\cup B\left(w_{4},\frac{d}{8}\right)\right)\right)\geq\sqrt{(\frac{d}{8})^{2}-(\frac{31d}{480})^{2}}>\frac{d}{10}.$
Thus, we have that $w_{4}\in\mathbb{O}(w_{1})$. This implies that $\frac{d}{16}\leq d(\mathbb{O}(w_{1}),w_{3})\leq d(w_{4},w_{3})<\frac{d}{16}$, which is a contradiction. We conclude that $\Omega-\Omega_{\frac{d}{8}}\subset B\left(\mathbb{O}(w_{1}),\frac{d}{16}\right)$. Therefore, $\mathbb{O}(w_{1})$ contains $\mathbb{Y}$, and $w_{2}\in\mathbb{O}(w_{1})$, which shows that there exists a zigzag between $w_{1}$ and $w_{2}$ satisfying the conditions mentioned in Lemma 2.5. ∎
Next, we introduce a lemma which is similar to Proposition 5.1 in [2]:
Lemma 2.6.
Suppose that $\Omega$ satisfies Assumption A. Given $\delta$, $d$ in Lemma 2.1, $v_{M}>0\in\mathbb{R}$, and $\epsilon>0$. Then there exists $t_{\epsilon}(v_{M})>0$ such that given $0<\tau_{2}\leq t_{\epsilon}(v_{M})$, one can find $l_{\epsilon}(\tau_{2})>0$ such that for all $1\leq i\leq m_{1}$, $(x,v)\in B(x^{0}_{i},d)\times\overline{B(0,v_{M})}$,
(2.21) $\forall\,s\in[0,\tau_{2}],\,X_{s}(x,v)\in\Omega_{l_{\epsilon}(\tau_{2})}\cap B(x_{i}^{0},d)$
implies
(2.22) $\forall\,s\in[0,t_{\epsilon}(v_{M})],\,|V_{s}(x,v)-v|<\epsilon.$
In fact, the numbers $t_{\epsilon}(v_{M})$ and $l_{\epsilon}(\tau_{2})$ can be chosen as follows.
(2.23) $t_{\epsilon}(v_{M}):=\frac{\epsilon}{8\widetilde{C}(\Omega)v_{M}^{2}},\,l_{\epsilon}(\tau_{2}):=\frac{\tau_{2}\epsilon}{64},$
where $\widetilde{C}(\Omega)$ is defined in Lemma 2.1.
Remark 2.3.
Unlike Proposition 5.1,[2], we do not need the lower bound of the velocity $v$.
Proof.
Let $\epsilon>0$ and $0\leq\tau_{2}\leq t_{\epsilon}(v_{M})$. Assume that
(2.24) $\forall s\in[0,\tau_{2}],X_{s}(x,v)\in\Omega_{l_{\epsilon}(\tau_{2})}\cap B(x_{i}^{0},d),$
where $t_{\epsilon}(v_{M})$ and $l_{\epsilon}(\tau_{2})$ are defined in (2.23).
For almost every $(x,v)\in B(x^{0}_{i},d)\times\overline{B(0,v_{M})}$, the trajectory $X_{s}(x,v)$ admits finitely many rebounds at times
$0<t_{1}<t_{2}<...<t_{r(x,v)}\leq t_{\epsilon}(v_{M}),$
with corresponding reflection angles
$\theta_{i}:=\frac{\pi}{2}-\arccos{|n(X_{t_{i}}(x,v))\cdot V_{t_{i}}(x,v)|},\,\forall 1\leq i\leq r(x,v).$
If $r(x,v)=0$, then we observe that $V_{s}(x,v)=v$ for all $s\in[0,t_{\epsilon}(v_{M})]$. Therefore, we assume that $r(x,v)\geq 1$.
We notice that the estimate
(2.25) $\forall s\in[0,t_{\epsilon}(v_{M})],\ |V_{s}(x,v)-v|<\epsilon$
holds if
(2.26) $\sum_{i=1}^{r(x,v)}2\sin{\theta_{i}}<\frac{\epsilon}{|v|}.$
Next, we show that for any $1\leq i\leq r(x,v)-1$
(2.27) $|X_{t_{i}}(x,v)-X_{t_{i-1}}(x,v)|\geq\frac{\sin{\theta_{i}}}{\widetilde{C}(\Omega)}.$
We invoke the Remark 2.1 and use the uniform interior sphere condition to deduce that for any $0\leq s\leq\frac{\sin{\theta_{i}}}{\widetilde{C}(\Omega)}$, we have
(2.28) $X_{t_{i}}(x,v)-s\frac{v}{|v|}\in B\left(X_{t_{i}}(x,v),\frac{1}{2\widetilde{C}(\Omega)}\right)\subset\Omega,$
which directly implies (2.27).
Therefore, it suffices to show that
(2.29) $\sum_{i=2}^{r(x,v)}|X_{t_{i}}(x,v)-X_{t_{i-1}}(x,v)|\leq\frac{\epsilon}{8|v|\widetilde{C}(\Omega)},$
and that
(2.30) $\sin{\theta_{1}}<\frac{\epsilon}{4|v|}.$
The (2.29) holds since we have
(2.31) $\sum_{i=2}^{r(x,v)}|X_{t_{i}}(x,v)-X_{t_{i-1}}(x,v)|\leq t_{\epsilon}(v_{M})|v|\leq\frac{\epsilon|v|}{8\widetilde{C}(\Omega)v_{M}^{2}}\leq\frac{\epsilon}{8|v|\widetilde{C}(\Omega)}.$
Here, we use the fact that $|v|\leq v_{M}$. For the (2.30), we again use the uniform interior sphere condition to deduce that for any $y\in\partial\Omega$ and $w\in\mathbb{R}^{3}$ with $n(y)\cdot w<0$, $\theta_{w}:=\frac{\pi}{2}-\arccos{|n(y)\cdot w|}$, we have
(2.32) $d\left(y+sw,\partial\Omega\right)\geq d\left(y+sw,\partial B\left(y-\frac{1}{2\widetilde{C}(\Omega)}n(y),\frac{1}{2\widetilde{C}(\Omega)}\right)\right)\geq\frac{s\sin{\theta_{w}}}{2}|w|,$
for any $0\leq s\leq\frac{\sin{\theta_{w}}}{2\widetilde{C}(\Omega)|w|}$.
Now, suppose that $\sin{\theta_{1}}\geq\frac{\epsilon}{4|v|}$. We consider the following estimate
(2.33) $\tau_{2}\leq t_{\epsilon}(v_{M}):=\frac{\epsilon}{8\widetilde{C}(\Omega)v_{M}^{2}}\leq\frac{1}{2\widetilde{C}(\Omega)|v|}\times\frac{\epsilon}{4|v|}\leq\frac{\sin{\theta_{1}}}{2\widetilde{C}(\Omega)|v|}.$
Notice that if $r(x,v)\geq 2$, then $t_{2}\geq\frac{\sin{\theta_{1}}}{\widetilde{C}(\Omega)|v|}>t_{\epsilon}$, which is a contradiction.
For the case $r(x,v)=1$, since $X_{t_{1}}(x,v)\in\partial\Omega$, we deduce from (2.32) and (2.33) , by setting $y:=X_{t_{1}}(x,v)$, that
(2.34) $d(X_{s}(x,v),\partial\Omega)\geq\frac{|s-t_{1}|\sin{\theta_{1}}}{8}|v|\geq\frac{|s-t_{1}|\epsilon}{32},$
for any $0\leq s\leq\tau_{2}$. Hence, we obtain
(2.35) $\sup_{0\leq s\leq\tau_{2}}d(X_{s}(x,v),\partial\Omega)\geq\sup_{0\leq s\leq\tau_{2}}\frac{|s-t_{1}|\epsilon}{32}\geq\frac{\frac{\tau_{2}}{2}\epsilon}{32}=l_{\epsilon}(\tau_{2}),$
which contradicts (2.24). We conclude the proof of Lemma 2.6.
Now, we introduce the following notation:
Definition 2.1.
Given $x,y\in\Omega$ and $d\in(0,1]$, we define
[…]
and
(2.37) $\operatorname{Conn}_{d}(\Omega):=\max\{\operatorname{Conn}_{d}(x,y)\mid x,y\in\Omega\}<+\infty.$
The last inequality is due to the fact that $\mathbb{Y}$ is finite.
3. Operator estimates and spreading properties
We introduce a new constant $n_{b}:=\int_{\mathbb{S}^{2}}b(\cos{\theta})\,d\sigma$ and quote a key lemma (Corollary 2.2, [11]), which is very useful throughout this article:
Lemma 3.1.
Given a measurable function g on $\mathbb{R}^{3}$, suppose that the collision operator satisfies Assumption B with $\nu<0$. Then, there exists $C_{g}^{L}>0$ which depends only on $n_{b}$, $C_{\Phi}$, $\rho_{g}$, and $e_{g}$ (and $l_{g,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$) such that
(3.1) $|L[g](v)|\leq C_{g}^{L}\left\langle v\right\rangle^{\gamma^{+}}.$
Definition 3.1.
We define $C_{L}:=\sup_{(t,x)\in[0,T]\times\mathbb{R}^{3}}C^{L}_{f(t,x,\cdot)}$.
We also quote a very useful lemma which describes the spreading property of the operator $Q$ (Lemma 2.4, [11]):
Lemma 3.2.
Suppose that the collision operator satisfies Assumption B with $\nu<0$. Then, there exists a constant $C>0$, which depends only on $\gamma,\ \nu$, and $b_{0}$, such that for any $v_{0}\in\mathbb{R}^{3}$, $0<r\leq R$, $\xi\in(0,1)$, we have
(3.2) $Q^{+}\left[\mathbf{1}_{B(v_{0},R)},\mathbf{1}_{B(v_{0},r)}\right]\geq Cl_{b}c_{\Phi}R^{3+\gamma}\xi^{\frac{1}{2}}\mathbf{1}_{B(v_{0},\sqrt{r^{2}+R^{2}}(1-\xi))},$
where $l_{b}:=\inf\limits_{\frac{1}{4}\pi\leq\theta\leq\frac{3}{4}\pi}b(\cos{\theta})$.
In particular, when $r=R=\delta$, we have
(3.3) $Q^{+}\left[\mathbf{1}_{B(v_{0},\delta)},\mathbf{1}_{B(v_{0},\delta)}\right]\geq Cl_{b}c_{\Phi}\delta^{3+\gamma}\xi^{\frac{1}{2}}\mathbf{1}_{B(v_{0},\delta\sqrt{2}(1-\xi))}.$
For simplicity, we define $C_{Q}:=Cl_{b}c_{\Phi}$
Now, we will adapt an important lemma introduced in Lemma 3.3 in [1]:
Proposition 3.1.
Let $\Omega\subset\mathbb{R}^{3}$ satisfy Assumption A and let the collision kernel $B$ satisfy Assumption B with $\nu<0$. Suppose that $f(t,x,v)$ is a continuous mild solution to (1.1)–(1.3) with $M>0$ and $0<E_{f}<\infty$. Then, there exist $x_{I}\in\Omega$ with $\Delta^{0}\in\left(0,\;\min\left\{\frac{d}{10},\,\frac{1}{2}d(x_{I},\partial\Omega)\right\}\right]$, $v_{I}\in\mathbb{R}^{3}$, $\alpha^{\prime}_{0}>0$, which depend on $\Omega,M,E_{f}$ and the modulus of continuity of $f_{0}$, such that for all $t\in[0,\Delta^{0}]$, $x\in B(x_{I},\Delta^{0})\cap\Omega$, we have
(3.4) $\forall v\in\mathbb{R}^{3},\ f(t,x,v)\geq\alpha^{\prime}_{0}\mathbf{1}_{B(v_{I},\Delta^{0})}(v).$
Remark 3.1.
The assumption here is weaker than that in Lemma 3.3 of [1]: we require $f(t,x,v)$ to be continuous only on $\Gamma_{conti}$, while Lemma 3.3 of [1] assumes continuity on the larger set $0,T)\times(\overline{\Omega}\times\mathbb{R}^{3}-\Gamma_{0})$.
Proof.
The following proof is an adaptation of the proof of Lemma 3.3 in [1].
First, we construct a lower bound at a point in $\Omega$. Note that by translation, we can always assume that $0\in\Omega$. Since $\Omega$ is bounded, we have for any $t>0$
(3.5) $\int_{\mathbb{R}^{3}}\int_{\Omega}(|x|^{2}+|v|^{2})f(t,x,v)\,dx\,dv\leq M(\text{\rm diam}(\Omega))^{2}+|\Omega|E_{f}<\infty.$
Define $R_{0}:=\max\left\{1,\sqrt{\frac{2(M(\text{\rm diam}(\Omega))^{2}+|\Omega|E_{f})}{M}}\right\}$. Notice that $\Omega\subset B(0,R_{0})$. Then, we have
(3.6) $\int_{B(0,R_{0})}\int_{B(0,R_{0})\cap\Omega}f_{0}(x,v)\,dx\,dv=\int_{B(0,R_{0})}\int_{\Omega}f_{0}(x,v)\,dx\,dv\geq\frac{M}{2}>0.$
Otherwise, we would have
(3.7) $\begin{split}&\int\int_{\Omega\times\mathbb{R}^{3}-(B(0,R_{0})\cap\Omega)\times B(0,R_{0})}(|x|^{2}+|v|^{2})f_{0}(x,v)\,dx\,dv\\ >&R_{0}^{2}\frac{M}{2}\geq M(\text{\rm diam}(\Omega))^{2}+|\Omega|E_{f},\end{split}$
which contradicts (3.5).
Hence, we can take a point $(x_{I},v_{I})\in\Omega\times B(0,R_{0})$ such that $f(0,x_{I},v_{I})>\frac{M}{3|B(0,R_{0})||\Omega|}$. By the uniform continuity of $f_{0}(x,v)$ on $\overline{\Omega}\times\overline{B(0,2R_{0})}$, there exist $0<\Delta_{I}<\min\left\{\frac{d(x_{I},\partial\Omega)}{2},R_{0}\right\}$ such that
$f_{0}(x,v)\geq\frac{M}{6|B(0,R_{0})||\Omega|}$
for any $(x,v)\in B(x_{I},\Delta_{I})\times B(v_{I},\Delta_{I})$.
Next, we observe that for any $(t,x,v)\in\left[0,\frac{\Delta_{I}}{4R_{0}}\right]\times B\left(x_{I},\frac{\Delta_{I}}{2}\right)\times B(v_{I},\Delta_{I})$, we have
$(x-tv,v)\in B(x_{I},\Delta_{I})\times B(v_{I},\Delta_{I})$
. As a result, we deduce from (1.12) and Lemma 3.1 that
(3.8) $\begin{split}f(t,x,v)&\geq f_{0}(X_{0,t}(x,v),v)\exp\left(-\int_{0}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right)\\ &\geq f_{0}(x-tv,v)e^{-tC_{L}\left\langle v\right\rangle^{\gamma^{+}}}\\ &\geq\frac{M}{6|B(0,R_{0})||\Omega|}e^{-\frac{\Delta_{I}C_{L}}{4R_{0}}\left\langle 2R_{0}\right\rangle^{\gamma^{+}}}\end{split}$
for any $(t,x,v)\in\left[0,\frac{\Delta_{I}}{4R_{0}}\right]\times B\left(x_{I},\frac{\Delta_{I}}{2}\right)\times B(v_{I},\Delta_{I})$, where the constant $C_{L}$ depends only on $n_{b}$$C_{\Phi}$ and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$).
We conclude the proof by defining
$\Delta^{0}:=\min\left\{\frac{\Delta_{I}}{4R_{0}},\frac{\Delta_{I}}{2},\frac{d}{10}\right\}.$
To demonstrate the fact that given an initial lower bound around a point $x\in\Omega$, multiple diluting lower bounds can be generated at the same point, we quote Lemma 3.3, [1]:
Proposition 3.2.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A. Let kernel $B$ satisfy Assumption B with $\nu<0$ and $\alpha\in[0,1]$. We consider a continuous mild solution $f$ of (1.1)–(1.3). Suppose that there exist $A>0$, $\Delta_{1},\ \Delta_{2}>0$, $(\tau,x^{\prime},v^{\prime})\in[0,T)\times\Omega\times\mathbb{R}^{3}$ such that $B(x^{\prime},\Delta_{2})\subset\Omega$ and
(3.9) $f(t,x,v)\geq A,\ \forall(t,x,v)\in[\tau,\tau+\Delta_{1}]\times B(x^{\prime},\Delta_{2})\times B(v^{\prime},\Delta_{2}).$
Then, we have for $n\in\mathbb{N}\cup\{0\}$, $t\in[\tau,\tau+\Delta_{1}]$, $x\in B(x^{\prime},\frac{\Delta_{2}}{2^{n}})$,
(3.10) $f(t,x,v)\geq\alpha_{n}(\tau,t,\Delta_{2},A,|v^{\prime}|)\mathbf{1}_{B(v^{\prime},r_{n}(\Delta_{2}))}(v),\,\forall\,v\in\mathbb{R}^{3},$
where the numbers $\{r_{n}(\Delta_{2})\}_{n=0}^{\infty}\in\mathbb{R}$, $\{t_{n}(t,\Delta_{2},|v^{\prime}|)\}_{n=0}^{\infty}\in\mathbb{R}$ and $\{\alpha_{n}(\tau,t,\Delta_{2},A,|v^{\prime}|)\}_{n=1}^{\infty}\in\mathbb{R}$ are defined as below:
$r_{0}(\Delta_{2}):=\Delta_{2},\ r_{n+1}(\Delta_{2}):=\frac{3\sqrt{2}}{4}r_{n}(\Delta_{2}),$
$t_{n}(\tau,t,\Delta_{2},|v^{\prime}|):=\max\left\{\tau,t-\frac{\Delta_{2}}{2^{n+1}(2r_{n}(\Delta_{2})+|v^{\prime}|)}\right\},$
$\alpha_{0}:=A,$
(3.11) $\begin{split}&\alpha_{n+1}(\tau,t,\Delta_{2},A,|v^{\prime}|)\\ :=&\frac{C_{Q}}{2}r_{n}^{3+\gamma}(\Delta_{2})\int_{t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}e^{-(t-s)C_{L}\left\langle 2r_{n}(\Delta_{2})+|v^{\prime}|\right\rangle^{\gamma^{+}}}\alpha_{n}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}ds\\ =&\frac{C_{Q}}{2}r_{n}^{3+\gamma}(\Delta_{2})\int_{0}^{t-t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}e^{-uC_{L}\left\langle 2r_{n}(\Delta_{2})+|v^{\prime}|\right\rangle^{\gamma^{+}}}\alpha_{n}(\tau,t-u,\Delta_{2},A,|v^{\prime}|)^{2}du.\end{split}$
Proof.
The following proof is also a summary of the proof of Lemma 3.3 in [1]. We prove the proposition by induction on $n$. The case $n=0$ is exactly the assumption. Assume that Proposition 3.2 holds for $n=k$. Given $t\in[\tau,\tau+\Delta_{1}]$, $x\in B(x^{\prime},\frac{\Delta_{2}}{2^{k+1}})$, $v\in B(0,|v^{\prime}|+2r_{k}(\Delta_{2}))$, we first notice that when $s\in\left[\max\left\{\tau,t-\frac{\Delta_{2}}{2^{k+1}(|v^{\prime}|+2r_{k}(\Delta_{2}))}\right\},t\right]$, we have
$|x^{\prime}-X_{s,t}(x,v)|=|x^{\prime}-x+tv-sv|\leq\frac{\Delta_{2}}{2^{k+1}}+|t-s|(|v^{\prime}|+2r_{k}(\Delta_{2}))\leq\frac{\Delta_{2}}{2^{k}},$
which implies that $X_{s,t}(x,v)\in B(x^{\prime},\frac{\Delta_{2}}{2^{k}})\subset\Omega$.
We consider the second term of the right hand side of (1.12) to obtain the following lower bound:
(3.12) $\begin{split}&f(t,x,v)\\ \geq&\int_{\tau}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v))\,ds^{\prime}\right)Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds.\end{split}$
Furthermore, by Lemma 3.1, we have
(3.13) $|L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)|\leq C_{L}\langle v\rangle^{\gamma^{+}}\leq C_{L}\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}},$
where the constant $C_{L}$ depends only on $n_{b}$$C_{\Phi}$ and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$).
Hence, the following estimate:
(3.14) $\begin{split}&f(t,x,v)\\ \geq&\int_{\tau}^{t}\exp\left(-(t-s)C_{L}\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds.\end{split}$
Since we have $|X_{s,t}(x,v)-x^{\prime}|\leq\frac{\Delta_{2}}{2^{k}}$, we can use the induction hypothesis:
(3.15) $f(s,X_{s,t}(x,v),w)\geq\alpha_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)1_{B(v^{\prime},r_{k}(\Delta_{2}))}(w),\ \forall w\in\mathbb{R}^{3}$
to deduce that
(3.16) $\begin{split}&f(t,x,v)\\ \geq&\int_{\max\left\{\tau,t-\frac{\Delta_{2}}{2^{k+1}(2r_{k}(\Delta_{2})+|v^{\prime}|)}\right\}}^{t}\exp\left(-(t-s)C_{L}\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\alpha^{2}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)Q^{+}[1_{B(v^{\prime},r_{k}(\Delta_{2}))}(\cdot),1_{B(v^{\prime},r_{k}(\Delta_{2}))}(\cdot)](v)\,ds.\end{split}$
Then we use Lemma 3.2 to ”spread” the lower bound:
For any $\xi\in(0,1)$ and $w\in\mathbb{R}^{3}$
(3.17) $Q^{+}\left[1_{B(v^{\prime},r_{k}(\Delta_{2}))},1_{B(v^{\prime},r_{k}(\Delta_{2}))}\right](w)\geq C_{Q}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}}1_{B(v^{\prime},r_{k}(\Delta_{2})\sqrt{2}(1-\xi))}(w).$
Hence, for any $\xi\in(0,1)$
(3.18) $\begin{split}&f(t,x,v)\\ \geq&\int_{\max\left\{\tau,t-\frac{\Delta_{2}}{2^{k+1}(2r_{k}(\Delta_{2})+|v^{\prime}|)}\right\}}^{t}\exp\left(-(t-s)C_{L}\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\alpha^{2}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)C_{Q}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}}1_{B(v^{\prime},r_{k}(\Delta_{2})\sqrt{2}(1-\xi))}(v)\,ds.\end{split}$
Set $\xi=\frac{1}{4}$ and notice that $B(v^{\prime},r_{k+1}(\Delta_{2})))\subset B(0,|v^{\prime}|+2r_{k}(\Delta_{2}))$. We conclude thus the proof of Proposition 3.2.
Remark 3.2.
Notice that by the last line of (3.11) and the fact that for any $\tau<t<t^{\prime}<\infty$
$t-t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)\leq t^{\prime}-t_{n}(\tau,t^{\prime},\Delta_{2},|v^{\prime}|).$
Hence, we notice that by induction on $n$, we can show that when $\tau<t$, $\alpha_{n+1}(\tau,t,\Delta_{2},A,|v^{\prime}|)$ is strictly decreasing with respect to $|v^{\prime}|$ and strictly increasing with respect to $t$.
Now, we show that we can generate lower bounds starting from one point $x^{\prime}\in\Omega$ and reaching another point $y\in\Omega$, as long as $\overline{x^{\prime}y}\subset\Omega$. Without loss of generality, we may assume that $\Delta_{2}\leq 2$.
Proposition 3.3.
Let $\Omega\subset\mathbb{R}^{3}$ satisfy Assumption A and let the collision kernel $B$ satisfy Assumption B with $\nu<0$, $\alpha\in[0,1]$. We consider a continuous mild solution $f$ of (1.1)–(1.3). Suppose that there exists $A>0$, $\Delta_{1},\Delta_{2}\in(0,1]$, $(\tau,x^{\prime},v^{\prime})\in[0,T)\times\Omega\times\mathbb{R}^{3}$ with $B(x^{\prime},\Delta_{2})\subset\Omega$ such that
$f(t,x,v)\geq A,\ \forall(t,x,v)\in[\tau,\tau+\Delta_{1}]\times B(x^{\prime},\Delta_{2})\times B(v^{\prime},\Delta_{2}).$
Then, given $\tau^{\prime}\in(\tau,\tau+\Delta_{1})$, $y\in\Omega$ with $\overline{x^{\prime}y}\subset\Omega$, we have that $B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)\subset\Omega$. Moreover, for all $x\in\Omega$, $v\in\mathbb{R}^{3}$
(3.19) $f(t,x,v)\geq\mathbb{B}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2},A,|v^{\prime}|)\mathbf{1}_{B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)\times B\left(\mathbb{V}^{\prime}(\tau,\tau^{\prime},y,x^{\prime}),\frac{\Delta_{2}}{2^{m+1}}\right)}(x,v)$
for any $m\geq m^{\prime}(\tau,\tau^{\prime},\Delta_{2},|v^{\prime}|,d(\overline{x^{\prime}y},\partial\Omega))$ and $\tau^{\prime}\leq t\leq\min\{\tau+\Delta_{1},\tau^{\prime}+\mathbb{D}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2})\}$.
Here, the functions $m^{\prime}(\tau,\tau^{\prime},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2})$, $\mathbb{B}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2},A,\mathfrak{y}_{3})$, $\mathbb{V}^{\prime}(\tau,\tau^{\prime},y,x^{\prime})$, and $\mathbb{D}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2})$ are defined as
$m^{\prime}(\tau,\tau^{\prime},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2}):=\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+\mathfrak{a}_{1}+1}{\Delta_{2}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta_{2}+1}{\mathfrak{y}_{2}}}\right\rceil\right\},$
$\mathbb{B}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2},A,\mathfrak{y}_{3}):=\alpha_{m}\left(\tau,\frac{\tau^{\prime}+\tau}{2},\Delta_{2},A,\mathfrak{y}_{3}\right)e^{-C_{L}\left\langle\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1\right\rangle^{\gamma^{+}}},$
$\mathbb{V}^{\prime}(\tau,\tau^{\prime},y,x^{\prime}):=\frac{2(y-x^{\prime})}{\tau^{\prime}-\tau},$
$\mathbb{D}^{\prime}(m,\tau,\tau^{\prime},\Delta_{2}):=\frac{\Delta_{2}}{2^{m+1}\left(\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1\right)}=\frac{\Delta_{2}(\tau^{\prime}-\tau)}{2^{m+1}(2d_{\Omega}+\tau^{\prime}-\tau)}.$
Here, $C_{L}$ is defined in Definition 3.1, $\alpha_{m}\left(\tau,\frac{\tau^{\prime}+\tau}{2},\Delta_{2},A,\mathfrak{y}_{3}\right)$ is as defined in Proposition 3.2, and we denote $d_{\Omega}:=\text{\rm diam}(\Omega)$ as the diameter of $\Omega$.
Proof.
For $0\leq\tau<\tau^{\prime}<\tau+\Delta_{1}$, we define $v^{\prime\prime}:=\frac{2(y-x^{\prime})}{\tau^{\prime}-\tau}$ and take $m\in\mathbb{N}\cup\{0\}$ with
$m\geq\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+|v^{\prime}|+1}{\Delta_{2}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta_{2}+1}{d(\overline{x^{\prime}y},\partial\Omega)}}\right\rceil\right\},$
then we see that
$r_{m}(\Delta_{2}):=(\frac{3\sqrt{2}}{4})^{m}\Delta_{2}\geq\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+|v^{\prime}|+1$
and $\frac{\Delta_{2}}{2^{m}}<d(\overline{x^{\prime}y},\partial\Omega)$. Here, the number $r_{m}(\Delta_{2})$ is as defined in Proposition 3.2 and we notice that $\frac{3\sqrt{2}}{4}>1$. Observe that we have
$\frac{\Delta_{2}}{2^{m+1}}\leq\frac{1}{2}d(\overline{x^{\prime}y},\partial\Omega)<d(y,\partial\Omega),$
so that $B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)\subset\Omega$.
Now, we prove the lower bound (3.19). As long as $x\in B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)$ and $v\in B\left(v^{\prime\prime},\min\left\{\frac{\Delta_{2}}{2^{m+1}(t-\tau)},1\right\}\right)$, for
$\tau^{\prime}\leq t\leq\min\left\{\tau+\Delta_{1},\tau^{\prime}+\frac{\Delta_{2}}{2^{m+1}\left(\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1\right)}\right\},$
we have
$\begin{split}&\left|x-\frac{t-\tau}{2}v-x^{\prime}\right|\\ \leq&|x-y|+\left|\frac{t-\tau}{2}v-(y-x^{\prime})\right|<\frac{\Delta_{2}}{2^{m+1}}+\left|\frac{t-\tau}{2}v-\frac{\tau^{\prime}-\tau}{2}v^{\prime\prime}\right|\\ \leq&\frac{\Delta_{2}}{2^{m+1}}+\left|\frac{t-\tau}{2}v-\frac{\tau^{\prime}-\tau}{2}v\right|+\left|\frac{\tau^{\prime}-\tau}{2}v-\frac{\tau^{\prime}-\tau}{2}v^{\prime\prime}\right|\\ <&\frac{\Delta_{2}}{2^{m+1}}+\left|\frac{t-\tau^{\prime}}{2}v\right|+\frac{\Delta_{2}}{2^{m+2}}\\ <&\frac{\Delta_{2}}{2^{m+1}}+\frac{\Delta_{2}}{2^{m+2}}+\frac{\Delta_{2}}{2^{m+2}}\\ =&\frac{\Delta_{2}}{2^{m}},\end{split}$
by which we deduce that $x-\frac{t-\tau}{2}v\in B\left(x^{\prime},\frac{\Delta_{2}}{2^{m}}\right)$. Notice that we have
$r_{m}(\Delta_{2})\geq|v^{\prime}|+\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1\geq|v^{\prime}|+|v^{\prime\prime}|+1>|v^{\prime}|+|v|,$
so we have $v\in B(v^{\prime},r_{m}(\Delta_{2}))$.
Note that for any $0\leq s\leq 1$, we have
$\begin{split}&|sx+(1-s)\left(x-\frac{t-\tau}{2}v\right)-(sy+(1-s)x^{\prime})|\\ \leq&s|x-y|+(1-s)\left|x-\frac{t-\tau}{2}v-x^{\prime}\right|\\ \leq&\frac{\Delta_{2}}{2^{m}}<d(\overline{x^{\prime}y},\partial\Omega).\end{split}$
Hence, the segment $\{x-\sigma v|\ 0\leq\sigma\leq\frac{t-\tau}{2}\}$ also lies within $\Omega$. From this, we deduce by using the first term of the right hand side of (1.12), Lemma 3.1, and Proposition 3.2 that
(3.20) $\begin{split}f(t,x,v)&\geq f\left(\frac{t+\tau}{2},x-\frac{t-\tau}{2}v,v\right)e^{-\frac{t-\tau}{2}C_{L}\left\langle v\right\rangle^{\gamma^{+}}},\\ &\geq\alpha_{m}\left(\tau,\frac{t+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)\mathbf{1}_{B\left(x^{\prime},\frac{\Delta_{2}}{2^{m}}\right)\times B(v^{\prime},r_{m}(\Delta_{2}))}(x-\frac{t-\tau}{2}v,v)e^{-\frac{t-\tau}{2}C_{L}\left\langle v\right\rangle^{\gamma^{+}}}\\ &=\alpha_{m}\left(\tau,\frac{t+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)e^{-\frac{t-\tau}{2}C_{L}\left\langle v\right\rangle^{\gamma^{+}}},\end{split}$
where we use the fact that $x-\frac{t-\tau}{2}v\in B(x^{\prime},\frac{\Delta_{2}}{2^{m}})$ and $v\in B(v^{\prime},r_{m}(\Delta_{2}))$. Hence, we have the following inequality for $\tau^{\prime}\leq t\leq\min\bigg\{\tau+\Delta_{1},\tau^{\prime}+\frac{\Delta_{2}}{2^{m+1}(\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1)}\bigg\}$, $x\in\Omega$, $v\in\mathbb{R}^{3}$ :
$f(t,x,v)$
$\geq$ $\alpha_{m}\left(\tau,\frac{t+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)e^{-\frac{t-\tau}{2}C_{L}\left\langle v\right\rangle^{\gamma^{+}}}\mathbf{1}_{B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)\times B\left(v^{\prime\prime},\min\left\{\frac{\Delta_{2}}{2^{m+1}(t-\tau)},1\right\}\right)}(x,v)$
(3.21) $\geq$ $\alpha_{m}\left(\tau,\frac{\tau^{\prime}+\tau}{2},\Delta_{2},A,|v^{\prime}|\right)e^{-C_{L}\left\langle\frac{2d_{\Omega}}{\tau^{\prime}-\tau}+1\right\rangle^{\gamma^{+}}}\mathbf{1}_{B\left(y,\frac{\Delta_{2}}{2^{m+1}}\right)\times B\left(v^{\prime\prime},\frac{\Delta_{2}}{2^{m+1}}\right)}(x,v).$
Here, we use Remark 3.2 in the last inequality above. The line (3.21) corresponds to (3.19). Thus, we have shown the statement of Proposition 3.3.
The following corollary can be derived by modifying the starting time $\tau^{\prime}$ in Proposition 3.3:
Corollary 3.4.
Let $\Omega\subset\mathbb{R}^{3}$ satisfy Assumption A and let the collision kernel $B$ satisfy Assumption B with $\nu<0$, $\alpha\in[0,1]$. We consider a continuous mild solution $f$ of (1.1)–(1.3). Suppose that there exists $A>0$, $\Delta_{1},\Delta_{1}^{\prime},\Delta_{2}\in(0,\frac{1}{2}]$, $(\tau,x^{\prime},v^{\prime})\in[0,T)\times\Omega\times\mathbb{R}^{3}$ with $B(x^{\prime},\Delta_{2})\subset\Omega$ such that
(3.22) $f(t,x,v)\geq A,\ \forall(t,x,v)\in[\tau,\tau+\Delta_{1}+\Delta_{1}^{\prime}]\times B(x^{\prime},\Delta_{2})\times B(v^{\prime},\Delta_{2}).$
Then, given $y\in\Omega$ such that $\overline{x^{\prime}y}\subset\Omega$, we have $B(y,\frac{\Delta_{2}}{2^{m+1}})\subset\Omega$. Furthermore, for any $x\in\Omega,v\in\mathbb{R}^{3}$
$f(t,x,v)\geq\mathbb{B}(m,\tau,\Delta_{1},\Delta_{2},A,|v^{\prime}|)\mathbf{1}_{B(y,\frac{\Delta_{2}}{2^{m+1}})\times B\left(\mathbb{V}(\Delta_{1},\Delta_{2},\tau,m,y,x^{\prime}),\frac{\Delta_{2}}{2^{m+1}}\right)}(x,v)$
for any $m\geq\mathbb{M}(\Delta_{1},\Delta_{2},|v^{\prime}|,d(\overline{x^{\prime}y},\partial\Omega))$ and $\mathbb{T}(\Delta_{1},\Delta_{2},\tau,m)\leq t\leq\min\{\tau+\Delta_{1}+\Delta_{1}^{\prime},\tau+\Delta_{1}+\mathbb{D}(\Delta_{1},\Delta_{2},m)\}$.
Here, the functions $\mathbb{M}(\Delta_{1},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2})$, $\mathbb{B}(m,\tau,\tau^{\prime},\Delta_{2},A,\mathfrak{y}_{3})$, $\mathbb{D}(\Delta_{1},\Delta_{2},m)$, $\mathbb{T}(\Delta_{1},\Delta_{2},\tau,m)$, and $\mathbb{V}(\Delta_{1},\Delta_{2},\tau,m,y,x^{\prime})$ are defined as
$\mathbb{M}(\Delta_{1},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2}):=\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\Delta_{1}}+\mathfrak{y}_{1}+1}{\Delta_{2}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta_{2}+1}{\mathfrak{y}_{2}}}\right\rceil\right\},$
$\mathbb{B}(m,\tau,\Delta_{1},\Delta_{2},A,\mathfrak{y}_{3}):=\alpha_{m}\left(\tau,\tau+\frac{\Delta_{1}}{4},\Delta_{2},A,\mathfrak{y}_{3}\right)e^{-C_{L}\left\langle\frac{4d_{\Omega}}{\Delta_{1}}+1\right\rangle^{\gamma^{+}}},$
$\mathbb{D}(\Delta_{1},\Delta_{2},m):=\frac{\Delta_{1}\Delta_{2}}{2^{m+2}(4d_{\Omega}+\Delta_{1})},$
$\mathbb{T}(\Delta_{1},\Delta_{2},\tau,m):=\max\left\{\tau+\frac{\Delta_{1}}{2},\Delta_{1}+\tau-\frac{\Delta_{1}\Delta_{2}}{2^{m+2}(4d_{\Omega}+\Delta_{1})}\right\},$
$\mathbb{V}(\Delta_{1},\Delta_{2},\tau,m,y,x^{\prime}):=\frac{2(y-x^{\prime})}{\mathbb{T}(\Delta_{1},\Delta_{2},\tau,m)-\tau}=\frac{2(y-x^{\prime})}{\max\left\{\frac{\Delta_{1}}{2},\Delta_{1}-\frac{\Delta_{1}\Delta_{2}}{2^{m+2}(4d_{\Omega}+\Delta_{1})}\right\}}.$
Remark 3.3.
By Remark 3.2, we notice that
  1. (1)
    $\mathbb{B}(m,\tau,\Delta_{1},\Delta_{2},A,\mathfrak{y}_{3})$ is strictly decreasing in $\mathfrak{y}_{3}$.
  2. (2)
    $\mathbb{M}(\Delta_{1},\Delta_{2},\mathfrak{y}_{1},\mathfrak{y}_{2})$ is increasing in $\mathfrak{y}_{1}$ and decreasing in $\mathfrak{y}_{2}$.
  3. (3)
    $\mathbb{T}(\Delta_{1},\Delta_{2},\tau,m)$ is increasing in $m$.
  4. (4)
    $|\mathbb{V}(\Delta_{1},\Delta_{2},\tau,m,y,x^{\prime})|$ is decreasing in $m$ with the upper bound $\frac{4d_{\Omega}}{\Delta_{1}}$.
Now, we will use the multiple lower bounds around the initial point $x_{I}$ from Proposition 3.1 to derive lower bounds around the points $\{y^{0}_{i}\}$.
Proposition 3.5.
Let $\Omega\subset\mathbb{R}^{3}$ satisfy Assumption A and let the collision kernel $B$ satisfy Assumption B with $\nu<0$. Fix $d$ with $0<d<\min\{1,\delta\}$ as in Lemma 2.1 and a collection of points $\{y_{i}^{0}\}_{i=1}^{m_{1}+m_{2}}\subset\Omega$ defined in Lemmas 2.1 and 2.2. Let $x_{I}$ be as defined in Proposition 3.1.
Suppose that, for fixed $\alpha\in[0,1]$, there exists a continuous mild solution to (1.1)–(1.3). Then, for any $\tau^{\prime\prime}\in(0,\Delta^{0})$, with $\Delta^{0}\in(0,\min\{\frac{d}{10},\frac{1}{2}d(x_{I},\partial\Omega)\}]$ as in Proposition 3.1, there exist positive numbers $\mathfrak{R}(\tau^{\prime\prime})$, $\mathfrak{B}(\tau^{\prime\prime})$, and $\mathfrak{d}(\tau^{\prime\prime})$, depending on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$), and a family of velocity vectors
$\{v_{i}(\tau^{\prime\prime})\}_{i=1}^{m_{1}+m_{2}}\subset B(0,\mathfrak{R}(\tau^{\prime\prime}))$
such that
$\bigcup_{i=1}^{m_{1}+m_{2}}B\left(y_{i}^{0},\mathfrak{d}(\tau^{\prime\prime})\right)\subset\Omega.$
Moreover, for each $1\leq i\leq m_{1}+m_{2}$, we have
(3.23) $f(t,x,v)\geq\ \mathfrak{B}(\tau^{\prime\prime})\mathbf{1}_{B\left(y_{i}^{0},\mathfrak{d}(\tau^{\prime\prime})\right)\times B\left(v_{i}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime})\right)}(x,v),$
for all
$\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})\leq t\leq\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime}).$
Here, we use the same radius $\mathfrak{d}(\tau^{\prime\prime})$ for the space and velocity balls and for the time window.
Proof.
First, by Proposition 3.1, we have
(3.24) $\forall v\in\mathbb{R}^{3},\ f(t,x,v)\geq\alpha^{\prime}_{0}\mathbf{1}_{B(v_{I},\Delta^{0})}(v),$
for all $t\in[0,\Delta^{0}]$, $x\in B(x_{I},\Delta^{0})$. Here, the quantities $\alpha^{\prime}_{0}$, $x_{I}$, $v_{I}$, $\Delta^{0}$ are defined in Proposition 3.1. Note that $B(x_{I},\Delta^{0})\subset\Omega$. Consider a good zigzag $\{x_{I},y_{1},y_{2},...,y_{\operatorname{Conn}_{d}(x_{I},y_{i}^{0})}=y_{i}^{0}\}$ in $\Omega$ (cf. Lemma 2.5).
Given $0<\tau^{\prime\prime}<\Delta^{0}$, for any $j\in\mathbb{N}$, and a $j$-tuple
$\mathfrak{M}_{j}:=(\mathfrak{m}_{1},\mathfrak{m}_{2},...,\mathfrak{m}_{j})\in\mathbb{N}^{j},$
we define $\mathbb{B}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|)$, $\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})$, $\mathbb{D}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})$, $\mathbb{V}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})$, $\mathbb{U}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})$, $\mathfrak{n}_{j}$ as
$\mathbb{T}_{1}(\mathfrak{M}_{1},\tau^{\prime\prime},\Delta^{0}):=\mathbb{T}(\tau^{\prime\prime},\Delta^{0},0,\mathfrak{m}_{1}),$
$\mathbb{D}_{1}(\mathfrak{M}_{1},\tau^{\prime\prime},\Delta^{0}):=\mathbb{D}(\tau^{\prime\prime},\Delta^{0},\mathfrak{m}_{1}),$
$\mathbb{V}_{1}(\mathfrak{M}_{1},\tau^{\prime\prime},\Delta^{0}):=\mathbb{V}(\tau^{\prime\prime},\Delta^{0},0,\mathfrak{m}_{1},y_{1},x_{I}),$
$\mathbb{B}_{1}(\mathfrak{M}_{1},\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|):=\mathbb{B}(\mathfrak{m}_{1},0,\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|),$
$\mathbb{U}_{1}:=\Delta^{0}.$
$\mathbb{T}_{j+1}(\mathfrak{M}_{j+1},\tau^{\prime\prime},\Delta^{0}):=\mathbb{T}\left(\tau^{\prime\prime}-\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{\sum_{k=1}^{j}(\mathfrak{m}_{k}+1)}},\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),\mathfrak{m}_{j+1}\right),$
$\mathbb{D}_{j+1}(\mathfrak{M}_{j+1},\tau^{\prime\prime},\Delta^{0})=\mathbb{D}\left(\tau^{\prime\prime}-\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{\sum_{k=1}^{j}(\mathfrak{m}_{k}+1)}},\mathfrak{m}_{j+1}\right),$
$\mathbb{V}_{j+1}(\mathfrak{M}_{j+1},\tau^{\prime\prime},\Delta^{0}):=\mathbb{V}\left(\tau^{\prime\prime}-\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{\sum_{k=1}^{j}(\mathfrak{m}_{k}+1)}},\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),\mathfrak{m}_{j+1},y_{j+1},y_{j}\right)$
$\hskip 93.89418pt=\frac{2(y_{j+1}-y_{j})}{\mathbb{T}_{j+1}(\mathfrak{M}_{j+1},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})},$
$\mathbb{B}_{j+1}(\mathfrak{M}_{j+1},\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|)$
$\hskip 28.45274pt:=\mathbb{B}\Bigg(\mathfrak{m}_{j+1},\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),\tau^{\prime\prime}-\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{\sum_{k=1}^{j}(\mathfrak{m}_{k}+1)}},$
$\hskip 62.59596pt\mathbb{B}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|),\frac{2d_{\Omega}}{\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{j-1}(\mathfrak{M}_{j-1},\tau^{\prime\prime},\Delta^{0})}\Bigg),$
$\mathbb{U}_{j+1}(\mathfrak{M}_{j+1},\tau^{\prime\prime},\Delta^{0}):=\min\{\mathbb{U}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),\tau^{\prime\prime}+\mathbb{D}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})\}.$
Here, the notation $\mathbb{B}$, $\mathbb{T}$, $\mathbb{D}$, $\mathbb{V}$ are defined in Corollary 3.4. We also define $\mathbb{T}_{0}:=0$ by convention. By Remark 3.4, we have $\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})<\tau^{\prime\prime}$ and that all denominators $\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{j-1}(\mathfrak{M}_{j}-1,\tau^{\prime\prime},\Delta^{0})$ are positive. We also define
$\mathfrak{n}_{1}:=\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\tau^{\prime\prime}}+|v_{I}|+1}{\Delta^{0}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta^{0}+1}{\min\{d(x_{I},\partial\Omega),\frac{d}{10}\}}}\right\rceil\right\},$
$\mathfrak{N}_{j}(m):=(\mathfrak{n}_{1},\mathfrak{n}_{2},...,\mathfrak{n}_{j-1},m),\,\mathfrak{N}_{j}:=\mathfrak{N}_{j}(\mathfrak{n}_{j}),$
$\mathfrak{n}_{j+1}:=\max\Bigg\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\tau^{\prime\prime}-\mathbb{T}_{j}(\mathfrak{N}_{j},\tau^{\prime\prime},\Delta^{0})}+\frac{2d_{\Omega}}{\mathbb{T}_{j}(\mathfrak{N}_{j},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{j-1}(\mathfrak{N}_{j-1},\tau^{\prime\prime},\Delta^{0})}+1}{\frac{\Delta^{0}}{2^{\sum_{k=1}^{j}(\mathfrak{n}_{k}+1)}}}\right)\right\rceil,$
$\hskip 68.28644pt\left\lceil\log_{2}{\frac{\left(\frac{\Delta^{0}}{2^{\sum_{k=1}^{j}(\mathfrak{n}_{k}+1)}}+1\right)}{\frac{d}{10}}}\right\rceil\Bigg\}.$
Now, we claim that for $1\leq j\leq\operatorname{Conn}_{d}(x_{I},y_{i}^{0})$, we have
(3.25) $\begin{split}&f(t,x,v)\geq\ \mathbb{B}_{j}(\mathfrak{N}_{j}(m),\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|)\\ &\hskip 56.9055pt\mathbf{1}_{B\left(y_{j},\frac{\Delta^{0}}{2^{m+1+\sum_{k=1}^{j-1}(\mathfrak{n}_{k}+1)}}\right)\times B\left(\mathbb{V}_{j}(\mathfrak{N}_{j}(m),\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{m+1+\sum_{k=1}^{j-1}(\mathfrak{n}_{k}+1)}}\right)}(x,v),\end{split}$
for any
$\begin{split}&\mathbb{T}_{j}(\mathfrak{N}_{j}(m),\tau^{\prime\prime},\Delta^{0})\leq t\\ \leq&\min\{\mathbb{U}_{j}(\mathfrak{N}_{j}(m),\tau^{\prime\prime},\Delta^{0}),\tau^{\prime\prime}+\mathbb{D}_{j}(\mathfrak{N}_{j}(m),\tau^{\prime\prime},\Delta^{0})\}\end{split}$
and $m\geq\mathfrak{n}_{j}$. Here, $\alpha^{\prime}_{0}$ is the constant defined in Proposition 3.1.
This will be proved by induction on $j$.
Step 1: base case: propagation of the lower bound from $x_{I}$ to $y_{1}$.
For the case $j=1$, we consider the Corollary 3.4 with $A=\alpha^{\prime}_{0}$, $\tau=0$, $\Delta_{1}=\tau^{\prime\prime}$, $\Delta_{1}^{\prime}=\Delta^{0}-\tau^{\prime\prime}$, $x^{\prime}=x_{I}$, $y=y_{1}$, $v^{\prime}=v_{I}$, and $\Delta_{2}=\Delta^{0}$. Notice that all the assumptions of Corollary 3.4 are satisfied since $B(x_{I},\Delta^{0})\subset B(x_{I},\frac{1}{2}d(x_{I},\partial\Omega))\subset\Omega$ and $\overline{x_{I}y_{1}}\in\Omega$ and the lower bound (3.22) holds. Thus we derive a lower bound at $y_{1}$ from $x_{I}$ by using Corollary 3.4:
$f(t,x,v)\geq\mathbb{B}(m,0,\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|)\mathbf{1}_{B\left(y_{1},\frac{\Delta^{0}}{2^{m+1}}\right)\times B\left(\mathbb{V}(\tau^{\prime\prime},\Delta^{0},0,m,y_{1},x_{I}),\frac{\Delta^{0}}{2^{m+1}}\right)}(x,v)$
for any $\mathbb{T}(\tau^{\prime\prime},\Delta^{0},0,m)\leq t\leq\min\big\{\Delta^{0},\tau^{\prime\prime}+\mathbb{D}(\tau^{\prime\prime},\Delta^{0},m)\big\}$
and $m\geq\mathbb{M}(\tau^{\prime\prime},\Delta^{0},|v_{I}|,d(\overline{x_{I}y_{1}},\partial\Omega))$. Notice that
$\mathbb{M}(\tau^{\prime\prime},\Delta^{0},|v_{I}|,d(\overline{x_{I}y_{1}},\partial\Omega))$
$=$ $\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\tau^{\prime\prime}}+|v_{I}|+1}{\Delta^{0}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta^{0}+1}{d(\overline{x_{I}y_{1}},\partial\Omega)}}\right\rceil\right\}$
$\leq$ $\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\tau^{\prime\prime}}+|v_{I}|+1}{\Delta^{0}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta^{0}+1}{\min\{d(x_{I},\partial\Omega),\frac{d}{10}\}}}\right\rceil\right\}$
$=$ $\mathfrak{n}_{1},$
which implies that as long as $m\geq\mathfrak{n}_{1}$, we have $m\geq\mathbb{M}(\tau^{\prime\prime},\Delta^{0},|v_{I}|,d(\overline{x_{I}y_{1}},\partial\Omega))$, and this finishes the case for $j=1$.
Step 2: Induction: propagation of the lower bound from $y_{k}$ to $y_{k+1}$.
Now, suppose that the case $j=k$ holds, i.e.,
(3.26) $\begin{split}&f(t,x,v)\geq\ \mathbb{B}_{k}(\mathfrak{N}_{k}(\widetilde{m}),\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|)\\ &\hskip 56.9055pt\mathbf{1}_{B\left(y_{k},\frac{\Delta^{0}}{2^{\widetilde{m}+1+\sum_{l=1}^{k-1}(\mathfrak{n}_{l}+1)}}\right)\times B\left(\mathbb{V}_{k}(\mathfrak{N}_{k}(\widetilde{m}),\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{\widetilde{m}+1+\sum_{l=1}^{k-1}(\mathfrak{n}_{l}+1)}}\right)}(x,v),\end{split}$
for any
$\begin{split}&\mathbb{T}_{k}(\mathfrak{N}_{k}(\widetilde{m}),\tau^{\prime\prime},\Delta^{0})\leq t\\ \leq&\min\{\mathbb{U}_{k}(\mathfrak{N}_{k}(\widetilde{m}),\tau^{\prime\prime},\Delta^{0}),\tau^{\prime\prime}+\mathbb{D}_{k}(\mathfrak{N}_{k}(\widetilde{m}),\tau^{\prime\prime},\Delta^{0})\}\end{split}$
and $\widetilde{m}\geq\mathfrak{n}_{k}$.
By taking $\widetilde{m}=\mathfrak{n}_{k}$, we use Corollary 3.4 again by considering
$A$ $=\mathbb{B}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|),$
$\tau$ $=\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0}),$
$\Delta_{1}$ $=\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0}),$
$\Delta_{1}^{\prime}$ $=\min\{\mathbb{U}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0}),\tau^{\prime\prime}+\mathbb{D}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})\}-\tau^{\prime\prime},$
$x^{\prime}$ $=y_{k},\,y=y_{k+1},$
$v^{\prime}$ $=\mathbb{V}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0}),$
$\Delta_{2}$ $=\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}}.$
We notice that by Remark 3.4, we have $\Delta_{1}>0$ and $\Delta_{1}^{\prime}>0$.
Also, all the assumptions of Corollary 3.4 are satisfied again since
$B\left(y_{k},\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}}\right)\subset B\left(y_{k},\frac{d}{10}\right)\subset\Omega,$
$\overline{y_{k}y_{k+1}}\in\Omega$, $\Delta_{1}\leq\Delta^{0}\leq\frac{1}{2}$, and
$\Delta_{1}^{\prime}$ $=\min\{\mathbb{U}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0}),\tau^{\prime\prime}+\mathbb{D}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})\}-\tau^{\prime\prime}$
$\leq\mathbb{U}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})-\tau^{\prime\prime}$
$\leq\mathbb{U}_{1}-\tau^{\prime\prime}\leq\Delta^{0}-\tau^{\prime\prime}\leq\Delta^{0}\leq\frac{1}{2}.$
Here, we recall that $\mathbb{U}_{1}=\Delta^{0}$. Hence, by noticing that
$|\mathbb{V}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})|$
$=$ $\frac{2|y_{k}-y_{k-1}|}{\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{k-1}(\mathfrak{N}_{k-1},\tau^{\prime\prime},\Delta^{0})}$
$\leq$ $\frac{2d_{\Omega}}{\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{k-1}(\mathfrak{N}_{k-1},\tau^{\prime\prime},\Delta^{0})},$
we have
(3.27) $\begin{split}&f(t,x,v)\\ \geq&\ \mathbb{B}\Bigg(m,\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0}),\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0}),\\ &\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}},\mathbb{B}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|),|\mathbb{V}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})|\Bigg)\\ &\mathbf{1}_{B\left(y_{k+1},\frac{\Delta^{0}}{2^{m+1+\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}}\right)\times B\left(\mathbb{V}_{k+1}(\mathfrak{N}_{k+1}(m),\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{m+1+\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}}\right)}(x,v)\\ \geq&\mathbb{B}_{k+1}(\mathfrak{N}_{k+1}(m),\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|)\\ &\mathbf{1}_{B\left(y_{k+1},\frac{\Delta^{0}}{2^{m+1+\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}}\right)\times B\left(\mathbb{V}_{k+1}(\mathfrak{N}_{k+1}(m),\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{m+1+\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}}\right)}(x,v),\end{split}$
for any
$m\geq\mathbb{M}(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0}),$
$\hskip 14.22636pt\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}},|\mathbb{V}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})|,d(\overline{y_{k}y_{k+1}},\partial\Omega)).$
We also notice that
$\mathbb{M}\Bigg(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0}),$
$\hskip 17.07182pt\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}},|\mathbb{V}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})|,d(\overline{y_{k}y_{k+1}},\partial\Omega)\Bigg)$
$\leq$ $\max\Bigg\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})}+\frac{2d_{\Omega}}{\mathbb{T}_{k}(\mathfrak{N}_{k},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{k-1}(\mathfrak{N}_{k-1},\tau^{\prime\prime},\Delta^{0})}+1}{\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}}}\right)\right\rceil$
$\hskip 34.14322pt\left\lceil\log_{2}{\frac{\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{n}_{l}+1)}}+1}{\frac{d}{10}}}\right\rceil\Bigg\}$
$=$ $\mathfrak{n}_{k+1}.$
Hence, $m\geq\mathfrak{n}_{k+1}$ is sufficient for (3.27) to hold.
Step 3: Conclusion.
Finally, by induction, the estimate holds for $j=1,...,\operatorname{Conn}_{d}(x_{I},y^{0}_{i})$ and we conclude the proof of Proposition 3.5 by selecting $j=\operatorname{Conn}_{d}(x_{I},y^{0}_{i})$ and defining
$\mathfrak{R}(\tau^{\prime\prime}):=\max_{1\leq j\leq\operatorname{Conn}_{d}(\Omega)}\left\{\frac{2d_{\Omega}}{\mathbb{T}_{j}(\mathfrak{N}_{j},\tau^{\prime\prime},\Delta^{0})-\mathbb{T}_{j-1}(\mathfrak{N}_{j-1},\tau^{\prime\prime},\Delta^{0})}\right\},$
$\mathfrak{B}(\tau^{\prime\prime}):=\min_{1\leq j\leq\operatorname{Conn}_{d}(\Omega)}\{\mathbb{B}_{j}(\mathfrak{N}_{j},\tau^{\prime\prime},\Delta^{0},\alpha^{\prime}_{0},|v_{I}|)\},$
$\mathfrak{d}_{X}(\tau^{\prime\prime}):=\frac{\Delta^{0}}{2^{\sum_{i=1}^{\operatorname{Conn}_{d}(\Omega)}(\mathfrak{n}_{i}+1)}}\leq\frac{d}{10},$
$\mathfrak{d}_{T}(\tau^{\prime\prime}):=\min\bigg\{\frac{\Delta^{0}-\tau^{\prime\prime}}{2},\tau^{\prime\prime}-\mathbb{T}_{\operatorname{Conn}_{d}(\Omega)}(\mathfrak{N}_{\operatorname{Conn}_{d}(\Omega)},\tau^{\prime\prime},\Delta^{0}),$
$\hskip 71.13188pt\mathbb{U}_{\operatorname{Conn}_{d}(\Omega)}(\mathfrak{N}_{\operatorname{Conn}_{d}(\Omega)},\tau^{\prime\prime},\Delta^{0})-\tau^{\prime\prime},\mathbb{D}_{\operatorname{Conn}_{d}(\Omega)}(\mathfrak{N}_{\operatorname{Conn}_{d}(\Omega)},\tau^{\prime\prime},\Delta^{0})\bigg\},$
$\mathfrak{d}(\tau^{\prime\prime}):=\min\{\mathfrak{d}_{X}(\tau^{\prime\prime}),\mathfrak{d}_{T}(\tau^{\prime\prime})\}\leq\frac{d}{10},$
$v_{i}(\tau^{\prime\prime}):=\mathbb{V}_{\operatorname{Conn}_{d}(x_{I},y^{0}_{i})}(\mathfrak{N}_{\operatorname{Conn}_{d}(x_{I},y^{0}_{i})},\tau^{\prime\prime},\Delta^{0}).$
Remark 3.4.
Notice that for any $j\geq 1$, we have
(3.28) $\tau^{\prime\prime}>\mathbb{T}_{j+1}(\mathfrak{M}_{j+1},\tau^{\prime\prime},\Delta^{0})>\mathbb{T}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})\geq\frac{\tau^{\prime\prime}}{2}$
(with possibly equality when $j=1$ in the lower bound $\mathbb{T}_{j}\geq\frac{\tau^{\prime\prime}}{2}$)
(3.29) $0<\mathbb{D}_{j+1}(\mathfrak{M}_{j+1},\tau^{\prime\prime},\Delta^{0})<\mathbb{D}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),$
and
(3.30) $\tau^{\prime\prime}<\mathbb{U}_{j+1}(\mathfrak{M}_{j+1},\tau^{\prime\prime},\Delta^{0})\leq\mathbb{U}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0}),$
for $\mathfrak{m}_{1}\geq\mathfrak{n}_{1}$, $\mathfrak{m}_{i}\geq 0$ for $i\geq 2$.
Proof.
We prove it by induction on $j$. For $j=1$, assume $\mathfrak{m}_{1}\geq\mathfrak{n}_{1}$. We notice that
(3.31) $\frac{\tau^{\prime\prime}}{2}\leq\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})=\mathbb{T}(\tau^{\prime\prime},\Delta^{0},0,\mathfrak{m}_{1})=\max\left\{\frac{\tau^{\prime\prime}}{2},\tau^{\prime\prime}-\frac{\tau^{\prime\prime}\Delta^{0}}{2^{\mathfrak{m}_{1}+2}(4d_{\Omega}+\tau^{\prime\prime})}\right\}<\tau^{\prime\prime}$
and that
(3.32) $\begin{split}&\mathbb{T}_{2}((\mathfrak{m}_{1},\mathfrak{m}_{2}),\tau^{\prime\prime},\Delta^{0})=\mathbb{T}(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{\mathfrak{m}_{1}+1}},\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}),\mathfrak{m}_{2})\\ =&\max\left\{\frac{\tau^{\prime\prime}}{2}+\frac{\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})}{2},\tau^{\prime\prime}-\frac{(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}))\frac{\Delta^{0}}{2^{\mathfrak{m}_{1}+1}}}{2^{\mathfrak{m}_{2}+2}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})))}\right\}.\end{split}$
By (3.31), we have
$\frac{\tau^{\prime\prime}}{2}+\frac{\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})}{2}>\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}).$
Next, we notice that the following inequality
(3.33) $\tau^{\prime\prime}-\frac{(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}))\frac{\Delta^{0}}{2^{\mathfrak{m}_{1}+1}}}{2^{\mathfrak{m}_{2}+2}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})))}>\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})$
is equivalent to the inequality below:
(3.34) $\Delta^{0}<2^{\mathfrak{m}_{1}+\mathfrak{m}_{2}+3}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}))).$
We notice that by Lemma 2.1, we have $d_{\Omega}\geq d$. Indeed, we have
$x_{i}^{0}-2dn(x_{i}^{0}),\,x_{i}^{0}-\frac{1}{2}dn(x_{i}^{0})\in\Omega,\,\left|x_{i}^{0}-2dn(x_{i}^{0})-(x_{i}^{0}-\frac{1}{2}dn(x_{i}^{0}))\right|=\frac{3d}{2}.$
Notice that
(3.35) $2^{\mathfrak{m}_{1}}\geq\frac{\Delta^{0}+1}{\min\{d(x_{I},\partial\Omega),\frac{d}{10}\}}.$
Here we use the fact that
$\mathfrak{m}_{1}\geq\mathfrak{n}_{1}=\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\tau^{\prime\prime}}+|v_{I}|+1}{\Delta^{0}}\right)\right\rceil,\left\lceil\log_{2}{\frac{\Delta^{0}+1}{\min\{d(x_{I},\partial\Omega),\frac{d}{10}\}}}\right\rceil\right\}.$
As a result, we deduce that
(3.36) $\begin{split}&2^{\mathfrak{m}_{1}+\mathfrak{m}_{2}+3}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})))\\ \geq&\frac{\Delta^{0}+1}{\min\{d(x_{I},\partial\Omega),\frac{d}{10}\}}2^{\mathfrak{m}_{2}+3}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})))\\ \geq&\frac{32(\Delta^{0}+1)d}{\min\{d(x_{I},\partial\Omega),\frac{d}{10}\}}\\ >&\frac{d(\Delta^{0}+1)}{\frac{d}{10}}\\ >&\Delta^{0}.\end{split}$
Hence, we conclude that $\mathbb{T}_{2}((\mathfrak{m}_{1},\mathfrak{m}_{2}),\tau^{\prime\prime},\Delta^{0})>\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})$.
For the number $\mathbb{D}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})$, we have
$\mathbb{D}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})=\mathbb{D}(\tau^{\prime\prime},\Delta^{0},\mathfrak{m}_{1})=\frac{\tau^{\prime\prime}\Delta^{0}}{2^{\mathfrak{m}_{1}+2}(4d_{\Omega}+\tau^{\prime\prime})}>0.$
We also notice that
$\mathbb{D}_{2}((\mathfrak{m}_{1},\mathfrak{m}_{2}),\tau^{\prime\prime},\Delta^{0})=\mathbb{D}(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{\mathfrak{m}_{1}+1}},\mathfrak{m}_{2})$
$=$ $\frac{(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}))\frac{\Delta^{0}}{2^{\mathfrak{m}_{1}+1}}}{2^{\mathfrak{m}_{2}+2}(4d_{\Omega}+\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}))}$
$=$ $\frac{(\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}))\Delta^{0}}{2^{\mathfrak{m}_{1}+\mathfrak{m}_{2}+3}(4d_{\Omega}+\tau^{\prime\prime}-\mathbb{T}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}))}$
$<$ $\frac{\tau^{\prime\prime}\Delta^{0}}{2^{\mathfrak{m}_{1}+2}(4d_{\Omega}+\tau^{\prime\prime})}=\mathbb{D}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0}).$
For $\mathbb{U}_{j}(\mathfrak{M}_{j},\tau^{\prime\prime},\Delta^{0})$, we have
$\mathbb{U}_{1}:=\Delta^{0}>\tau^{\prime\prime},$
$\mathbb{U}_{2}((\mathfrak{m}_{1},\mathfrak{m}_{2}),\tau^{\prime\prime},\Delta^{0})=\min\{\mathbb{U}_{1},\tau^{\prime\prime}+\mathbb{D}_{1}(\mathfrak{m}_{1},\tau^{\prime\prime},\Delta^{0})\}\leq\mathbb{U}_{1}.$
So we finished the case $j=1$.
Suppose that (3.28), (3.29), (3.30) hold for $j=k$, we have
$\begin{split}&(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}))(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0})))\\ <&2^{\mathfrak{m}_{k+1}+1}(\tau^{\prime\prime}-\mathbb{T}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0}))(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}))),\end{split}$
from which we deduce that
$\begin{split}&\frac{(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}))\Delta^{0}}{2^{1+\sum_{l=1}^{k+1}(\mathfrak{m}_{l}+1)}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})))}\\ <&\frac{(\tau^{\prime\prime}-\mathbb{T}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0}))\Delta^{0}}{2^{1+\sum_{l=1}^{k}(\mathfrak{m}_{l}+1)}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0})))}.\end{split}$
Hence, we have
(3.37) $\begin{split}&\mathbb{T}_{k+1}(\mathfrak{M}_{k+1},\tau^{\prime\prime},\Delta^{0})\\ =&\mathbb{T}(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{m}_{l}+1)}},\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}),\mathfrak{m}_{k+1})\\ =&\max\begin{Bmatrix}\frac{\tau^{\prime\prime}}{2}+\frac{\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})}{2},\\ \tau^{\prime\prime}-\frac{(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}))\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{m}_{l}+1)}}}{2^{\mathfrak{m}_{k+1}+2}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})))}\end{Bmatrix}\\ =&\max\Bigg\{\frac{\tau^{\prime\prime}}{2}+\frac{\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})}{2},\\ &\hskip 34.14322pt\tau^{\prime\prime}-\frac{(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}))\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{m}_{l}+1)}}}{2^{\mathfrak{m}_{k+1}+2}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})))}\Bigg\}\\ =&\max\Bigg\{\frac{\tau^{\prime\prime}}{2}+\frac{\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})}{2},\\ &\hskip 34.14322pt\tau^{\prime\prime}-\frac{(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}))\Delta^{0}}{2^{1+\sum_{l=1}^{k+1}(\mathfrak{m}_{l}+1)}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})))}\Bigg\}\\ >&\max\Bigg\{\frac{\tau^{\prime\prime}}{2}+\frac{\mathbb{T}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0})}{2},\\ &\hskip 34.14322pt\tau^{\prime\prime}-\frac{(\tau^{\prime\prime}-\mathbb{T}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0}))\Delta^{0}}{2^{1+\sum_{l=1}^{k}(\mathfrak{m}_{l}+1)}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0})))}\Bigg\}\\ =&\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}).\end{split}$
Also, we notice that by induction hypothesis
(3.38) $\begin{split}&\mathbb{T}_{k+1}(\mathfrak{M}_{k+1},\tau^{\prime\prime},\Delta^{0})\\ =&\max\Bigg\{\frac{\tau^{\prime\prime}}{2}+\frac{\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})}{2},\\ &\hskip 34.14322pt\tau^{\prime\prime}-\frac{(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}))\Delta^{0}}{2^{1+\sum_{l=1}^{k+1}(\mathfrak{m}_{l}+1)}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})))}\Bigg\}\\ <&\tau^{\prime\prime}.\end{split}$
Next, we notice that
$0$ $<\mathbb{D}_{k+1}(\mathfrak{M}_{k+1},\tau^{\prime\prime},\Delta^{0})$
$=\mathbb{D}(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}),\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{m}_{l}+1)}},\mathfrak{m}_{k+1})$
$=\frac{(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}))\frac{\Delta^{0}}{2^{\sum_{l=1}^{k}(\mathfrak{m}_{l}+1)}}}{2^{\mathfrak{m}_{k+1}+2}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})))}$
$<\frac{(\tau^{\prime\prime}-\mathbb{T}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0}))\frac{\Delta^{0}}{2^{\sum_{l=1}^{k-1}(\mathfrak{m}_{l}+1)}}}{2^{\mathfrak{m}_{k}+2}(4d_{\Omega}+(\tau^{\prime\prime}-\mathbb{T}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0})))}$
$=\mathbb{D}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}),$
and that
$\tau^{\prime\prime}<$ $\mathbb{U}_{k+1}(\mathfrak{M}_{k+1},\tau^{\prime\prime},\Delta^{0})$
$=$ $\min\{\mathbb{U}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}),\tau^{\prime\prime}+\mathbb{D}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0})\}$
$<$ $\min\{\mathbb{U}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0}),\tau^{\prime\prime}+\mathbb{D}_{k-1}(\mathfrak{M}_{k-1},\tau^{\prime\prime},\Delta^{0})\}$
$=$ $\mathbb{U}_{k}(\mathfrak{M}_{k},\tau^{\prime\prime},\Delta^{0}).$
Hence, we show that (3.28), (3.29), (3.30) holds for $j=k+1$, and we conclude the proof of the statement of Remark 3.4.
Remark 3.5.
The numbers $\{\mathfrak{n}_{i}\}_{i=1}^{\operatorname{Conn}_{d}(\Omega)}$ and $\mathfrak{R}(\tau^{\prime\prime})$ do not depend on the choice of zigzag.
Proposition 3.6.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$, $\alpha\in[0,1]$. Let $f(t,x,v)$ be a continuous mild solution to (1.1)–(1.3). We consider $0<\tau^{\prime\prime}<\Delta^{0}$, where $\Delta^{0}$ is given in Proposition 3.1. Then, there exist $0<\delta_{X}(\tau^{\prime\prime})=\delta_{V}(\tau^{\prime\prime})\leq\frac{d}{56}$, $\delta_{T}(\tau^{\prime\prime})$, $a_{0}(\tau^{\prime\prime})>0$, which depend on $\Omega$, $M$, $E_{f}$, and a modulus of continuity of $f_{0}$ (and $L_{f,p}$ if $\gamma<0$); and $\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq 2$, which depends on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$), such that $\delta_{X}(\tau^{\prime\prime})\geq\delta_{T}(\tau^{\prime\prime})$ and that for all $N\in\mathbb{N}$, there exist $\mathcal{N}(N,\tau^{\prime\prime})$ in $\mathbb{N}$, $\{x_{i}\}_{i=1}^{\mathcal{N}(N,\tau^{\prime\prime})}$ in $\Omega$ and $\{\overline{v}_{i}(\tau^{\prime\prime})\}_{i=1}^{\mathcal{N}(N,\tau^{\prime\prime})}$ in $B(0,\mathfrak{R}_{\min}(\tau^{\prime\prime}))$, such that
(3.39) $\overline{\Omega}\subset\underset{1\leq i\leq\mathcal{N}(N,\tau^{\prime\prime})}{\bigcup}B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{N}}\right),$
(3.40) $f(t,x,v)\geq a_{0}(\tau^{\prime\prime})\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),\delta_{V}(\tau^{\prime\prime}))}(v)$
for any $1\leq i\leq\mathcal{N}(N,\tau^{\prime\prime})$, and any $(t,x,v)\in[\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T}(\tau^{\prime\prime})]\times[B(x_{i},\delta_{X}(\tau^{\prime\prime}))\cap\overline{\Omega}]\times\mathbb{R}^{3}$.
Proof.
Given $0<\tau^{\prime\prime}<\Delta^{0}$, we choose $0<d<\min\{1,\delta\}$ and notice that by Proposition 3.5 there exist $\mathfrak{B}(\tau^{\prime\prime})>0$, $\frac{d}{10}\geq\mathfrak{d}(\tau^{\prime\prime})>0$, $\mathfrak{R}(\tau^{\prime\prime})$, $\{v_{i}(\tau^{\prime\prime})\}_{i=1}^{m_{1}+m_{2}}\in B(0,\mathfrak{R}(\tau^{\prime\prime}))$ such that
(3.41) $f(t,x,v)\geq\ \mathfrak{B}(\tau^{\prime\prime})\mathbf{1}_{B(y^{0}_{i},\mathfrak{d}(\tau^{\prime\prime}))\times B(v_{i}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}))}(v),\ \forall t\in[\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})],$
for any $1\leq i\leq m_{1}+m_{2}$.
Notice that, by the definition and Remark 3.4, we have
(3.42) $\begin{split}\tau^{\prime\prime}&>\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})\geq\tau^{\prime\prime}-(\tau^{\prime\prime}-\mathbb{T}_{\operatorname{Conn}_{d}(\Omega)}(\mathfrak{N}_{\operatorname{Conn}_{d}(\Omega)},\tau^{\prime\prime},\Delta^{0}))\\ &=\mathbb{T}_{\operatorname{Conn}_{d}(\Omega)}(\mathfrak{N}_{\operatorname{Conn}_{d}(\Omega)},\tau^{\prime\prime},\Delta^{0})>\frac{\tau^{\prime\prime}}{2},\end{split}$
(3.43) $\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})\leq\tau^{\prime\prime}+\frac{\Delta^{0}-\tau^{\prime\prime}}{2}=\frac{\Delta^{0}+\tau^{\prime\prime}}{2}<\Delta^{0}.$
Now, we define
(3.44) $\mathfrak{n}^{\prime\prime}:=\max\left\{\mathbb{M}\left(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}),\frac{d}{16}\right),\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\mathfrak{d}(\tau^{\prime\prime})}+\mathfrak{R}(\tau^{\prime\prime})+1}{\mathfrak{d}(\tau^{\prime\prime})}\right)\right\rceil\right\}.$
Notice that
$r_{\mathfrak{n}^{\prime\prime}}(\mathfrak{d}(\tau^{\prime\prime}))=\left(\frac{3\sqrt{2}}{4}\right)^{\mathfrak{n}^{\prime\prime}}\mathfrak{d}(\tau^{\prime\prime})\geq\frac{4d_{\Omega}}{\mathfrak{d}(\tau^{\prime\prime})}+\mathfrak{R}(\tau^{\prime\prime})+1.$
We also define
(3.45) $\begin{split}\delta_{X}(\tau^{\prime\prime})=\delta_{V}(\tau^{\prime\prime}):=\min\left\{\frac{1}{4}\min\limits_{1\leq i\leq m_{1}+m_{2}}d(y_{i}^{0},\partial\Omega),\frac{d}{56},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+2}},\frac{2d_{\Omega}}{\mathfrak{d}(\tau^{\prime\prime})}\right\}.\end{split}$
Given $N\in\mathbb{N}$, we notice that by compactness of $\overline{\Omega}$, there exists $\{x_{i}\}_{i=1}^{\mathcal{N}(N,\tau^{\prime\prime})}\in\Omega$ such that $\overline{\Omega}\subset\underset{1\leq i\leq\mathcal{N}(N,\tau^{\prime\prime})}{\bigcup}B(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{N}})$. Given $1\leq i\leq\mathcal{N}(N,\tau^{\prime\prime})$, we consider two cases.
1. Point near the boundary: If $d(x_{i},\partial\Omega)\leq\frac{d}{7}$, then there exists $x^{\prime}_{i}\in\partial\Omega$ such that $|x^{\prime}_{i}-x_{i}|\leq\frac{d}{6}$. Also, by (2.1), there exists $k(i)\in\{1,2,...,m_{1}\}$ such that $|x_{k(i)}^{0}-x^{\prime}_{i}|\leq\frac{d}{8}$, so that $|x_{i}-x_{k(i)}^{0}|\leq|x_{i}-x_{i}^{\prime}|+|x_{i}^{\prime}-x_{k(i)}^{0}|\leq\frac{7}{24}d$.
We consider
(3.46) $T^{\prime\prime}:=\max\left\{\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{n}^{\prime\prime}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\},$
$\overline{v}_{i}(\tau^{\prime\prime}):=\frac{2(x_{i}-y^{0}_{k(i)})}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}=\frac{2(x_{i}-y^{0}_{k(i)})}{\max\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime})-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{n}^{\prime\prime}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}}.$
$\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime}):=\frac{2d_{\Omega}}{\max\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime})-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{n}^{\prime\prime}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}}$
Here, we recall the definition of $y^{0}_{j}$ in Lemma 2.1. As long as $x\in B(x_{i},\min\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}},\frac{d}{56}\})\cap\Omega$, $v\in B\left(\overline{v}_{i}(\tau^{\prime\prime}),\min\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}(t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\right\}\right)$, $T^{\prime\prime}\leq t\leq T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$, we have
(3.47) $x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v\in B\left(y^{0}_{k(i)},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}}}\right).$
Furthermore, by definition of $\mathfrak{n}^{\prime\prime}$ in (3.44), we have
(3.48) $r_{\mathfrak{n}^{\prime\prime}}(\mathfrak{d}(\tau^{\prime\prime}))\geq\mathfrak{R}(\tau^{\prime\prime})+\frac{4d_{\Omega}}{\mathfrak{d}(\tau^{\prime\prime})}+1\geq|v_{k(i)}(\tau^{\prime\prime})|+\left|\frac{2(x_{i}-y^{0}_{k(i)})}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}\right|+1\geq|v_{k(i)}(\tau^{\prime\prime})|+|v|,$
which implies
(3.49) $v\in B(v_{k(i)}(\tau^{\prime\prime}),r_{\mathfrak{n}^{\prime\prime}}(\mathfrak{d}(\tau^{\prime\prime}))).$
Here, $\mathfrak{R}(\tau^{\prime\prime})$ is defined in Proposition 3.5. Next, notice that
(3.50) $|x-x^{0}_{k(i)}|\leq|x-x_{i}|+|x_{i}-x^{0}_{k(i)}|\leq\frac{d}{56}+\frac{7d}{24}<\frac{d}{3},$
so we have $x\in B(x^{0}_{k(i)},\frac{d}{3})$. Now, we parametrize $x$ and $x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v$ as below:
[…]
We notice by (3.50) that $\widetilde{x}_{3}<\frac{d}{3}$.
Since we have (notice that $\mathfrak{d}(\tau^{\prime\prime})\leq\frac{d}{8}$)
$x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v\in B(y^{0}_{k(i)},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}}})\subset B(y^{0}_{k(i)},\frac{d}{8}),$
we deduce that $\overline{x}_{3}>\frac{3d}{4}-\frac{d}{8}$.
As a result, we have
$|x-[x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v]|$
$\leq$ $|x-x^{0}_{k(i)}|+|x^{0}_{k(i)}-y^{0}_{k(i)}|+|y^{0}_{k(i)}-[x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v]|$
$\leq$ $\frac{d}{3}+\frac{3d}{4}+\frac{d}{8}=\frac{29d}{24},$
which implies that
(3.53) $\begin{split}&n(x^{0}_{k(i)})\cdot\frac{v}{|v|}\\ =&n(x^{0}_{k(i)})\cdot\frac{x-[x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v]}{|x-[x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v]|}\\ \geq&n(x^{0}_{k(i)})\cdot\frac{[(\widetilde{x}_{1}-\overline{x}_{1})e^{1}_{k(i)}+(\widetilde{x}_{2}-\overline{x}_{2})e^{2}_{k(i)}-(\widetilde{x}_{3}-\overline{x}_{3})n(x^{0}_{k(i)})]}{\frac{29d}{24}}\\ =&\frac{\overline{x}_{3}-\widetilde{x}_{3}}{\frac{29d}{24}}\\ \geq&\frac{\frac{3d}{4}-\frac{d}{8}-\frac{d}{3}}{\frac{29d}{24}}=\frac{7}{29}.\end{split}$
Hence, by using the method as before, we have
(3.54) $\left\{\,x-sv\middle|\ 0\leq s\leq\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}\,\right\}\subset\Omega.$
We also notice that
(3.55) $\begin{split}&T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}\\ \geq&\max\left\{\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{n}^{\prime\prime}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}+\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{n}^{\prime\prime}+1}\left(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime})\right)}\\ \geq&\tau^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{n}^{\prime\prime}+2}\left(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime})\right)}.\end{split}$
Since $B(y^{0}_{k(i)},\mathfrak{d}(\tau^{\prime\prime}))\subset B(y^{0}_{k(i)},\frac{d}{8})\subset\Omega$ by Lemma 2.2, from (4.2), we can apply Proposition 3.2 with $\tau=\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})$, $\Delta_{1}=2\mathfrak{d}(\tau^{\prime\prime}),\ \Delta_{2}=\mathfrak{d}(\tau^{\prime\prime})$, $x^{\prime}=y^{0}_{k(i)}$, and $v^{\prime}=v_{k(i)}(\tau^{\prime\prime})$.
Hence, we have for $\tau^{\prime\prime}\leq t\leq T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$ (see Remark 3.2 and the fact that $T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}\leq\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})$)
(3.56) $\begin{split}&f\left(\frac{t+(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2},x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)\\ \geq&\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)\\ &\mathbf{1}_{B\left(y^{0}_{k(i)},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}}}\right)\times B(v_{k(i)}(\tau^{\prime\prime}),r_{\mathfrak{n}^{\prime\prime}}(\mathfrak{d}(\tau^{\prime\prime})))}\left(x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right),\end{split}$
which implies by the Duhamel formula that for any $x\in B(x_{i},\min\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}},\frac{d}{56}\})\cap\Omega$, $v\in B(\overline{v}_{i}(\tau^{\prime\prime}),\min\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}(t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\})$, $\tau^{\prime\prime}\leq t\leq T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$, we have
(3.57) $\begin{split}&f(t,x,v)\\ \geq&f\left(\frac{t+(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2},x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)\\ &\mathbf{1}_{B\left(y^{0}_{k(i)},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}}}\right)\times B(v_{k(i)}(\tau^{\prime\prime}),r_{\mathfrak{n}^{\prime\prime}}(\mathfrak{d}(\tau^{\prime\prime})))}\left(x-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}.\end{split}$
Here, the last line is from (3.47) and (3.49). Now, we define
$\delta_{T,1}(\tau^{\prime\prime}):=\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{n}^{\prime\prime}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}.$
Notice that $\delta_{T,1}(\tau^{\prime\prime})\leq\mathfrak{d}(\tau^{\prime\prime})$, and we see by (3.55) that
$\tau^{\prime\prime}+\delta_{T,1}(\tau^{\prime\prime})\leq T^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}\left(\frac{2d_{\Omega}}{T^{\prime\prime}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}.$
As a result, we deduce from (3.57) that for all $(x,v)\in\Omega\times\mathbb{R}^{3}$ and $\tau^{\prime\prime}\leq t\leq\tau^{\prime\prime}+\delta_{T,1}(\tau^{\prime\prime})$, we have
$f(t,x,v)\geq\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}$
$\mathbf{1}_{B(x_{i},\min\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}},\frac{d}{56}\})\times B(\overline{v}_{i}(\tau^{\prime\prime}),\min\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}(t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\})}(x,v)$
$\geq\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}$
$\hskip 14.22636pt\mathbf{1}_{B(x_{i},\delta_{X})\times B(\overline{v}_{i}(\tau^{\prime\prime}),\delta_{V}(\tau^{\prime\prime}))}(x,v),$
where we used (3.45) to derive that $\delta_{X}\leq\max\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}},\frac{d}{56}\}$. In the final line we used the fact that
$\min\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}(t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\right\}$
$\geq$ $\min\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}((\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime}))-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\right\}\geq\min\left\{\frac{1}{2^{\mathfrak{n}^{\prime\prime}+2}},1\right\}.$
2. Point far from the boundary.
If $d(x_{i},\partial\Omega)>\frac{d}{7}$, by Lemma 2.2, we can find $m_{1}+1\leq k(i)\leq m_{1}+m_{2}$ such that $x_{i}\in B(y^{0}_{k(i)},\frac{d}{16})$, which implies $\overline{y^{0}_{k(i)}x_{i}}\subset B(y^{0}_{k(i)},\frac{d}{16})\subset\Omega-\Omega_{\frac{d}{16}}$.
We can use Corollary 3.4, since by (4.2), (3.22) is satisfied with $\tau=\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})$, $\Delta_{1}=\Delta_{1}^{\prime}=\Delta_{2}=\mathfrak{d}(\tau^{\prime\prime})$, $x^{\prime}=y^{0}_{k(i)}$, $y=x_{i}$, $v^{\prime}=v_{k(i)}(\tau^{\prime\prime})$. Notice as before that all the assumptions of Corollary 3.4 are satisfied by the fact that $B(y^{0}_{k(i)},\mathfrak{d}(\tau^{\prime\prime}))\subset B(y^{0}_{k(i)},\frac{d}{8})\subset\Omega$ and that $\overline{y^{0}_{k(i)}x_{i}}\subset B\left(y^{0}_{k(i)},\frac{d}{16}\right)\subset\Omega$.
Thus, we have
(3.58) $\begin{split}&f(t,x,v)\\ &\geq\mathbb{B}(\widetilde{m},\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),|v_{k(i)}(\tau^{\prime\prime})|)\\ &\hskip 11.38092pt\times\mathbf{1}_{B(x_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\widetilde{m}+1}})\times B\left(\mathbb{V}(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\widetilde{m},x_{i},y^{0}_{k(i)}),\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\widetilde{m}+1}}\right)}(x,v).\end{split}$
[…]
(3.59) $\mathfrak{n}^{\prime\prime}=\mathbb{M}\left(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}),\frac{d}{16}\right)\geq\mathbb{M}\left(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),|v_{k(i)}(\tau^{\prime\prime})|,d\left(\overline{y^{0}_{k(i)}x_{i}},\partial\Omega\right)\right).$
As a result, by (3.58), we have for $(x,v)\in\Omega\times\mathbb{R}^{3}$ and $\mathbb{T}(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{n}^{\prime\prime})\leq t\leq\min\{\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}+\mathbb{D}(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{n}^{\prime\prime}))\}=\tau^{\prime\prime}+\delta_{T,2}(\tau^{\prime\prime})$
$f(t,x,v)$
$\geq$ $\mathbb{B}(\mathfrak{n}^{\prime\prime},\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),|v_{k(i)}(\tau^{\prime\prime})|)$
$\mathbf{1}_{B(x_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}})\times B\left(\mathbb{V}(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{n}^{\prime\prime},x_{i},y^{0}_{k(i)}),\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{n}^{\prime\prime}+1}}\right)}(x,v)$
$\geq$ $\mathbb{B}(\mathfrak{n}^{\prime\prime},\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),|v_{k(i)}(\tau^{\prime\prime})|)$
$\mathbf{1}_{B(x_{i},\delta_{X})\times B\left(\mathbb{V}(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{n}^{\prime\prime},x_{i},y^{0}_{k(i)}),\delta_{V}(\tau^{\prime\prime})\right)}(x,v)$
$\geq$ $\mathbb{B}(\mathfrak{n}^{\prime\prime},\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))$
$\mathbf{1}_{B(x_{i},\delta_{X})\times B\left(\mathbb{V}(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{n}^{\prime\prime},x_{i},y^{0}_{k(i)}),\delta_{V}(\tau^{\prime\prime})\right)}(x,v),$
where we used (3.45). We now define
(3.60) $\overline{v}_{i}(\tau^{\prime\prime}):=\mathbb{V}(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{n}^{\prime\prime},x_{i},y^{0}_{k(i)})=\frac{2(x_{i}-y^{0}_{k(i)})}{\max\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime})-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{n}^{\prime\prime}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}},$
$\mathfrak{R}^{2}_{\min}(\tau^{\prime\prime}):=\frac{2d_{\Omega}}{\max\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime})-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{n}^{\prime\prime}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}}$
and finish the case by noticing that
$\mathbb{T}(\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{n}^{\prime\prime})<\tau^{\prime\prime}$
and that
$|\overline{v}_{i}(\tau^{\prime\prime})|\leq\mathfrak{R}^{2}_{\min}(\tau^{\prime\prime}).$
We conclude the proof by defining
(3.61) $\delta_{T}(\tau^{\prime\prime}):=\min\{\delta_{T,1}(\tau^{\prime\prime}),\delta_{T,2}(\tau^{\prime\prime}),\delta_{X}(\tau^{\prime\prime})\},$
(3.62) $\mathfrak{R}_{\min}(\tau^{\prime\prime}):=\max\{\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime}),\mathfrak{R}^{2}_{\min}(\tau^{\prime\prime}),2\},$
and
$\begin{split}&a_{0}(\tau^{\prime\prime}):=\min\bigg\{\alpha_{\mathfrak{n}^{\prime\prime}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)e^{-C_{L}\left\langle\mathfrak{R}^{1}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}},\\ &\hskip 71.13188pt\mathbb{B}(\mathfrak{n}^{\prime\prime},\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))\bigg\}.\end{split}$
4. Maxwellian bound for the cut-off case for non-fully specular reflection condition
In this section, we consider the lower bound problem in the cut-off case with a non-fully specular boundary condition.
Proposition 4.1.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\alpha\in[0,1]$. We consider a continuous mild solution $f(t,x,v)$ of (1.1)–(1.3) . Given $\tau^{\prime\prime}\in(0,\Delta^{0})$ with $\Delta^{0}>0$ given in Proposition 3.1, there exist $b(\tau^{\prime\prime})$, $\delta_{T,I}(\tau^{\prime\prime})>0$, which depends on $\Omega,M,E_{f}$, and the modulus of continuity of $f_{0}$, such that for $t\in[\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime})]$,
(4.1) $\forall x\in\partial\Omega,\ \int_{v_{*}\cdot n(x)>0}f(t,x,v_{*})(v_{*}\cdot n(x))\,dv_{*}>b(\tau^{\prime\prime}).$
Remark 4.1.
Although the estimate remains valid for $\alpha=1$, in our analysis, we only employ it for $\alpha\in[0,1)$.
Proof.
Given $0<\tau^{\prime\prime}<\Delta^{0}$, we choose $0<d<\min\{1,\delta\}$ and notice that by Proposition 3.5 there exist $\mathfrak{B}(\tau^{\prime\prime})>0$, $\frac{d}{10}\geq\mathfrak{d}(\tau^{\prime\prime})>0$, $\mathfrak{R}(\tau^{\prime\prime})$, $\{v_{i}(\tau^{\prime\prime})\}_{i=1}^{m_{1}+m_{2}}\in B(0,\mathfrak{R}(\tau^{\prime\prime}))$ such that
(4.2) $f(t,x,v)\geq\ \mathfrak{B}(\tau^{\prime\prime})\mathbf{1}_{B(y^{0}_{i},\mathfrak{d}(\tau^{\prime\prime}))\times B(v_{i}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}))}(x,v),\ \forall t\in[\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})],$
for any $1\leq i\leq m_{1}+m_{2}$. Given $x_{\partial}\in\partial\Omega$, by Lemma 2.1, there exists $1\leq i\leq m_{1}$ such that $x_{\partial}\in B(x^{0}_{i},\frac{d}{8})$.
Then, we proceed with the same method as in Proposition 3.6, this time we propagate the lower bound from $x^{0}_{i}$ to $x_{\partial}$
We define
(4.3) $\mathfrak{m}_{\partial}:=\max\left\{1,\left\lceil\log_{\frac{3\sqrt{2}}{4}}\left(\frac{\frac{4d_{\Omega}}{\mathfrak{d}(\tau^{\prime\prime})}+\mathfrak{R}(\tau^{\prime\prime})+1}{\mathfrak{d}(\tau^{\prime\prime})}\right)\right\rceil\right\},$
and define
(4.4) $T_{\partial}:=\max\left\{\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\},$
(4.5) $v_{x_{\partial}}(\tau^{\prime\prime}):=\frac{2(x_{\partial}-y^{0}_{i})}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}=\frac{2(x_{\partial}-y^{0}_{i})}{\max\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime})-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}},$
(4.6) $R^{\partial}_{\min}(\tau^{\prime\prime}):=\frac{2d_{\Omega}}{\max\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime})-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}}.$
Notice that, by the definition of notation of $R^{\partial}_{\min}(\tau^{\prime\prime})$, we have
(4.7) $|v_{x_{\partial}}(\tau^{\prime\prime})|\leq R^{\partial}_{\min}(\tau^{\prime\prime}).$
As long as $v\in B\left(v_{x_{\partial}}(\tau^{\prime\prime}),\min\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}(t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\right\}\right)$, $T_{\partial}\leq t\leq T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$, we have
(4.8) $\begin{split}&\left|x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v-y^{0}_{i}\right|<\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}}.\end{split}$
Furthermore, by the definition of $\mathfrak{m}_{\partial}$ in (4.3), we have
(4.9) $r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime}))\geq\mathfrak{R}(\tau^{\prime\prime})+\frac{4d_{\Omega}}{\mathfrak{d}(\tau^{\prime\prime})}+1\geq|v_{i}(\tau^{\prime\prime})|+\left|\frac{2(x_{\partial}-y^{0}_{i})}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}\right|+1\geq|v_{i}(\tau^{\prime\prime})|+|v|,$
which implies $v\in B(v_{i}(\tau^{\prime\prime}),r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime})))$. Here, we recall $\mathfrak{R}(\tau^{\prime\prime})$ again in Proposition 3.5. Next, notice that
(4.10) $|x_{\partial}-x^{0}_{i}|\leq\frac{d}{8},$
and that
(4.11) $x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v\in B(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}})\subset B(y^{0}_{i},\mathfrak{d}(\tau^{\prime\prime}))\subset B(y^{0}_{i},\frac{d}{8})$
by the fact that $\mathfrak{d}(\tau^{\prime\prime})\leq\frac{d}{8}$. As before, we note that the segment $\left\{x_{\partial}-sv|\ 0\leq s\leq\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}\right\}$ between $x_{\partial}$ and $x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v\in B\left(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}}}\right)$ also lies within $\Omega$. We also notice that
(4.12) $\begin{split}&T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}\\ \geq&\max\left\{\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}+\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+1}\left(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime})\right)}\\ \geq&\tau^{\prime\prime}+\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+2}\left(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime})\right)}.\end{split}$
Then, using Proposition 3.2 (note that as before by (4.2), the assumption (3.9) holds with $\tau=\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})$, $\Delta_{1}=2\mathfrak{d}(\tau^{\prime\prime}),\ \Delta_{2}=\mathfrak{d}(\tau^{\prime\prime})$, $x^{\prime}=y^{0}_{i}$, and $v^{\prime}=v_{i}(\tau^{\prime\prime})$, we also have $B(y^{0}_{i},\mathfrak{d}(\tau^{\prime\prime}))\subset B(y^{0}_{i},\frac{d}{8})\subset\Omega$ by Lemma 2.2), we have for $\tau^{\prime\prime}\leq t\leq T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$ (see Remark 3.2 and the fact that $T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}\leq\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})$)
(4.13) $\begin{split}&f\left(\frac{t+(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2},x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)\\ \geq&\alpha_{\mathfrak{m}_{\partial}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)\\ &\mathbf{1}_{B\left(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}}}\right)\times B\left(v_{i}(\tau^{\prime\prime}),r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime}))\right)}\left(x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right),\end{split}$
which implies by the Duhamel formula that for any $v\in B(v_{x_{\partial}}(\tau^{\prime\prime}),\min\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}(t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\})$, $\tau^{\prime\prime}\leq t\leq T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$, we have (notice that by (4.7), we have $|v|\leq|v_{x_{\partial}}(\tau^{\prime\prime})|+1\leq R^{\partial}_{\min}(\tau^{\prime\prime})+1$)
(4.14) $\begin{split}&f(t,x_{\partial},v)\\ \geq&f\left(\frac{t+(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2},x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{m}_{\partial}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)\\ &\mathbf{1}_{B\left(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}}}\right)\times B(v_{i}(\tau^{\prime\prime}),r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime})))}\left(x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}.\end{split}$
Next, we recall (4.8) and (4.9):
(4.15) $x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v\in B\left(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}}}\right),v\in B(v_{i}(\tau^{\prime\prime}),r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime}))).$
So, we have
(4.16) $\begin{split}&\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))\\ &\mathbf{1}_{B\left(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}}}\right)\times B(v_{i}(\tau^{\prime\prime}),r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime})))}(x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v)e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}.\end{split}$
Now, we define $\delta_{T,I}(\tau^{\prime\prime}):=\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}$(notice as before that $\delta_{T,I}(\tau^{\prime\prime})\leq\mathfrak{d}(\tau^{\prime\prime})$), and notice by (4.12) that
(4.17) $T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}\geq\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime}).$
As a result, we have by (4.14) that for all $(x,v)\in\Omega\times\mathbb{R}^{3}$ and $\tau^{\prime\prime}\leq t\leq\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime})$ ,
$f(t,x_{\partial},v)$
$\geq$ $\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}$
$\mathbf{1}_{B(v_{x_{\partial}}(\tau^{\prime\prime}),\min\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}(t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\})}(v)$
$\geq$ $\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\mathbf{1}_{B(v_{x_{\partial}}(\tau^{\prime\prime}),\frac{1}{2^{\mathfrak{m}_{\partial}+2}})}(v).$
In the final line, we use the fact again that
$\min\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}(t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\right\}\geq\min\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}((\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime}))-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})))},1\right\}\geq\frac{1}{2^{\mathfrak{m}_{\partial}+2}}.$
Next, we define
$A^{\prime}:=\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}},$
$\delta_{V,I}(\tau^{\prime\prime}):=\frac{1}{2^{\mathfrak{m}_{\partial}+2}}.$
We conclude that
[…]
for $\tau^{\prime\prime}\leq t\leq\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime})$. Here we recall as in (4.5) that $v_{x_{\partial}}(\tau^{\prime\prime})=\frac{2(x_{\partial}-y^{0}_{i})}{\max\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime})-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}}$.
Next, by parameterizing
$x_{\partial}=x_{i}^{0}+x_{\partial,1}e^{1}_{i}+x_{\partial,2}e^{2}_{i}-\phi_{i}(x_{\partial,1},x_{\partial,2})n(x_{i}^{0}),$
we find that
(4.19) $\begin{split}&v_{x_{\partial}}(\tau^{\prime\prime})\cdot n(x_{\partial})\\ =&\frac{2}{\max\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime})-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}}[(x_{i}^{0}-y_{i}^{0})+(x_{\partial}-x_{i}^{0})]\cdot n(x_{\partial})\\ \geq&\frac{2}{\max\left\{\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime})-\frac{\mathfrak{d}(\tau^{\prime\prime})^{2}}{2^{\mathfrak{m}_{\partial}+2}(4d_{\Omega}+\mathfrak{d}(\tau^{\prime\prime}))}\right\}\sqrt{|\nabla\phi_{i}(x_{\partial,1},x_{\partial,2})|^{2}+1}}\\ &\times\left[\left(0,0,-\frac{3d}{4}\right)\cdot(\partial_{x}\phi_{i}(x_{\partial,1},x_{\partial,2}),\partial_{y}\phi_{i}(x_{\partial,1},x_{\partial,2}),-1)-\left|(x_{\partial}-x_{i}^{0})\right|\right]\\ \geq&\frac{5d}{8\mathfrak{d}(\tau^{\prime\prime})}.\end{split}$
Then, we have
$v\cdot n(x_{\partial})\geq\frac{5d}{16\mathfrak{d}(\tau^{\prime\prime})},\ \forall v\in B\left(v_{x_{\partial}}(\tau^{\prime\prime}),\frac{5d}{16\mathfrak{d}(\tau^{\prime\prime})}\right).$
Hence, for $t\in[\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime})]$, $x_{\partial}\in\partial\Omega$,
(4.20) $\begin{split}&\int_{v_{*}\cdot n(x_{\partial})>0}f(t,x_{\partial},v_{*})(v_{*}\cdot n(x_{\partial}))\,dv_{*}\\ \geq&\int_{B\left(v_{x_{\partial}}(\tau^{\prime\prime}),\min\{\delta_{V,I}(\tau^{\prime\prime}),\frac{5d}{16\mathfrak{d}(\tau^{\prime\prime})}\}\right)}A^{\prime}\frac{5d}{16\mathfrak{d}(\tau^{\prime\prime})}\,dv_{*}\\ \geq&A^{\prime}\frac{5d}{16\mathfrak{d}(\tau^{\prime\prime})}\frac{4\pi(\min\{\delta_{V,I}(\tau^{\prime\prime}),\frac{5d}{16\mathfrak{d}(\tau^{\prime\prime})}\})^{3}}{3}\\ =:&b(\tau^{\prime\prime}),\end{split}$
and we finish the proof.
The rest of the proof of existence of lower bound is exactly the same as in Briant’s paper [1](from Prop. 3.9, pp. 21–25). Before we introduce the next proposition, we define some constants here:
Definition 4.1.
Given $\tau^{\prime\prime},\tau^{\prime}>0$, $N\in\mathbb{N}$, $0\leq\alpha<1$, $\delta_{X}(\tau^{\prime\prime})$, $\delta_{V}(\tau^{\prime\prime})$, $\mathfrak{R}_{\min}(\tau^{\prime\prime})$ as in Prop. 3.6 and a sequence $\{\xi_{n}\}_{n=1}^{\infty}$ satisfying $0\leq\xi_{n}\leq\frac{1}{4}$ for $n\in\mathbb{N}$, we define the following numbers:
(4.21) $r^{\prime\prime}_{n}:=\begin{cases}\delta_{V}(\tau^{\prime\prime}),\ n=0\\ \sqrt{2}(1-\xi_{n})r^{\prime\prime}_{n-1},\ n\geq 1,\end{cases}$
(4.22) $b_{n,\tau^{\prime\prime}}(\tau^{\prime}):=\begin{cases}\min\{b(\tau^{\prime\prime}),a_{0}(\tau^{\prime\prime})\}&,\ n=0\\ \min\begin{Bmatrix}(1-\alpha)b(\tau^{\prime\prime})e^{-C_{L}(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1}\rangle^{\gamma^{+}}}\\ \times\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1})^{2}}{2T_{B}}},\\ (\min(a_{0}(\tau^{\prime\prime}),b(\tau^{\prime})))^{2}C_{Q}r_{0}^{{}^{\prime\prime}3+\gamma}\xi^{\frac{1}{2}}_{1}\\ \times\frac{\tau^{\prime}-\tau^{\prime\prime}}{2^{2}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1})}e^{-C_{L}\frac{(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1}\rangle^{\gamma^{+}}}{2(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1})}}\end{Bmatrix}&,\ n=1\\ (1-\alpha)b(\tau^{\prime\prime})e^{-C_{L}(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n}\rangle^{\gamma^{+}}}\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})^{2}}{2T_{B}}}&,\ n\geq 2,\end{cases}$
where $b(\tau^{\prime\prime})$ is introduced in Prop. 4.1, and
(4.23) $a_{n,\tau^{\prime\prime}}(\tau^{\prime}):=\begin{cases}a_{0}(\tau^{\prime\prime})&,\ n=0\\ (\min(a_{n-1,\tau^{\prime\prime}}(\tau^{\prime}),b_{n-1,\tau^{\prime\prime}}(\tau^{\prime})))^{2}C_{Q}r_{n-1}^{{}^{\prime\prime}3+\gamma}\xi^{\frac{1}{2}}_{n}\\ \times\frac{\tau^{\prime}-\tau^{\prime\prime}}{2^{n+1}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})}e^{-C_{L}\frac{(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n}\rangle^{\gamma^{+}}}{2^{n}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})}}&,\ n\geq 1,\end{cases}$
where $a_{0}(\tau^{\prime\prime})$ is defined in Proposition 3.6.
We now introduce the following proposition.
Proposition 4.2.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$. Let $f$ be the continuous mild solution of (1.1)–(1.3) with $0\leq\alpha<1$. We fix $\tau^{\prime\prime}\in(0,\Delta^{0})$, where $\Delta^{0}>0$ is given in Proposition 3.1. Then, there exist $\delta_{T,D}(\tau^{\prime\prime})>0$, which depends on $\Omega,M,E_{f}$, and the modulus of continuity of $f_{0}$, such that for every $\tau^{\prime}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})]$ and every $N\in\mathbb{N}$, for each $0\leq n\leq N$ and each $1\leq i\leq\mathcal{N}(N,\tau^{\prime\prime})$, we have
(4.24) $\begin{split}&\ \ \ \ \ f(t,x,v)\geq\min\left(a_{n,\tau^{\prime\prime}}(\tau^{\prime}),b_{n,\tau^{\prime\prime}}(\tau^{\prime})\right)\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{\prime\prime}_{n})}(v),\\ &\forall(t,x)\in\left[\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{n+1}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})}\right),\tau^{\prime}\right]\times\left[B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{n}}\right)\cap\overline{\Omega}\right],\end{split}$
where $x_{i}$ and $\overline{v}_{i}(\tau^{\prime\prime})$ are defined in Proposition 3.6 and depend on $N$. Here, we recall that the notation $\delta_{X}(\tau^{\prime\prime})$, $\mathfrak{R}_{\min}(\tau^{\prime\prime})$, $\mathcal{N}(N,\tau^{\prime\prime})$ are defined in Proposition 3.6 and that $a_{n,\tau^{\prime\prime}}(\tau^{\prime})>0,\,b_{n,\tau^{\prime\prime}}(\tau^{\prime})>0$ and $r^{\prime\prime}_{n}$ are defined in Definition 4.1.
Proof.
The proof follows the same induction-on-$n$ scheme as Proposition 3.9 of [1]. Notice that we have $\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq 1$ in Proposition 3.6. Then, we define (we recall that the notion $\delta_{T,I}(\tau^{\prime\prime})$ provided in Proposition 4.1)
(4.25) $\delta_{T,D}(\tau^{\prime\prime}):=\min\{\delta_{X}(\tau^{\prime\prime}),\delta_{T}(\tau^{\prime\prime}),\delta_{T,I}(\tau^{\prime\prime})\}$
and pick $\tau^{\prime}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})]$. This choice ensures $\delta_{T,D}(\tau^{\prime\prime})\leq\delta_{X}(\tau^{\prime\prime})$ (for the interior displacement), $\delta_{T,D}(\tau^{\prime\prime})\leq\delta_{T}(\tau^{\prime\prime})$ (for the boundary integral) and $\delta_{T,D}(\tau^{\prime\prime})\leq\delta_{T,I}(\tau^{\prime\prime})$(for the time slab). The proof uses an induction on $n$. The case $n=0$ is a direct consequence of Proposition 3.6.
We now assume that the case $n=k$ holds. Given $1\leq i\leq\mathcal{N}(N,\tau^{\prime\prime})$, consider
(4.26) $\begin{split}&(t,x,v)\in\left[\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+2}(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime}))}\right),\tau^{\prime}\right]\\ &\hskip 56.9055pt\times\left[B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}}\right)\cap\overline{\Omega}\right]\times B\left(\overline{v}_{i}(\tau^{\prime\prime}),r^{\prime\prime}_{k+1}\right).\end{split}$
We generate the announced lower bound depending on whether the characteristic line $\{X_{s,t}(x,v)\}_{\tau^{\prime\prime}\leq s\leq t}$ touches the boundary or not.
First case: $X_{s,t}(x,v)\in\overline{\Omega}$ for $\tau^{\prime\prime}\leq s\leq t$. We use the Duhamel formula (1.12) and apply Lemma 3.1 to deduce that
(4.27) $\begin{split}&f(t,x,v)\\ \geq&\int_{\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+1}\left(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}\right)}^{\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+2}\left(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}\right)}e^{-C_{L}(t-s)\left\langle r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right\rangle^{\gamma^{+}}}\\ &Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds.\end{split}$
Next, we notice that in the assumption that
(4.28) $\tau^{\prime}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})],$
for any $(t,x,v)$ which satisfies (4.26), for all
(4.29) $\begin{split}&\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+1}(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime}))}\right)\leq s\\ \leq&\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+2}(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime}))}\right),\end{split}$
we have
(4.30) $\begin{split}|X_{s,t}(x,v)-x_{i}|\leq\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k}}.\end{split}$
Indeed, we have
(4.31) $|X_{s,t}(x,v)-x_{i}|\leq|X_{s,t}(x,v)-x|+|x-x_{i}|\leq|v|(t-s)+\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}}.$
Plus, we have
(4.32) $|v|\leq r^{\prime\prime}_{k+1}+|\overline{v}_{i}(\tau^{\prime\prime})|\leq r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime}),$
$t-s\leq\tau^{\prime}-\left[\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+1}(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime}))}\right)\right]=\frac{\tau^{\prime}-\tau^{\prime\prime}}{2^{k+1}(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime}))}.$
Hence, we have (we also use (4.28) and (4.25))
(4.33) $\begin{split}&|X_{s,t}(x,v)-x_{i}|\leq\frac{\tau^{\prime}-\tau^{\prime\prime}}{2^{k+1}}+\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}}\leq\frac{\delta_{T,D}(\tau^{\prime\prime})}{2^{k+1}}+\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}}\leq\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k}}.\end{split}$
Since $r^{\prime\prime}_{k+1}(\tau^{\prime\prime})>r^{\prime\prime}_{k}(\tau^{\prime\prime})$, we have
$\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+1}\left(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}\right)>\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+1}\left(r^{\prime\prime}_{k}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}\right).$
Thus we can apply the induction assumption and deduce that
(4.34) $f(s,X_{s,t}(x,v),w)\geq\min\left(a_{k,\tau^{\prime\prime}}(\tau^{\prime}),b_{k,\tau^{\prime\prime}}(\tau^{\prime})\right)\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{\prime\prime}_{k})}(w),$
for any $w\in\mathbb{R}^{3}$, $s$ satisfying (4.29), $i=1,...,\mathcal{N}(N,\tau^{\prime\prime})$, and $(t,x,v)$ satisfying (4.26). Thus, we deduce from (4.34) and Lemma 3.2 that
(4.35) $\begin{split}&f(t,x,v)\\ \geq&\int_{\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+1}\left(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}\right)}^{\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+2}\left(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}\right)}\min\left(a_{k,\tau^{\prime\prime}}(\tau^{\prime}),b_{k,\tau^{\prime\prime}}(\tau^{\prime})\right)^{2}\\ &\times e^{-C_{L}(t-s)\langle r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\rangle^{\gamma^{+}}}C_{Q}(r^{\prime\prime}_{k})^{3+\gamma}\xi_{k+1}^{\frac{1}{2}}\mathbf{1}_{B\left(\overline{v}_{i}(\tau^{\prime\prime}),r^{\prime\prime}_{k+1}\right)}(v)\,ds\\ \geq&\int_{\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+1}\left(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}\right)}^{\tau^{\prime\prime}+(\tau^{\prime}-\tau^{\prime\prime})\left(1-\frac{1}{2^{k+2}\left(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}\right)}\min\left(a_{k,\tau^{\prime\prime}}(\tau^{\prime}),b_{k,\tau^{\prime\prime}}(\tau^{\prime})\right)^{2}\\ &\times e^{-C_{L}(t-s)\langle r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\rangle^{\gamma^{+}}}C_{Q}(r^{\prime\prime}_{k})^{3+\gamma}\xi_{k+1}^{\frac{1}{2}}\,ds\\ \geq&\min\left(a_{k,\tau^{\prime\prime}}(\tau^{\prime}),b_{k,\tau^{\prime\prime}}(\tau^{\prime})\right)^{2}C_{Q}(r^{\prime\prime}_{k})^{3+\gamma}\xi_{k+1}^{\frac{1}{2}}\frac{\tau^{\prime}-\tau^{\prime\prime}}{2^{k+2}(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime}))}\\ &\times e^{\frac{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\langle r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime})\rangle^{\gamma^{+}}}{2^{k+1}(r^{\prime\prime}_{k+1}+\mathfrak{R}_{\min}(\tau^{\prime\prime}))}}\\ =&a_{k+1,\tau^{\prime\prime}}(\tau^{\prime})\geq\min\left(a_{k+1,\tau^{\prime\prime}}(\tau^{\prime}),b_{k+1,\tau^{\prime\prime}}(\tau^{\prime})\right).\end{split}$
Hence, we showed that (4.24) holds for $n=k+1$ when $X_{s,t}(x,v)\in\Omega$ for $\tau^{\prime\prime}\leq s\leq t$.
Second case: $X_{s,t}(x,v)\notin\overline{\Omega}$ for some $s\in[\tau^{\prime\prime},t]$.
First, we observe that by definition of $t_{\partial}(x,v)$ in Definition 1.1, we have $t_{\partial}(x,v)\leq t-s\leq t-\tau^{\prime\prime}\leq\tau^{\prime}-\tau^{\prime\prime}$. Next, we notice by (4.25) that $\tau^{\prime}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime})]$. Then, since $\alpha<1$, the diffuse part yields a strictly positive contribution, thus we use (1.13), Proposition 4.1, and Lemma 3.1 to derive that for any $(t,x,v)$ satisfying (4.26), we have
(4.36) $\begin{split}&f(t,x,v)\\ \geq&(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v))>0}f(t,X_{t-t_{\partial}(x,v),t}(x,v),w)(w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v)))dw\right)\\ &\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}\exp\left[-\int_{t-t_{\partial}(x,v)}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right]\\ \geq&(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}e^{-t_{\partial}(x,v)C_{L}\left\langle v\right\rangle^{\gamma^{+}}}\\ \geq&(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{k+1})^{2}}{2T_{B}}}e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}{\left\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{k+1}\right\rangle}^{\gamma^{+}}}\\ \geq&b_{k+1,\tau^{\prime\prime}}(\tau^{\prime})\geq\min\left(a_{k+1,\tau^{\prime\prime}}(\tau^{\prime}),b_{k+1,\tau^{\prime\prime}}(\tau^{\prime})\right),\end{split}$
which shows that (4.24) holds for $n=k+1$ in the case when the backward characteristic $s\mapsto X_{s,t}(x,v)$ touches the boundary.
In conclusion, we showed that (4.24) holds for $n=k+1$ and by induction that (4.24) holds for $0\leq n\leq N$. ∎
Finally, we end the proof of Theorem 1.1 for case $\alpha\in[0,1)$. We pick $0<\xi<\frac{1}{4}$ and setting $\xi_{n}:=\xi^{n}$, by which we can deduce that (we recall $r_{n}^{\prime\prime}$ from (4.21))
(4.37) $c_{r}(\xi)2^{\frac{n}{2}}\leq r^{\prime\prime}_{n}=\delta_{V}(\tau^{\prime\prime})2^{\frac{n}{2}}\Pi_{i=1}^{n}(1-\xi^{i})\leq\delta_{V}(\tau^{\prime\prime})2^{\frac{n}{2}},$
where $c_{r}(\xi):=\delta_{V}(\tau^{\prime\prime})\Pi_{i=1}^{\infty}(1-\xi^{i})>0$.
We first show that given $\tau_{2}\in\left[\tau^{\prime\prime}+\min\left\{\frac{\delta_{T,D}(\tau^{\prime\prime})-\tau^{\prime\prime}}{2},\tau^{\prime\prime}\right\},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})\right]$ there exists $\mathcal{A}>0$ such that
(4.38) $f(\tau_{2},x,v)\geq\mathcal{A}^{2^{n}},\ \forall n\in\mathbb{N},\ (x,v)\in\overline{\Omega}\times B(0,c_{r}(\xi)2^{\frac{n}{2}}),$
to get a lower maxwellian lower bound of a continuous mild solution of (1.1)–(1.3).
We will use the proof from [1, 11] to show the existence of $\mathcal{A}_{1}>0$ and $\mathcal{A}_{2}>0$ such that
(4.39) $b_{n,\tau^{\prime\prime}}(\tau_{2})\geq\mathcal{A}_{1}^{2^{n}},\ a_{n,\tau^{\prime\prime}}(\tau_{2})\geq\mathcal{A}_{2}^{2^{n}},$
for any $0<\tau^{\prime\prime}<\tau_{2}\leq\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})$.
In this article, we only show the sketch of the proof. For details, see [1, 11]. First, we observe that for $n\geq 2$ (we recall that $\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq 1$)
$b_{n,\tau^{\prime\prime}}(\tau_{2})$
$=$ $(1-\alpha)b(\tau^{\prime\prime})e^{-C_{L}(\tau_{2}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n}\rangle^{\gamma^{+}}}\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})^{2}}{2T_{B}}}$
$=$ $(1-\alpha)b(\tau^{\prime\prime})e^{-C_{L}(\tau_{2}-\tau^{\prime\prime})(1+\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})^{\gamma^{+}}}\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})^{2}}{2T_{B}}}$
$\geq$ $(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}(2+r^{\prime\prime}_{n})^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})^{2}}{2T_{B}}}$
$\geq$ $(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-8C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}}e^{-\frac{\mathfrak{R}_{\min}(\tau^{\prime\prime})^{2}}{T_{B}}}e^{-\left((2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}2^{n}\right)}$
$\geq$ $C_{1}e^{-\left((2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}2^{n}\right)},$
where we defined $C_{1}:=(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-8C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}}e^{-\frac{\mathfrak{R}_{\min}(\tau^{\prime\prime})^{2}}{T_{B}}}$.
By defining
$\mathcal{A}_{1}(\tau_{2}):=\begin{cases}\min\left\{b_{0,\tau^{\prime\prime}}(\tau_{2}),\sqrt{b_{1,\tau^{\prime\prime}}(\tau_{2})},e^{-\left((2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}\right)}\right\},\ C_{1}\geq 1\\ \min\left\{b_{0,\tau^{\prime\prime}}(\tau_{2}),\sqrt{b_{1,\tau^{\prime\prime}}(\tau_{2})},C_{1}e^{-\left((2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}\right)}\right\},\ C_{1}<1,\end{cases}$
we conclude that for $n\in\mathbb{N}$
(4.40) $b_{n,\tau^{\prime\prime}}(\tau_{2})\geq\mathcal{A}_{1}^{2^{n}}(\tau_{2}).$
For $a_{n,\tau^{\prime\prime}}(\tau_{2})$, we also recall the definition (for $n\geq 1$):
$a_{n,\tau^{\prime\prime}}(\tau_{2})=(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}C_{Q}r_{n-1}^{{}^{\prime\prime}3+\gamma}\xi^{\frac{n}{2}}\frac{\tau_{2}-\tau^{\prime\prime}}{2^{n+1}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})}$
$\hskip 54.06006pt\times e^{-C_{L}\frac{(\tau_{2}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n}\rangle^{\gamma^{+}}}{2^{n}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})}}.$
We notice that for $n\geq 1$, we have
$\frac{(\tau_{2}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n}\rangle^{\gamma^{+}}}{2^{n+1}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})}\leq\frac{\tau_{2}(1+\mathfrak{R}_{\min}(\tau^{\prime\prime}))^{\gamma^{+}}}{\mathfrak{R}_{\min}(\tau^{\prime\prime})}+\frac{\tau_{2}(\delta_{V}(\tau^{\prime\prime}))^{\gamma^{+}}}{\mathfrak{R}_{\min}(\tau^{\prime\prime})}.$
By defining
$C_{2}:=C_{Q}\min\left\{\frac{\delta_{T,D}(\tau^{\prime\prime})-\tau^{\prime\prime}}{2},\tau^{\prime\prime}\right\}e^{-C_{L}\left(\frac{\tau_{2}(1+\mathfrak{R}_{\min}(\tau^{\prime\prime}))^{\gamma^{+}}}{\mathfrak{R}_{\min}(\tau^{\prime\prime})}+\frac{\tau_{2}(\delta_{V}(\tau^{\prime\prime}))^{\gamma^{+}}}{\mathfrak{R}_{\min}(\tau^{\prime\prime})}\right)},$
we have the following inequality:
$a_{n,\tau^{\prime\prime}}(\tau_{2})$
$\geq$ $C_{2}(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}r_{n-1}^{{}^{\prime\prime}3+\gamma}\xi^{\frac{n}{2}}\frac{1}{2^{n+1}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})}$
$\geq$ $C_{2}(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}(c_{r}(\xi)2^{\frac{n-1}{2}})^{3+\gamma}\xi^{\frac{n}{2}}\frac{1}{2^{n+1}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+\delta_{V}(\tau^{\prime\prime})2^{\frac{n}{2}})}$
$\geq$ $C_{2}(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}(\frac{c_{r}(\xi)}{2})^{3+\gamma}(2^{\frac{3+\gamma}{2}})^{n}\xi^{\frac{n}{2}}\frac{1}{2^{n+1}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+\delta_{V}(\tau^{\prime\prime})2^{\frac{n}{2}})}$
$\geq$ $C_{2}(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}(\frac{c_{r}(\xi)}{2})^{3+\gamma}(2^{\frac{3+\gamma}{2}})^{n}\xi^{\frac{n}{2}}\frac{1}{4}\frac{1}{2^{n}}\min\{\frac{1}{\mathfrak{R}_{\min}(\tau^{\prime\prime})},\frac{1}{\delta_{V}(\tau^{\prime\prime})2^{\frac{n}{2}}}\}$
$\geq$ $C_{2}(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}(\frac{c_{r}(\xi)}{2})^{3+\gamma}\frac{1}{4}(2^{\frac{1+\gamma}{2}}\xi^{\frac{1}{2}})^{n}\min\{\frac{1}{\mathfrak{R}_{\min}(\tau^{\prime\prime})},\frac{1}{\delta_{V}(\tau^{\prime\prime})2^{\frac{n}{2}}}\}$
$\geq$ $C_{2}(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}(\frac{c_{r}(\xi)}{2})^{3+\gamma}\frac{1}{4}(2^{\frac{1+\gamma}{2}}\xi^{\frac{1}{2}})^{n}$
$\times\min\left\{\frac{1}{\max\{\mathfrak{R}_{\min}(\tau^{\prime\prime}),\delta_{V}(\tau^{\prime\prime})\}},\frac{1}{\max\{\mathfrak{R}_{\min}(\tau^{\prime\prime}),\delta_{V}(\tau^{\prime\prime})\}2^{\frac{n}{2}}}\right\}$
$\geq$ $C_{2}(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}(\frac{c_{r}(\xi)}{2})^{3+\gamma}\frac{1}{4(\mathfrak{R}_{\min}(\tau^{\prime\prime})+\delta_{V}(\tau^{\prime\prime}))}(2^{\frac{1+\gamma}{2}}\xi^{\frac{1}{2}})^{n}\frac{1}{2^{\frac{n}{2}}}$
$\geq$ $C_{2}(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}\left(\frac{c_{r}(\xi)}{2}\right)^{3+\gamma}\frac{1}{4(\mathfrak{R}_{\min}(\tau^{\prime\prime})+\delta_{V}(\tau^{\prime\prime}))}(2^{\frac{\gamma}{2}}\xi^{\frac{1}{2}})^{n}.$
By defining
(4.41) $C_{3}:=C_{2}(\frac{c_{r}(\xi)}{2})^{3+\gamma}\frac{1}{4(\mathfrak{R}_{\min}(\tau^{\prime\prime})+\delta_{V}(\tau^{\prime\prime}))},\ \lambda:=\min\{1,2^{\frac{\gamma}{2}}\xi^{\frac{1}{2}},C_{3}2^{\frac{\gamma}{2}}\xi^{\frac{1}{2}}\},$
we have
(4.42) $a_{n,\tau^{\prime\prime}}(\tau_{2})\geq(\min(a_{n-1,\tau^{\prime\prime}}(\tau_{2}),b_{n-1,\tau^{\prime\prime}}(\tau_{2})))^{2}\lambda^{n}.$
Next, note that by (4.22) and (4.23), we have
$b_{0,\tau^{\prime\prime}}(\tau_{2})\leq a_{0,\tau^{\prime\prime}}(\tau_{2}),\ b_{1,\tau^{\prime\prime}}(\tau_{2})\leq a_{1,\tau^{\prime\prime}}(\tau_{2}).$
Then, we define the following numbers:
(4.43) $l_{n}:=\min\{1\leq l\leq n-1\mid a_{n-l,\tau^{\prime\prime}}(\tau_{2})\geq b_{n-l,\tau^{\prime\prime}}(\tau_{2})\}.$
In doing so, we can iterate (4.42) and use (4.40) to derive:
(4.44) $\begin{split}a_{n,\tau^{\prime\prime}}(\tau_{2})&\geq\lambda^{n+2(n-1)+...+2^{l_{n}-1}(n-l_{n}+1)}(b_{n-l_{n},\tau^{\prime\prime}}(\tau_{2}))^{2^{l_{n}}}\\ &\geq\lambda^{n+2(n-1)+...+2^{l_{n}-1}(n-l_{n}+1)}\mathcal{A}_{1}^{2^{n-l_{n}}2^{l_{n}}}\\ &\geq\lambda^{\sum_{i=0}^{l_{n}-1}2^{i}(n-i)}\mathcal{A}_{1}^{2^{n}}.\end{split}$
Next, we notice that
(4.45) $\begin{split}&\sum_{i=0}^{l_{n}-1}2^{i}(n-i)\leq 2^{l_{n}}(n-l_{n}+2).\end{split}$
Then, we notice that for any $0\leq m\leq n-1<\infty$
(4.46) $\begin{split}&2^{n-1-m}(m+3)\leq 2^{n+1}.\end{split}$
By replacing $m$ in (4.46) by $n-l_{n}-1$, we have
$2^{l_{n}}(n-l_{n}+2)\leq 2^{n+1}.$
As a result, we conclude from (4.44), (4.45), and (4.46) that
$a_{n,\tau^{\prime\prime}}(\tau_{2})\geq\lambda^{2^{l_{n}}(n-l_{n}+2)}\mathcal{A}_{1}^{2^{n}}\geq\lambda^{2^{n+1}}\mathcal{A}_{1}^{2^{n}}.$
Notice that we used the fact that $\lambda\leq 1$. Hence, by defining $\mathcal{A}_{2}(\tau_{2}):=\lambda^{2}\mathcal{A}_{1}(\tau_{2})$, we conclude the proof of Theorem 1.1 in the case $0\leq\alpha<1$ where
$t\in\left[\tau^{\prime\prime}+\min\left\{\frac{\delta_{T,D}(\tau^{\prime\prime})-\tau^{\prime\prime}}{2},\tau^{\prime\prime}\right\},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})\right].$
To deduce the Theorem 1.2 from Theorem 1.1. We observe that
We consider the first term of (1.12) and (1.13). Then we use Lemma 3.1 to control the damping effect along the characteristic line $X_{s,t}(x,v)$. Thus, the solution is a super solution of the damp transport equation. That is, we either have
(4.47) $\begin{split}f(t,x,v)&\geq f_{0}(X_{0,t}(x,v),v)e^{-tC_{L}(1+|v|^{2})},\end{split}$
when $t\leq t_{\partial}(x,v)$, or
(4.48) $\begin{split}f(t,x,v)&=\alpha f(t-t_{\partial}(x,v),X_{t-t_{\partial}(x,v),t}(x,v),R(X_{t-t_{\partial}(x,v),t}(x,v),v))\\ &\hskip 14.22636pte^{-t_{\partial}(x,v)C_{L}(1+|v|^{2})}\\ &+(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v))>0}f(t,X_{t-t_{\partial}(x,v),t}(x,v),w)(w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v)))dw\right)\\ &\hskip 14.22636pt\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}e^{-t_{\partial}(x,v)C_{L}(1+|v|^{2})},\end{split}$
when $t\geq t_{\partial}(x,v)$. Then Theorem 1.2 can be deduced from the comparison principle.
5. Maxwellian bound in the cut-off case for fully specular reflection condition
In this chapter, we consider the fully specular reflection condition $\alpha=1$.
We first introduce with omitted proof the following estimate of the lower bound of away from the boundary, which is similar to Proposition 4.1, [2]. We note that the argument does not use the convexity of $\Omega$.
Proposition 5.1.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$. We consider a continuous mild solution $f(t,x,v)$ of (1.1)–(1.3) with $\alpha=1$. Fix $\tau^{\prime\prime}$ as in Proposition 3.6. Then, for any $\tau^{\prime}\in\left(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T}(\tau^{\prime\prime})\right]$ and $N\in\mathbb{N}$, the following holds: for any $1\leq n\leq N$, $0<l\leq\delta_{X}(\tau^{\prime\prime})$, and
$\mathfrak{R}>\max\left\{\frac{l}{\tau^{\prime}-\tau^{\prime\prime}},\mathfrak{R}_{\min}(\tau^{\prime\prime})+(\frac{3\sqrt{2}}{4})^{N}\delta_{V}(\tau^{\prime\prime})\right\},$
if $(t,x,v)\in\left[\tau^{\prime\prime}+\frac{l}{2^{n}\mathfrak{R}},\tau^{\prime}\right]\times\overline{\Omega}\times B(0,\mathfrak{R})$, and there exists $t_{1}\in\left[\tau^{\prime\prime},t-\frac{l}{2^{n}\mathfrak{R}}\right]$, we have $X_{t_{1},t}(x,v)\in(\Omega-\Omega_{l})\cap B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{n}}\right)$ for some $1\leq i\leq\mathcal{N}(N,\tau^{\prime\prime})$, then we have
(5.1) $\ \ \ \ \ f(t,x,v)\geq a^{S}_{n}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{n})}(V_{t_{1},t}(x,v)).$
Here, the notation $\delta_{X}(\tau^{\prime\prime})$, $\delta_{V}(\tau^{\prime\prime})$, $\delta_{T}(\tau^{\prime\prime})$, and $\overline{v}_{i}(\tau^{\prime\prime})$ are as in Proposition 3.6 and may depend on $N$. The sequence $r^{S}_{n}$ and $a^{S}_{n}(l,\tau^{\prime})$ are defined by
(5.2) $\begin{cases}r^{S}_{0}(\tau^{\prime\prime})=\delta_{V}(\tau^{\prime\prime}),\\ r^{S}_{n+1}(\tau^{\prime\prime})=\frac{3\sqrt{2}}{4}r^{S}_{n}(\tau^{\prime\prime}),\end{cases}$
(5.3) $\begin{cases}a^{S}_{0}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})=a_{0}(\tau^{\prime\prime})e^{-C_{L}(\tau^{\prime}-\tau^{\prime\prime})\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}},\\ a^{S}_{n+1}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})=C_{Q}\frac{(r^{S}_{n}(\tau^{\prime\prime}))^{3+\gamma}l}{2^{n+4}\mathfrak{R}}e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}a^{S}_{n}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2},\end{cases}$
where we recall that $a_{0}(\tau^{\prime\prime})$, $\mathfrak{R}_{\min}(\tau^{\prime\prime})$ are defined in Proposition 3.6.
Proof.
We prove (5.1) by using an induction on n. To show that (5.1) holds in the case $n=0$, we first consider the first term of the right hand side of (1.14), which implies by Lemma 3.1 that
(5.4) $\begin{split}f(t,x,v)\geq f(t_{1},X_{t_{1},t}(x,v),V_{t_{1},t}(x,v))e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}.\end{split}$
Then we use Proposition 3.6 to derive
(5.5) $\begin{split}&f(t_{1},X_{t_{1},t}(x,v),V_{t_{1},t}(x,v))\\ &\geq a_{0}(\tau^{\prime\prime})\mathbf{1}_{(B(x_{i},\delta_{X}(\tau^{\prime\prime}))\cap\overline{\Omega})\times B(\overline{v}_{i}(\tau^{\prime\prime}),\delta_{V}(\tau^{\prime\prime}))}(X_{t_{1},t}(x,v),V_{t_{1},t}(x,v))\\ &=a_{0}(\tau^{\prime\prime})\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),\delta_{V}(\tau^{\prime\prime}))}(V_{t_{1},t}(x,v)),\end{split}$
where we used the assumption that $X_{t_{1},t}(x,v)\in(\Omega-\Omega_{l})\cap B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{n}}\right)$. By combining (5.4) and (5.5), we proved the case $n=0$.
Now, we assume that the case $n=k\leq N-1$ holds. Given $0<l\leq\delta_{X}(\tau^{\prime\prime})$, $(t,x,v)\in[\tau^{\prime\prime}+\frac{l}{2^{k+1}\mathfrak{R}},\tau^{\prime}]\times\overline{\Omega}\times B(0,\mathfrak{R})$ with $t_{1}\in[\tau^{\prime\prime},t-\frac{l}{2^{k+1}\mathfrak{R}}]$ such that $X_{t_{1},t}(x,v)\in(\Omega-\Omega_{l})\cap B(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}})$, we consider the second term of the right hand side of (1.14) and use Lemma 3.1 to derive that
(5.6) $\begin{split}&f(t,x,v)\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}\\ &\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds,\end{split}$
where we use the fact that $t\geq t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}$. Next, we notice that
(5.7) $\begin{split}|X_{t_{1},t}(x,v)-X_{s,t}(x,v)|\leq(s-t_{1})|v|<\frac{l}{2^{k+2}}\end{split}$
for any $(t,x,v)\in[\tau^{\prime\prime}+\frac{l}{2^{k+1}\mathfrak{R}},\tau^{\prime}]\times\overline{\Omega}\times B(0,\mathfrak{R})$ with $t_{1}\in[\tau^{\prime\prime},t-\frac{l}{2^{k+1}\mathfrak{R}}]$, $s\in[t_{1}+\frac{l}{2^{k+3}\mathfrak{R}},t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}]$, which implies that
(5.8) $X_{s,t}(x,v)\in\Omega-\Omega_{l-\frac{l}{2^{k+2}}}.$
We also deduce that
(5.9) $V_{s,t}(x,v)=V_{t_{1},t}(x,v)$
since under specular reflection, the backward characteristic velocity is piecewise constant and identical on $t_{1},s]$ without collisions. Next, we notice that for $u\in B(0,\mathfrak{R})$,
(5.10) $\begin{split}&|X_{t_{1},s}(X_{s,t}(x,v),u)-X_{t_{1},t}(x,v)|\\ \leq&|X_{t_{1},s}(X_{s,t}(x,v),u)-X_{s,t}(x,v)|+|X_{s,t}(x,v)-X_{t_{1},t}(x,v)|\\ \leq&(s-t_{1})|u|+(s-t_{1})|v|\leq\frac{l}{2^{k+1}}\leq\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}},\end{split}$
which implies
(5.11) $X_{t_{1},s}(X_{s,t}(x,v),u)\in(\Omega-\Omega_{\frac{l}{8}})\cap B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k}}\right),$
and
(5.12) $V_{t_{1},s}(X_{s,t}(x,v),u)=u,$
for $u\in B(0,\mathfrak{R})$. In addition, for any $s\in[t_{1}+\frac{l}{2^{k+3}\mathfrak{R}},t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}]$, we have
(5.13) $\tau^{\prime\prime}+\frac{\frac{l}{8}}{2^{k}\mathfrak{R}}\leq s\leq\tau^{\prime},$
(5.14) $t_{1}\in\left[\tau^{\prime\prime},s-\frac{\frac{l}{8}}{2^{k}\mathfrak{R}}\right].$
As a result, the assumptions of induction are satisfied and we deduce that
(5.15) $\begin{split}f(s,X_{s,t}(x,v),u)\geq a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k})}(u).\end{split}$
for $(t,x,v)\in\left[\tau^{\prime\prime}+\frac{l}{2^{k+1}\mathfrak{R}},\tau^{\prime}\right]\times\overline{\Omega}\times B(0,\mathfrak{R})$ and $t_{1}\in[\tau^{\prime\prime},t-\frac{l}{2^{k+1}\mathfrak{R}}]$ with $s\in\left[t_{1}+\frac{l}{2^{k+3}\mathfrak{R}},t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}\right]$, $u\in B(0,\mathfrak{R})$. Here, we used (5.12).
Hence, we have
(5.16) $\begin{split}&Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\\ \geq&a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}Q^{+}[\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k})\cap B(0,\mathfrak{R})}(\cdot),\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k})\cap B(0,\mathfrak{R})}(\cdot)](V_{s,t}(x,v))\\ =&a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}Q^{+}[\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k})}(\cdot),\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k})}(\cdot)](V_{s,t}(x,v)).\end{split}$
Here, we use the fact that $\mathfrak{R}>\mathfrak{R}_{\min}(\tau^{\prime\prime})+(\frac{3\sqrt{2}}{4})^{N}\delta_{V}(\tau^{\prime\prime})\geq|\overline{v}_{i}(\tau^{\prime\prime})|+r^{S}_{k}$ for the derivation of the last line. Now, we use Lemma 3.2 with $\xi=\frac{1}{2}$ to derive
(5.17) $\begin{split}Q^{+}[\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k})}(\cdot),\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k})}(\cdot)](V_{s,t}(x,v))\geq C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{s,t}(x,v)).\end{split}$
Hence, we deduce by (5.6), (5.9), (5.16), and (5.17) that
(5.18) $\begin{split}&f(t,x,v)\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{s,t}(x,v))\,ds\\ =&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{s,t}(x,v))\,ds\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\frac{l}{2^{k+3}\mathfrak{R}}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{t_{1},t}(x,v))\\ \geq&a^{S}_{k+1}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{t_{1},t}(x,v)).\end{split}$
As a result, the induction hypothesis for $n=k+1$ holds. By induction, we deduce that (5.1) holds for $0\leq n\leq N$ and we conclude the result.
The following corollary is obtained by a slight modification of the previous proposition (Corollary 4.2, [2]):
Corollary 5.2.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha=1$. Fix $\tau^{\prime\prime}>0$ as in Proposition 3.6. Then, given $\Delta^{I}_{T}\in(0,\delta_{T}(\tau^{\prime\prime})]$ with $\tau^{A}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T}]$, for all $0<l\leq\delta_{X}(\tau^{\prime\prime})$, there exist $a^{S}(l,\tau^{A},\tau^{\prime\prime},\Delta^{I}_{T})>0$ and $\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T})\in\left(0,\frac{\tau^{A}-\tau^{\prime\prime}}{3}\right]$, depending on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$(and $L_{f,p}$ if $\gamma<0$) such that we have the following: for $(t,x,v)\in[\tau^{A},\tau^{\prime\prime}+\Delta^{I}_{T}]\times\overline{\Omega}\times\mathbb{R}^{3}$, if for some $t_{1}\in[\tau^{\prime\prime},t-\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T})]$ we have $X_{t_{1},t}(x,v)\in\Omega-\Omega_{l}$, then we have
(5.19) $\ \ \ \ \ f(t,x,v)\geq a^{S}(l,\tau^{A},\tau^{\prime\prime},\Delta^{I}_{T})\mathbf{1}_{B(0,3\mathfrak{R}_{\min}(\tau^{\prime\prime}))}(v).$
Here, $\mathfrak{R}_{\min},\delta_{T}(\tau^{\prime\prime})$, and $\delta_{X}(\tau^{\prime\prime})$ are given in Proposition 3.6.
Proof.
Given $0<l\leq\delta_{X}(\tau^{\prime\prime})$ and $\tau^{A}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T}]$. We define
(5.20) $\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A}):=\left\lceil\max\left\{\log_{\frac{3\sqrt{2}}{4}}\left(\frac{4\mathfrak{R}_{\min}(\tau^{\prime\prime})}{\delta_{V}(\tau^{\prime\prime})}\right),\log_{2}\frac{3l}{\tau^{A}-\tau^{\prime\prime}}\right\}\right\rceil,$
(5.21) $\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T}):=\frac{l}{2^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}\left[\max\left\{\frac{l}{\Delta^{I}_{T}},\mathfrak{R}_{\min}(\tau^{\prime\prime})+\left(\frac{3\sqrt{2}}{4}\right)^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}\delta_{V}(\tau^{\prime\prime})\right\}+1\right]},$
(5.22) $\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A}):=1+\max\left\{\frac{l}{\tau^{A}-\tau^{\prime\prime}},\mathfrak{R}_{\min}(\tau^{\prime\prime})+\left(\frac{3\sqrt{2}}{4}\right)^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}\delta_{V}(\tau^{\prime\prime})\right\}.$
By Proposition 3.6, we know that $\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq 2$ and $\delta_{V}(\tau^{\prime\prime})\leq\frac{1}{56}$, which imply $\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})\geq 8$. We observe that
(5.23) $\begin{split}\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T})\leq\frac{l}{2^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}}\leq\frac{\tau^{A}-\tau^{\prime\prime}}{3}.\end{split}$
The last inequality follows directly from the definition of $\widetilde{N}_{l}$, which leads to $2^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}\geq\frac{3l}{\tau^{A}-\tau^{\prime\prime}}$.
Now, given $(t,x,v)\in[\tau^{A},\tau^{\prime\prime}+\Delta^{I}_{T}]\times\overline{\Omega}\times\mathbb{R}^{3}$ with $t_{1}\in[\tau^{\prime\prime},t-\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T})]$ such that $X_{t_{1},t}(x,v)\in\Omega-\Omega_{l}$. We consider $\{x_{i}\}_{i=1}^{\mathcal{N}(\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A}),\tau^{\prime\prime})}$ from Proposition 3.6 such that
$\overline{\Omega}\subset\bigcup_{1\leq i\leq\mathcal{N}(\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A}),\tau^{\prime\prime})}B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}}\right)$
. There exists $1\leq i\leq\mathcal{N}(\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A}),\tau^{\prime\prime})$ such that $X_{t_{1},t}(x,v)\in B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}}\right)$. Next, we find that
(5.24) $\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T})\geq\frac{l}{2^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A})},$
which leads to
(5.25) $t_{1}\in\left[\tau^{\prime\prime},t-\frac{l}{2^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A})}\right].$
Moreover, we have
(5.26) $\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A})>\max\left\{\frac{l}{\Delta^{I}_{T}},\mathfrak{R}_{\min}(\tau^{\prime\prime})+(\frac{3\sqrt{2}}{4})^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}\delta_{V}(\tau^{\prime\prime})\right\}.$
Thus, we apply Proposition 5.1 with $\tau^{\prime}=\tau^{\prime\prime}+\Delta^{I}_{T}$, $n=N=\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})$, and $\mathfrak{R}=\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A})$ and deduce that for any
$(t,x,v)\in\left[\tau^{\prime\prime}+\frac{l}{2^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A})},\tau^{\prime\prime}+\Delta^{I}_{T}\right]\times\overline{\Omega}\times B(0,\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A})),$
we have
$f(t,x,v)$ $\geq a^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(l,\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T},\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A}))\mathbf{1}_{B\left(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(\tau^{\prime\prime})\right)}(V_{t_{1},t}(x,v))$
$\geq a^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(l,\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T},\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A}))\mathbf{1}_{B\left(0,3\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}(V_{t_{1},t}(x,v))$
$=a^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(l,\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T},\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A}))\mathbf{1}_{B(0,3\mathfrak{R}_{\min}(\tau^{\prime\prime}))}(v).$
Here, we used the fact that $r^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(\tau^{\prime\prime})\geq 4\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq 3\mathfrak{R}_{\min}(\tau^{\prime\prime})+|\overline{v}_{i}(\tau^{\prime\prime})|$ for the second line.
Moreover, we have
$\tau^{A},\tau^{\prime\prime}+\Delta^{I}_{T}]\subset\left[\tau^{\prime\prime}+\frac{l}{2^{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A})},\tau^{\prime\prime}+\Delta^{I}_{T}\right].$
Finally, by the choice of $\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})$, we have
$\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A})\geq 4\mathfrak{R}_{\min}(\tau^{\prime\prime}).$
As a result, by defining
(5.27) $a^{S}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T}):=a^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(l,\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T},\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A})).$
we obtain the desired lower bound (5.19), which completes the proof.
Before stating the next proposition, we recall the quantities $\delta_{X}(\tau^{\prime\prime})$, $\delta_{V}(\tau^{\prime\prime})$, $\delta_{T}(\tau^{\prime\prime})$, and $\mathfrak{R}_{\min}(\tau^{\prime\prime})$ from Proposition 3.6, and notice that $\delta_{T}(\tau^{\prime\prime})\leq\delta_{X}(\tau^{\prime\prime})=\delta_{V}(\tau^{\prime\prime})\leq\frac{d}{56}$ and $\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq 2$. We now introduce the following auxiliary sequences:
(5.28) $\begin{cases}r^{B}_{0}(\tau^{\prime\prime})=\delta_{V}(\tau^{\prime\prime}),\\ r^{B}_{n+1}(\tau^{\prime\prime})=\frac{3\sqrt{2}}{4}r^{B}_{n}(\tau^{\prime\prime})-\frac{\delta_{V}(\tau^{\prime\prime})}{40}.\end{cases}$
Observe that $\{r^{B}_{n}(\tau^{\prime\prime})\}_{n=0}^{\infty}$ is strictly increasing and unbounded, that
$r^{B}_{0}(\tau^{\prime\prime})<1\leq\mathfrak{R}_{\min}(\tau^{\prime\prime}),$
and that
(5.29) $\frac{2\mathfrak{R}_{\min}(\tau^{\prime\prime})}{\mathfrak{R}_{\min}(\tau^{\prime\prime})+1}\geq\frac{4}{3}>\frac{3\sqrt{2}}{4}.$
Hence, the following notation
(5.30) $N_{B}(\tau^{\prime\prime}):=\min\{n\mid 2\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq r^{B}_{n}(\tau^{\prime\prime})>\mathfrak{R}_{\min}(\tau^{\prime\prime})+1\}$
is well-defined. We also define
(5.31) $\vartheta_{n}^{B}(\tau^{\prime\prime}):=\frac{1}{2^{n+1}\left(r^{B}_{n}(\tau^{\prime\prime})+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}.$
(5.32) $\delta_{T}^{B}(\tau^{\prime\prime}):=\min\left\{t_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}\left(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{B}_{N_{B}(\tau^{\prime\prime})}\right),\delta_{T}(\tau^{\prime\prime}),\frac{d}{4\left(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r_{N_{B}(\tau^{\prime\prime})}^{B}(\tau^{\prime\prime})\right)}\right\},$
(5.33) $a^{B}_{n,\tau^{\prime\prime}}(\tau^{B}):=\begin{cases}a_{0}(\tau^{\prime\prime}),\ &n=0\\ \left(\min\{a^{B}_{n-1,\tau^{\prime\prime}}(\tau^{B}),b^{B}_{\tau^{\prime\prime}}(\tau^{B})\}\right)^{2}C_{Q}(r^{B}_{n-1}(\tau^{\prime\prime}))^{3+\gamma}\\ \frac{\tau^{B}-\tau^{\prime\prime}}{2^{n+2}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{B}_{n}(\tau^{\prime\prime}))}e^{-C_{L}\frac{(\tau^{B}-\tau^{\prime\prime})\left\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{B}_{n}(\tau^{\prime\prime})\right\rangle^{\gamma^{+}}}{2^{n}\left(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{B}_{n}(\tau^{\prime\prime})\right)}},\ &n\geq 1,\end{cases}$
(5.34) $b^{B}_{\tau^{\prime\prime}}(\tau^{B}):=a^{S}\left(l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}\left(\frac{\tau^{B}-\tau^{\prime\prime}}{6}\right),\tau^{B},\tau^{\prime\prime},\delta_{T}^{B}(\tau^{\prime\prime})\right).$
Here, the notation $a_{0}(\tau^{\prime\prime})$ is from Proposition 3.6 and the notation $a^{S}$ is from Corollary 5.2.
Proposition 5.3.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B. Fix $d$ with $0<d<\min\{1,\delta\}$ as in Lemma 2.1. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha=1$ and let $\tau^{\prime\prime}>0$ be as given in Proposition 3.6. The following holds: For any $\tau^{B}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T}^{B}(\tau^{\prime\prime})]$, there exist $a^{B}_{0}(\tau^{B})>0$, which depends on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$, such that
(5.35) $f(t,x,v)\geq\min\left\{a^{B}_{N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime}}(\tau^{B}),b^{B}_{\tau^{\prime\prime}}(\tau^{B})\right\}\mathbf{1}_{B(0,1)}(v)$
for any
$(t,x,v)\in\left[\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-\vartheta_{N_{B}(\tau^{\prime\prime})}^{B}(\tau^{\prime\prime})\right),\tau^{B}\right]\times\Omega\times\mathbb{R}^{3}.$
Here, $N_{B}(\tau^{\prime\prime})$ is defined in (5.30) and $\mathfrak{R}_{\min}(\tau^{\prime\prime})$ is introduced in Proposition 3.6 and $\vartheta_{N_{B}(\tau^{\prime\prime})}$ is from (5.31).
Proof.
We start in a similar way to the proof of Proposition 4.2. We apply Proposition 3.6 to $\overline{\Omega}$ to deduce the existence of points $\{x_{j}\}_{j=1}^{\mathcal{N}(N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime})}\subset\Omega$ such that $\overline{\Omega}\subset\underset{1\leq j\leq\mathcal{N}(N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime})}{\bigcup}B(x_{j},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{N_{B}(\tau^{\prime\prime})}})$ and
(5.36) $f(t,x,v)\geq a_{0}(\tau^{\prime\prime})\mathbf{1}_{B(\overline{v}_{j}(\tau^{\prime\prime}),\delta_{V}(\tau^{\prime\prime}))}(v)$
for any $1\leq j\leq\mathcal{N}(N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime})$ and any $(t,x,v)\in[\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T}(\tau^{\prime\prime})]\times[B(x_{j},\delta_{X}(\tau^{\prime\prime}))\cap\overline{\Omega}]\times\mathbb{R}^{3}$.
Then, we recall the notation $t_{\epsilon}(v_{M})$ in (2.23) and $\delta_{T}^{B}(\tau^{\prime\prime})$ in (5.32)
(5.37) $\delta_{T}^{B}(\tau^{\prime\prime}):=\min\left\{t_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}\left(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{B}_{N_{B}(\tau^{\prime\prime})}\right),\delta_{T}(\tau^{\prime\prime}),\frac{d}{4\left(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r_{N_{B}(\tau^{\prime\prime})}^{B}(\tau^{\prime\prime})\right)}\right\},$
Now, given $\tau^{B}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T}^{B}(\tau^{\prime\prime})]$ and $0\leq n\leq N_{B}(\tau^{\prime\prime})$, we show that
(5.38) $\begin{split}&\ \ \ \ \ f(t,x,v)\geq\min\{a^{B}_{n,\tau^{\prime\prime}}(\tau^{B}),b^{B}_{\tau^{\prime\prime}}(\tau^{B})\}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{B}_{n}(\tau^{\prime\prime}))}(v),\\ &\forall(t,x)\in\left[\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-\vartheta_{n}^{B}(\tau^{\prime\prime})\right),\tau^{B}\right]\times B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{n}}\right)\end{split}$
for $1\leq i\leq\mathcal{N}(N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime})$.
The proof is similar to the proof of Proposition 4.2 and we use an induction on $n$. The base case $n=0$ follows directly from (5.36). Assume that the case $n=k$ holds. Given
(5.39) $\begin{split}(t,x,v)\in&\left[\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right),\tau^{B}\right]\\ &\times B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}}\right)\times B(\overline{v}_{i}(\tau^{\prime\prime}),r^{B}_{k+1}(\tau^{\prime\prime})),\end{split}$
we use the Duhamel formula (1.12) and apply Lemma 3.1 to deduce that
(5.40) $\begin{split}&f(t,x,v)\\ \geq&\int_{\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-2\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right)}^{\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right)}e^{-C_{L}(t-s)\left\langle r^{B}_{k+1}(\tau^{\prime\prime})+\mathfrak{R}_{\min}(\tau^{\prime\prime})\right\rangle^{\gamma^{+}}}\\ &Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds.\end{split}$
Since for any $s\in\left[\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-2\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right),t\right]$, we have
(5.41) $\begin{split}&|X_{s,t}(x,v)-x_{i}|\leq|X_{s,t}(x,v)-x|+|x-x_{i}|\leq|v|(t-s)+\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}}\\ \leq&\frac{\tau^{B}-\tau^{\prime\prime}}{2^{k+2}}+\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}}\leq\frac{\delta_{T}^{B}(\tau^{\prime\prime})}{2^{k+2}}+\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k+1}}\leq\frac{\delta_{X}(\tau^{\prime\prime})}{2^{k}}<\frac{d}{8}.\end{split}$
Here recall the definition of $\delta_{T}^{B}(\tau^{\prime\prime})$ and $\delta_{T}(\tau^{\prime\prime})$ from (5.32) and (3.61). We next apply the induction assumption and Lemma 3.2 to derive that
(5.42) $\begin{split}&Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\\ \geq&C_{Q}(r^{B}_{k}(\tau^{\prime\prime}))^{3+\gamma}\frac{1}{2}\min\left(a^{B}_{k,\tau^{\prime\prime}}(\tau^{B}),b^{B}_{\tau^{\prime\prime}}(\tau^{B})\right)^{2}\mathbf{1}_{B\left(\overline{v}_{i}(\tau^{\prime\prime}),r^{B}_{k+1}(\tau^{\prime\prime})+\frac{\delta_{V}(\tau^{\prime\prime})}{40}\right)}(V_{s,t}(x,v)).\end{split}$
We will generate the lower bound again depending on whether the characteristic line $\{X_{s}(x,v)\}_{\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-2\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right)\leq s\leq t}$ touches the boundary or not.
If $|V_{s,t}(x,v)-v|<\frac{\delta_{V}(\tau^{\prime\prime})}{40}$ for $\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-2\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right)\leq s\leq t$, we have
(5.43) $\mathbf{1}_{B\left(\overline{v}_{i}(\tau^{\prime\prime}),r^{B}_{k+1}(\tau^{\prime\prime})+\frac{\delta_{V}(\tau^{\prime\prime})}{40}\right)}(V_{s,t}(x,v))=1,$
for $\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-2\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right)\leq s\leq t$. We deduce that
(5.44) $\begin{split}&f(t,x,v)\\ \geq&C_{Q}(r^{B}_{k}(\tau^{\prime\prime}))^{3+\gamma}\frac{1}{2}\min\left(a^{B}_{k,\tau^{\prime\prime}}(\tau^{B}),b^{B}_{\tau^{\prime\prime}}(\tau^{B})\right)^{2}\\ &\int_{\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-2\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right)}^{\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right)}e^{-C_{L}(t-s)\langle r^{B}_{k+1}(\tau^{\prime\prime})+\mathfrak{R}_{\min}(\tau^{\prime\prime})\rangle^{\gamma^{+}}}\,ds\\ \geq&C_{Q}(r^{B}_{k}(\tau^{\prime\prime}))^{3+\gamma}\frac{1}{2}\min\left(a^{B}_{k,\tau^{\prime\prime}}(\tau^{B}),b^{B}_{\tau^{\prime\prime}}(\tau^{B})\right)^{2}\frac{\tau^{B}-\tau^{\prime\prime}}{2^{k+2}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{B}_{k+1}(\tau^{\prime\prime}))}\\ &\times e^{-C_{L}\frac{(\tau^{B}-\tau^{\prime\prime})\left\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{B}_{k+1}(\tau^{\prime\prime})\right\rangle^{\gamma^{+}}}{2^{k+1}\left(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{B}_{k+1}(\tau^{\prime\prime})\right)}}\\ =&a^{B}_{k+1,\tau^{\prime\prime}}(\tau^{B}).\end{split}$
On the other hand, if $|V_{s,t}(x,v)-v|\geq\frac{\delta_{V}(\tau^{\prime\prime})}{40}$ for some $\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-2\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right)\leq s\leq t$, we have $X_{s,t}(x,v)\in\partial\Omega$ for some
$s\in\left[\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-2\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right),t\right].$
By (5.41), it follows that $x_{i}\in\Omega_{\frac{d}{8}}$, so by Remark 2.2, there exists $1\leq k(i)\leq m_{1}$ such that $x_{i}\in B(x^{0}_{k(i)},\frac{d}{4})$. Notice that $|V_{s,t}(x,v)-v|\geq\frac{\delta_{V}(\tau^{\prime\prime})}{40}$ for some $\tau^{\prime\prime}\leq s\leq t$ Using Lemma 2.6, we deduce that for any $0\leq\tau_{2}\leq t-\tau^{\prime\prime}$, we have
$X_{s^{\prime},t}(x,v)\notin\Omega_{l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}(\tau_{2})}\cap B(x_{i}^{0},d)$
for some
$s^{\prime}\in[\tau^{\prime\prime},\tau^{\prime\prime}+\tau_{2}],$
which implies since $X_{s^{\prime},t}(x,v)\in B(x^{0}_{i},d)$:
$X_{s^{\prime},t}(x,v)\in\Omega-\Omega_{l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}(\tau_{2})}.$
Since we have
(5.45) $\begin{split}t-\tau^{\prime\prime}\geq\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-\vartheta_{k+1}^{B}(\tau^{\prime\prime})\right)-\tau^{\prime\prime}=\frac{\tau^{B}-\tau^{\prime\prime}}{2},\end{split}$
we can take $\tau_{2}=\frac{\tau^{B}-\tau^{\prime\prime}}{6}$. Next, we observe that
$\begin{split}\tau^{\prime\prime}+\tau_{2}=\tau^{\prime\prime}+\frac{\tau^{B}-\tau^{\prime\prime}}{3}\leq t-\frac{\tau^{B}-\tau^{\prime\prime}}{3}\leq t-\widetilde{t}\left(l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}\left(\frac{\tau^{B}-\tau^{\prime\prime}}{6}\right),\tau^{\prime\prime},\tau^{B},\delta_{T}^{B}(\tau^{\prime\prime})\right).\end{split}$
Consequently, we have
$s^{\prime}\in\left[\tau^{\prime\prime},t-\widetilde{t}\left(l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}\left(\frac{\tau^{B}-\tau^{\prime\prime}}{6}\right),\tau^{\prime\prime},\tau^{B},\delta_{T}^{B}(\tau^{\prime\prime})\right)\right]$
and
$X_{s^{\prime},t}(x,v)\in\Omega-\Omega_{l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}(\tau_{2})}.$
We also notice that
$l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}\left(\frac{\tau^{B}-\tau^{\prime\prime}}{6}\right)=\frac{(\tau^{B}-\tau^{\prime\prime})}{15360}\delta_{V}(\tau^{\prime\prime})\leq\delta_{X}(\tau^{\prime\prime}).$
Therefore, we apply Corollary 5.2 with parameters
$(l,\tau^{A},\tau^{\prime\prime},\Delta^{I}_{T})\rightarrow\left(l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}\left(\frac{\tau^{B}-\tau^{\prime\prime}}{6}\right),\tau^{B},\tau^{\prime\prime},\delta_{T}^{B}(\tau^{\prime\prime})\right)$
to deduce that
$f(t,x,v)\geq a^{S}\left(l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}\left(\frac{\tau^{B}-\tau^{\prime\prime}}{6}\right),\tau^{B},\tau^{\prime\prime},\delta_{T}^{B}(\tau^{\prime\prime})\right)\mathbf{1}_{B(0,3\mathfrak{R}_{\min}(\tau^{\prime\prime}))}(v)$
$=a^{S}\left(l_{\frac{\delta_{V}(\tau^{\prime\prime})}{40}}\left(\frac{\tau^{B}-\tau^{\prime\prime}}{6}\right),\tau^{B},\tau^{\prime\prime},\delta_{T}^{B}(\tau^{\prime\prime})\right),$
which coincides with the definition of $b^{B}_{\tau^{\prime\prime}}(\tau^{B})$ given above. Here, we used the fact
$|v|\leq|\overline{v}_{i}(\tau^{\prime\prime})|+r^{B}_{k+1}(\tau^{\prime\prime})\leq\mathfrak{R}_{\min}(\tau^{\prime\prime})+2\mathfrak{R}_{\min}(\tau^{\prime\prime})\leq 3\mathfrak{R}_{\min}(\tau^{\prime\prime}).$
[…]
(5.46) $\begin{split}f(t,x,v)&\geq\min\left\{a^{B}_{N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime}}(\tau^{B}),b^{B}_{\tau^{\prime\prime}}(\tau^{B})\right\}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r_{N_{B}(\tau^{\prime\prime})}^{B}(\tau^{\prime\prime}))}(v),\\ &\geq\min\left\{a^{B}_{N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime}}(\tau^{B}),b^{B}_{\tau^{\prime\prime}}(\tau^{B})\right\}\mathbf{1}_{B(0,1)}(v),\\ &\forall(t,x)\in\left[\tau^{\prime\prime}+(\tau^{B}-\tau^{\prime\prime})\left(1-\vartheta_{N_{B}(\tau^{\prime\prime})}^{B}(\tau^{\prime\prime})\right),\tau^{B}\right]\times B\left(x_{i},\frac{\delta_{X}(\tau^{\prime\prime})}{2^{N_{B}(\tau^{\prime\prime})}}\right),\end{split}$
which implies (5.35). This completes the proof.
Now, we introduce new notations:
Definition 5.1.
Given $0<\tau^{\prime\prime}<\tau^{S}$ as in Proposition 5.3 , we define
(5.47) $\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}):=\tau^{\prime\prime}+(\tau^{S}-\tau^{\prime\prime})\left(1-\vartheta_{N_{B}(\tau^{\prime\prime})}^{B}(\tau^{\prime\prime})\right).$
Clearly, we see that $\tau^{\prime\prime}<\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})<\tau^{S}$.
Definition 5.2.
Given $0<\tau^{\prime\prime}<\tau^{S}$, and a sequence $\{\xi_{n}\}\in(0,1)^{\mathbb{N}}$, we define the following numbers:
(5.48) $r^{\prime\prime}_{n,S}:=\begin{cases}1,\ n=0\\ \sqrt{2}(1-\xi_{n})r^{\prime\prime}_{n-1,S},\ n\geq 1,\end{cases}$
and
(5.49) $a^{S}_{n,\tau^{\prime\prime}}(\tau^{S}):=\begin{cases}\min\left\{a^{B}_{N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime}}(\tau^{S}),b^{B}_{\tau^{\prime\prime}}(\tau^{S})\right\},\ &n=0\\ (a^{S}_{n-1,\tau^{\prime\prime}}(\tau^{S}))^{2}C_{Q}(r^{\prime\prime}_{n-1,S})^{3+\gamma}\xi_{n}^{\frac{1}{2}}\frac{\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})}{2^{n+1}r^{\prime\prime}_{n,S}}\\ \times e^{-C_{L}\frac{(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left\langle r^{\prime\prime}_{n,S}\right\rangle^{\gamma^{+}}}{2^{n}r^{\prime\prime}_{n,S}}}\ &n\geq 1,\end{cases}$
where $a^{B}_{0}(\tau^{B})$ is defined in Proposition 5.3.
Now, we present a proposition similar to Proposition 4.2:
Proposition 5.4.
Let $\Omega\subset\mathbb{R}^{3}$ satisfying Assumption A, kernel $B$ satisfying Assumption B, and let $f(t,x,v)$ be the continuous mild solution of (1.1)–(1.3) with $\alpha=1$, $\tau^{\prime\prime}>0$ given from Proposition 3.6. Then, given $\tau^{S}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T}^{B}(\tau^{\prime\prime})]$, we have
(5.50) $\begin{split}&\ \ \ \ \ f(t,x,v)\geq a^{S}_{n,\tau^{\prime\prime}}(\tau^{S})\mathbf{1}_{B(0,r^{\prime\prime}_{n,S})}(v),\\ &\forall(t,x)\in\left[\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left(1-\frac{1}{2^{n+1}(r^{\prime\prime}_{n,S})}\right),\tau^{S}\right]\times\Omega,\end{split}$
for any $n\in\mathbb{N}$.
Proof.
The proof is the same as the one of Proposition 4.2. We prove by induction on $n$. The base case holds by applying Proposition 5.3 with the choice $\tau^{B}:=\tau^{S}$: we have
(5.51) $f(t,x,v)\geq\min\left\{a^{B}_{N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime}}(\tau^{S}),b^{B}_{\tau^{\prime\prime}}(\tau^{S})\right\}\mathbf{1}_{B(0,1)}(v)$
for any
$(t,x,v)\in\left[\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}),\tau^{S}\right]\times\Omega\times\mathbb{R}^{3}.$
We now assume that the case $n=k$ holds. We consider
(5.52) $\begin{split}&(t,x,v)\in\left[\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left(1-\frac{1}{2^{k+2}(r^{\prime\prime}_{k+1,S})}\right),\tau^{S}\right]\\ &\times\Omega\times B(0,r^{\prime\prime}_{k+1,S}).\end{split}$
We recall that by the Duhamel formula (1.14), and apply Lemma 3.1 to deduce that
(5.53) $\begin{split}&f(t,x,v)\\ \geq&\int_{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+1}r^{\prime\prime}_{k+1,S}})}^{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+2}r^{\prime\prime}_{k+1,S}})}e^{-C_{L}(t-s)\left\langle r^{\prime\prime}_{k+1,S}\right\rangle^{\gamma^{+}}}\\ &Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds.\end{split}$
Since we have
$\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left(1-\frac{1}{2^{k+1}(r^{\prime\prime}_{k+1,S})}\right)\geq\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left(1-\frac{1}{2^{k+1}(r^{\prime\prime}_{k,S})}\right),$
we can apply the induction assumption and deduce that
(5.54) $f(s,X_{s,t}(x,v),w)\geq a^{S}_{k,\tau^{\prime\prime}}(\tau^{S})\mathbf{1}_{B(0,r^{\prime\prime}_{k,S})}(w),$
for any $w\in\mathbb{R}^{3}$,
$s\in\left[\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left(1-\frac{1}{2^{k+1}r^{\prime\prime}_{k+1,S}}\right),\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left(1-\frac{1}{2^{k+1}r^{\prime\prime}_{k+2,S}}\right)\right]$
, and $(t,x,v)$ satisfying (5.52). Thus, we deduce from (5.54) that
(5.55) $\begin{split}&Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)]\\ \geq&(a^{S}_{k,\tau^{\prime\prime}}(\tau^{S}))^{2}Q^{+}[\mathbf{1}_{B(0,r^{\prime\prime}_{k,S})},\mathbf{1}_{B(0,r^{\prime\prime}_{k,S})}].\end{split}$
Then, we use Lemma 3.2 with $\xi=\xi_{k+1}$ to get
(5.56) $\begin{split}&Q^{+}[\mathbf{1}_{B(0,r^{\prime\prime}_{k,S})},\mathbf{1}_{B(0,r^{\prime\prime}_{k,S})}]\\ \geq&C_{Q}(r^{\prime\prime}_{k,S})^{3+\gamma}\xi_{k+1}^{\frac{1}{2}}\mathbf{1}_{B(0,r^{\prime\prime}_{k,S}\sqrt{2}(1-\xi_{k+1}))}=C_{Q}(r^{\prime\prime}_{k,S})^{3+\gamma}\xi_{k+1}^{\frac{1}{2}}\mathbf{1}_{B(0,r^{\prime\prime}_{k+1,S})}.\end{split}$
As a result, we deduce from (5.53), (5.55), and (5.56) the following estimate: for any $(t,x,v)$ satisfying (5.52), we have
(5.57) $\begin{split}&f(t,x,v)\\ \geq&\int_{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+1}r^{\prime\prime}_{k+1,S}})}^{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+2}r^{\prime\prime}_{k+1,S}})}(a^{S}_{k,\tau^{\prime\prime}}(\tau^{S}))^{2}\\ &\times e^{-C_{L}(t-s)\langle r^{\prime\prime}_{k+1,S}\rangle^{\gamma^{+}}}C_{Q}(r^{\prime\prime}_{k,S})^{3+\gamma}\xi_{k+1}^{\frac{1}{2}}\mathbf{1}_{B(0,r^{\prime\prime}_{k+1,S})}(V_{s,t}(x,v))\,ds\\ \geq&\int_{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+1}r^{\prime\prime}_{k+1,S}})}^{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+2}r^{\prime\prime}_{k+1,S}})}(a^{S}_{k,\tau^{\prime\prime}}(\tau^{S}))^{2}\\ &\times e^{-C_{L}(t-s)\langle r^{\prime\prime}_{k+1,S}\rangle^{\gamma^{+}}}C_{Q}(r^{\prime\prime}_{k,S})^{3+\gamma}\xi_{k+1}^{\frac{1}{2}}\,ds\\ \geq&(a^{S}_{k,\tau^{\prime\prime}}(\tau^{S}))^{2}C_{Q}(r^{\prime\prime}_{k,S})^{3+\gamma}\xi_{k+1}^{\frac{1}{2}}\frac{\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})}{2^{k+2}r^{\prime\prime}_{k+1,S}}\\ &\times e^{-C_{L}\frac{(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left\langle r^{\prime\prime}_{k+1,S}\right\rangle^{\gamma^{+}}}{2^{k+1}r^{\prime\prime}_{k+1,S}}}\\ =&a^{S}_{k+1,\tau^{\prime\prime}}(\tau^{S}).\end{split}$
Hence, we showed that (5.50) holds for $n=k+1$ and conclude the proof by induction.
Thanks to Proposition 5.4, we can use the exact same method used at the end of the fourth chapter to derive the Maxwellian lower bound for the fully specular reflection boundary condition.
6. Lower bound for the non-cutoff case
In this chapter, we investigate lower bounds for the solution of the Boltzmann equation in the non-cutoff case, that is, $\nu\geq 0$. Here, the domain $\Omega$ is not necessarily convex. In contrast to the cutoff case studied in Chapter 2–4, the non-cutoff setting requires a refined decomposition of the collision operator to handle the angular singularity:
(6.1) $\begin{split}Q[h_{1},h_{2}](v)&=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})-h_{2}(v)h_{1}(v_{*})]\,dv_{*}\,d\sigma\\ &=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{1}(v^{\prime}_{*})(h_{2}(v^{\prime})-h_{2}(v))]\,dv_{*}\,d\sigma\\ &\hskip 42.67912pt-h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})(h_{1}(v_{*})-h_{1}(v^{\prime}_{*}))\,dv_{*}\,d\sigma\\ &=:Q^{1}_{b}[h_{1},h_{2}](v)-Q^{2}_{b}[h_{1},h_{2}](v),\end{split}$
where
$Q^{1}_{b}[h_{1},h_{2}](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{1}(v^{\prime}_{*})(h_{2}(v^{\prime})-h_{2}(v))]\,dv_{*}\,d\sigma,$
$Q^{2}_{b}[h_{1},h_{2}](v):=h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{1}(v_{*})-h_{1}(v^{\prime}_{*})]\,dv_{*}\,d\sigma.$
With this notation and (for some $\epsilon>0$) the decomposition of $B(|v-v_{*}|,\cos{\theta})$
(6.2) $B(|v-v_{*}|,\cos{\theta})=\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta\geq\epsilon}+\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta<\epsilon},$
we can consider the singular and non-singular parts of $Q$:
$Q[h_{1},h_{2}](v)=Q^{+}_{\epsilon}[h_{1},h_{2}](v)-Q^{-}_{\epsilon}[h_{1},h_{2}](v)+Q^{1}_{\epsilon}[h_{1},h_{2}](v)+Q^{2}_{\epsilon}[h_{1},h_{2}](v),$
where
$Q^{+}_{\epsilon}[h_{1},h_{2}](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta\geq\epsilon}[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})]\,dv_{*}\,d\sigma,$
$Q^{-}_{\epsilon}[h_{1},h_{2}](v):=h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta\geq\epsilon}h_{1}(v_{*})\,dv_{*}\,d\sigma,$
$Q^{1}_{\epsilon}[h_{1},h_{2}](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta<\epsilon}\left[h_{1}(v^{\prime}_{*})(h_{2}(v^{\prime})-h_{2}(v))\right]\,dv_{*}\,d\sigma,$
$Q^{2}_{\epsilon}[h_{1},h_{2}](v):=h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta<\epsilon}[h_{1}(v_{*})-h_{1}(v^{\prime}_{*})]\,dv_{*}\,d\sigma.$
We further introduce the following notations
(6.3) $b^{CO}_{\epsilon}(\cos{\theta}):=\mathbf{1}_{\theta\geq\epsilon}b(\cos{\theta}),\,b^{NCO}_{\epsilon}(\cos{\theta}):=\mathbf{1}_{\theta<\epsilon}b(\cos{\theta}),$
(6.4) $m_{b}:=\int_{\mathbb{S}^{2}}b(\cos{\theta})(1-\cos{\theta})\,d\sigma,$
(6.5) $n_{b^{CO}_{\epsilon}}:=\int_{\mathbb{S}^{2}}b^{CO}_{\epsilon}(\cos{\theta})\,d\sigma,\,m_{b^{NCO}_{\epsilon}}:=\int_{\mathbb{S}^{2}}b^{NCO}_{\epsilon}(\cos{\theta})(1-\cos{\theta})\,d\sigma,$
$L_{\epsilon}(h)(v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta\geq\epsilon}h(v_{*})\,dv_{*}\,d\sigma,$
$S[h](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})[h(v_{*})-h(v^{\prime}_{*})]\,dv_{*}\,d\sigma,$
$S_{\epsilon}[h](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta<\epsilon}[h(v_{*})-h(v^{\prime}_{*})]\,dv_{*}\,d\sigma.$
With this decomposition of the collision operator, we now introduce the definition of mild solutions in the non-cutoff setting:
Definition 6.1.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and the kernel $B$ satisfies Assumption B with $\nu\geq 0$. Given a non-negative continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$, we call a non-negative continuous function $f$ defined on $0,T)\times(\overline{\Omega}\times\mathbb{R}^{3})$ with $|f(t,x,v)|\leq C(1+|v|)^{-r}$ for some constant $C>0$ and $r>3$ for any $0<t\leq T$, $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$ a ”continuous mild” solution to (1.1)–(1.3) with initial data $f_{0}$ if $f$ is continuous on $\Gamma_{conti}$ and there exists a number $0<\epsilon_{0}<\frac{\pi}{4}$ such that for any $0<\epsilon<\epsilon_{0}$, $(t,x,v)\in[0,T)\times\Omega\times\mathbb{R}^{3}$, we have
(6.6) $\begin{split}f(t,x,v)&=f_{0}(X_{0,t}(x,v),v)\exp\bigg[-\int_{0}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\bigg]\\ &+\int_{0}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}+Q_{\epsilon}^{1})[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\end{split}$
when $t\leq t_{\partial}(x,v)$, and
(6.7) $\begin{split}f(t,x,v)&=\alpha f(t-t_{\partial}(x,v),X_{t-t_{\partial}(x,v),t}(x,v),R(X_{t-t_{\partial}(x,v),t}(x,v),v))\\ &\exp\left[-\int_{t-t_{\partial}(x,v)}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right]\\ &+(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v))>0}f(t,X_{t-t_{\partial}(x,v),t}(x,v),w)(w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v)))dw\right)\\ &\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}\exp\left[-\int_{t-t_{\partial}(x,v)}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right]\\ &+\int_{t-t_{\partial}(x,v)}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}+Q_{\epsilon}^{2})[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds\end{split}$
when $t\geq t_{\partial}(x,v)$. Here, we recall the definition of $t_{\partial}(x,v)$ from Definition 1.1.
Before introducing the main result in the non-cutoff case, we introduce the following constants (here $\widetilde{\gamma}:=(2+\gamma)^{+}$):
(6.8) $e^{\prime}_{f}(t,x):=\int_{v\in\mathbb{R}^{3}}|v|^{\widetilde{\gamma}}f(t,x,v)\,dv,$
(6.9) $E^{\prime}_{f}:=\sup\limits_{[0,T)\times\Omega}e^{\prime}_{f}(t,x),$
(6.10) $w_{f}(t,x):=\|f(t,x,\cdot)\|_{W^{2,\infty}_{v}},$
(6.11) $W_{f}:=\sup\limits_{[0,T)\times\Omega}w_{f}(t,x).$
Now, we introduce our main theorem for the non-cutoff case:
Theorem 6.1.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the kernel $B$ satisfies Assumption B with $\nu\geq 0$, $\alpha\in[0,1]$. We consider a non-negative function $f_{0}$ that is continuous on $\overline{\Omega}\times\mathbb{R}^{3}$. Let $f(t,x,v)$ be a continuous mild solution to (1.1)–(1.3) on $0,T)\times\overline{\Omega}\times\mathbb{R}^{3}$, with $T>0$ (and the initial condition $f_{0}$) which satisfies the following properties:
  1. (1)
    $M>0$;
  2. (2)
    $E_{f}<\infty$, where $p_{\gamma}>\frac{3}{3+\gamma}>0$, if $3<\gamma<0$;
  3. (3)
    $W_{f}<\infty$ and $E^{\prime}_{f}<\infty$.
Then, the following lower bound holds: There exists $0<\tau_{0}\leq T$ such that for any $\tau\in(0,\tau_{0})$ and $K>2\frac{\log{(2+\frac{2\nu}{2-\nu})}}{\log{2}}$, there exist $\Delta_{\tau_{0}}>0$, $\rho>0$ and $\theta>0$ depending on $\tau_{0}$, $C_{\Phi},c_{\Phi},\gamma,b_{0},\nu,E_{f}$ $W_{f}$$E^{\prime}_{f}$, $M$ (and $L_{f,p}$ if $\gamma<0$ ), $\tau$, $K$, and the modulus of continuity of $f_{0}$, such that
(6.12) $f(t,x,v)\geq\frac{\rho}{(2\pi\theta)^{\frac{3}{2}}}e^{-\frac{|v|^{K}}{2\theta}},\ \forall\,t\in[\tau,\Delta_{\tau_{0}}),\ \forall\,(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}.$
In the case when $\nu=0$, we can further take $K=2$.
To prove this theorem, we quote the following result extracted from Corollary 2.2 in [11]:
Lemma 6.1.
Consider $g$ a measurable function on $\mathbb{R}^{3}$, and assume that the collision operator satisfies Assumption B with $0\leq\nu<2$. Then, there exists $C_{g}^{S}>0$ which depends only on $m_{b}$, $C_{\Phi}$ and $e_{g}$ (and $l_{g,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$) such that
(6.13) $|S[g](v)|\leq C_{g}^{S}\left\langle v\right\rangle^{\gamma^{+}}.$
Here we recall the definition of $e_{g}$ in (1.16), $l_{g,p}$ in (1.17) and define $m_{b}:=\int_{\mathbb{S}^{2}}b(\cos{\theta})(1-\cos{\theta})\,d\sigma$.
Similarly to the cut-off case, we would like to use (6.6) and (6.7) to spread the initial lower bounds. However, the additional non-cutoff term $Q^{1}_{\epsilon}$ is not necessarily nonnegative. Hence, we need to quote Lemma 2.5 from [11]:
Lemma 6.2.
Suppose that the collision operator satisfies Assumption B with $0\leq\nu<2$. Then, there exists a constant $C>0$, which depends only on $\gamma,\ \nu,\ b_{0}$, such that for any measurable function $h_{1},h_{2}$, we have
(6.14) $|Q^{1}_{b}\left(h_{1},h_{2}\right)|\leq Cm_{b}c_{\Phi}\|h_{2}\|_{L^{1}_{\widetilde{\gamma}}}\|h_{1}\|_{W^{2,\infty}}.$
Now, we show that a result which is similar to the result of Proposition 3.2 can be obtained:
Proposition 6.2.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and the kernel $B$ satisfies Assumption B with $0\leq\nu<2$, $\alpha\in[0,1]$. We consider a mild continuous solution $f(t,x,v)$ of (1.1)–(1.3). Given $A>0$, $0<\Delta_{1}\leq 1,$ $\Delta_{2}<1$, and $(\tau,x^{\prime},v^{\prime})\in[0,T)\times\Omega\times\mathbb{R}^{3}$ such that $B(x^{\prime},\Delta_{2})\subset\Omega$ and
$f(t,x,v)\geq A,\ \forall(t,x,v)\in[\tau,\tau+\Delta_{1}]\times B(x^{\prime},\Delta_{2})\times B(v^{\prime},\Delta_{2}).$
Then, there exist $\{\epsilon_{i}>0\}_{i=0}^{\infty}$ such that for $n\in\mathbb{N}\cup\{0\}$, $t\in[\tau,\tau+\Delta_{1}]$, $x\in B(x^{\prime},\frac{\Delta_{2}}{2^{n}})$,
(6.15) $\forall v\in\mathbb{R}^{3},\ f(t,x,v)\geq\alpha^{NC}_{n}(\tau,t,\Delta_{2},A,|v^{\prime}|)\mathbf{1}_{B(v^{\prime},r_{n}(\Delta_{2}))}(v),$
where the numbers $\{r_{n}(\Delta_{2})\}_{n=0}^{\infty}\in\mathbb{R}$, $\{t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)\}_{n=0}^{\infty}\in\mathbb{R}$ and $\{\alpha^{NC}_{n}(\tau,t,\Delta_{2},A,|v^{\prime}|)\}_{n=0}^{\infty}\in\mathbb{R}$ are defined as follows:
(6.16) $r_{0}(\Delta_{2}):=\Delta_{2},\ r_{n+1}(\Delta_{2}):=\frac{3\sqrt{2}}{4}r_{n}(\Delta_{2}),$
(6.17) $t^{NC}_{n}(\tau,t,\Delta_{2},|v^{\prime}|):=\max\left\{\tau,t-\frac{\Delta_{2}}{2^{n+1}(2r_{n}(\Delta_{2})+|v^{\prime}|)}\right\},$
(6.18) $\alpha_{0}:=A$
(6.19) $\begin{split}&\alpha^{NC}_{n+1}(\tau,t,\Delta_{2},A,|v^{\prime}|)\\ \ :=&\frac{1}{4}\int_{t^{NC}_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon_{n}}}+C_{L,2}m_{b^{NCO}_{\epsilon_{n}}})\left\langle|v^{\prime}|+2r_{n}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,1}\alpha^{NC}_{n}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon_{n}}}c_{\Phi}(r_{n}(\Delta_{2}))^{3+\gamma}\,ds.\end{split}$
where $C_{L,1}$, $C_{L,2}$ and $C_{Q,1}$ are constants which depend on $\gamma,\ \nu,\ b_{0},\ n_{b}$, $m_{b}$, $C_{\Phi}$, $E_{f}$, $W_{f}$, $E^{\prime}_{f}$, MM(and $L_{f,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$).
Proof.
We prove the proposition by induction on $n$. The case $n=0$ is exactly the assumption. Assume that the statement of Proposition 6.2 holds for $n=k$. Given $t\in[\tau,\tau+\Delta_{1}]$, $x\in B\left(x^{\prime},\frac{\Delta_{2}}{2^{k+1}}\right)$, $v\in B\left(v^{\prime},r_{k+1}(\Delta_{2})\right)\subset B\left(0,|v^{\prime}|+2r_{k}(\Delta_{2})\right)$, we consider the second term on the right-hand side of (6.6) to obtain the following lower bound which holds for $0<\epsilon<\epsilon_{0}$ ($\epsilon_{0}$ is mentioned in Definition 6.1):
(6.20) $\begin{split}&f(t,x,v)\\ =&f(t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|),X_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|),t}(x,v),v)\\ &\exp\bigg[-\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\bigg]\\ &+\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}+Q_{\epsilon}^{1})[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\\ \geq&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v))\,ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}-|Q_{\epsilon}^{1}|)[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds.\end{split}$
Now, by Lemma 3.1 and Lemma 6.1, we have
(6.21) $|(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)|\leq(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}$
for some constants $C_{L,1},C_{L,2}$ depending only on $n_{b}$, $m_{b}$, $C_{\Phi}$ and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$).
Hence, the following estimate holds:
(6.22) $\begin{split}&f(t,x,v)\\ \geq&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &(Q^{+}_{\epsilon}-|Q_{\epsilon}^{1}|)[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\end{split}$
for all $\frac{\pi}{4}>\epsilon>0$.
We notice that when $s\in\left[t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|),t\right]$, we have
$|x^{\prime}-X_{s,t}(x,v)|=|x^{\prime}-x+tv-sv|\leq\frac{\Delta_{2}}{2^{k+1}}+|t-s|(|v^{\prime}|+2r_{k}(\Delta_{2}))\leq\frac{\Delta_{2}}{2^{k}},$
which implies that $X_{s,t}(x,v)\in B\left(x^{\prime},\frac{\Delta_{2}}{2^{k}}\right)\subset\Omega$ and we deduce by induction hypothesis that
(6.23) $\forall v\in\mathbb{R}^{3},\ f(s,X_{s,t}(x,v),v)\geq\alpha^{NC}_{k}(\tau,t,\Delta_{2},A,|v^{\prime}|)\mathbf{1}_{B(v^{\prime},r_{k}(\Delta_{2}))}(v).$
Next, we obtain from Lemma 3.2 by setting $\xi=\frac{1}{4}$: for any $v\in B\left(v^{\prime},r_{k+1}(\Delta_{2})\right)$, we have
(6.24) $\begin{split}&Q_{\epsilon}^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\\ \geq&C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\frac{1}{2}\mathbf{1}_{B\left(v^{\prime},\frac{3\sqrt{2}}{4}r_{k}(\Delta_{2})\right)}\\ =&C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\frac{1}{2},\end{split}$
for some $C_{Q,1}>0$ which only depends on $\gamma,\ \nu,\ b_{0}$.
Now, we use (6.14) to derive:
(6.25) $\begin{split}|Q_{\epsilon}^{1}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)|\leq C_{Q,2}m_{b^{NCO}_{\epsilon}}c_{\Phi}E^{\prime}_{f}W_{f},\end{split}$
for some $C_{Q,2}>0$, which depends only on $\gamma,\ \nu,\ b_{0}$.
As a result, we deduce that
(6.26) $\begin{split}&f(t,x,v)\\ \geq&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636pt\left(C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\frac{1}{2}-C_{Q,2}m_{b^{NCO}_{\epsilon}}c_{\Phi}E^{\prime}_{f}W_{f}\right)\,ds\\ \geq&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\frac{1}{2}\,ds\\ -&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,2}m_{b^{NCO}_{\epsilon}}c_{\Phi}E^{\prime}_{f}W_{f}\,ds,\end{split}$
for all $\frac{\pi}{4}>\epsilon>0$.
Then, we notice that when $\frac{\pi}{4}>\epsilon>0$, we ahve
(6.27) $l_{b^{CO}_{\epsilon}}:=\inf\limits_{\frac{1}{4}\pi\leq\theta\leq\frac{3}{4}\pi}b^{CO}_{\epsilon}(\cos{\theta})\geq l_{b},$
Thus, we deduce from (6.4) and (6.5) that
(6.28) $\begin{split}n_{b^{CO}_{\epsilon}}=\int_{\mathbb{S}^{2}}b^{CO}_{\epsilon}(\cos{\theta})\,d\sigma\sim\begin{cases}\frac{b_{0}}{\nu}\epsilon^{-\nu},\,\nu\in(0,2),\\ b_{0}|\log{\epsilon}|,\,\nu=0,\end{cases}\end{split}$
(6.29) $\begin{split}m_{b^{NCO}_{\epsilon}}=\int_{\mathbb{S}^{2}}b^{NCO}_{\epsilon}(\cos{\theta})(1-\cos{\theta})\,d\sigma\sim\begin{cases}\frac{b_{0}}{2-\nu}\epsilon^{2-\nu},\,\nu\in(0,2),\\ \frac{b_{0}}{2}\epsilon^{2},\,\nu=0.\end{cases}\end{split}$
Hence, for any $k\in\mathbb{N}$, we can choose $\frac{\pi}{4}>\epsilon_{k}>0$ such that
(6.30) $\begin{split}&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon_{k}}}+C_{L,2}m_{b^{NCO}_{\epsilon_{k}}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,2}m_{b^{NCO}_{\epsilon_{k}}}c_{\Phi}E^{\prime}_{f}W_{f}\,ds\\ \leq&\frac{1}{4}\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon_{k}}}+C_{L,2}m_{b^{NCO}_{\epsilon_{k}}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon_{k}}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\,ds\end{split}$
and therefore
(6.31) $\begin{split}&f(t,x,v)\\ \geq&\frac{1}{4}\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon_{k}}}+C_{L,2}m_{b^{NCO}_{\epsilon_{k}}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon_{k}}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\,ds\\ =&\alpha^{NC}_{k+1}(\tau,t,\Delta_{2},A,|v^{\prime}|),\end{split}$
for $v\in B(v^{\prime},r_{k+1}(\Delta_{2}))$ and we conclude the proof by induction.
With the help of Proposition 6.2, we can use the exact same argument from Proposition 3.3 to Proposition 3.5 to generate series of lower bounds on $\mathbb{Y}$. Then, we can work as in Proposition 4.1 to deduce the following.
Proposition 6.3.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the kernel $B$ satisfies Assumption B with $\nu\geq 0$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha\in[0,1]$. Fix $\tau\in(0,\Delta^{0})$, with $\Delta^{0}>0$ is given in Proposition 3.1. Then, there exist $b^{NC}(\tau)$, $\delta_{T}(\tau)>0$, which depend on $\gamma,\ \nu,\ b_{0},\ n_{b}$, $m_{b}$ $C_{\Phi}$ and $E_{f}$, $W_{f}$, $E^{\prime}_{f}$, and MM(and $L_{f,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$) such that for $t\in[\tau,\tau+\delta_{T}(\tau)]$,
(6.32) $\forall x\in\partial\Omega,\ \int_{v_{*}\cdot n(x)>0}f(t,x,v_{*})(v_{*}\cdot n(x))\,dv_{*}>b^{NC}(\tau).$
Using the same method as in [1, 2] (notice that the proof does not require the convexity of $\Omega$), we derive the following.
Lemma 6.3.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and the kernel $B$ satisfies Assumption B with $\nu\geq 0$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha\in[0,1]$. Fix $\tau\in(0,\Delta^{0})$, where $\Delta^{0}>0$ is given in Proposition 3.1 and $\{\xi_{n}\}_{n=1}^{\infty}\in(0,1)$. Then, there exist $r^{NC}_{0}>0$, $\{\Delta_{i}\}_{i=1}^{\infty}$ such that $\sum_{i=1}^{\infty}\Delta_{i}=1$ and that for any $n\in\mathbb{N}$, $t\in[(\sum_{i=1}^{n}\Delta_{i})\tau,\tau]$, $\forall(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$, we have
(6.33) $f(t,x,v)\geq a_{n}^{NC}(\tau)\mathbf{1}_{B(0,r^{\prime\prime}_{n})}(v)$
Here,
(6.34) $r^{NC}_{n+1}:=\sqrt{2}\,(1-\xi_{n})\,r^{NC}_{n},$
(6.35) $a_{n+1}^{NC}(\tau):=\begin{cases}C_{Q}\Delta_{n+1}e^{-\left[C_{L}\left(a^{NC}_{n}(\tau)\right)^{2}\left(r_{n}^{NC}\right)^{3+\gamma-\widetilde{\gamma}}\xi_{n}^{\frac{1}{2}}\right]^{\frac{-\nu}{2-\nu}}\left(\sum\limits_{k\geq n+1}\Delta_{k}\right)\left(r_{n+1}^{NC}\right)^{\gamma^{+}}}\\ \times\left(a^{NC}_{n}(\tau)\right)^{2}\left(r_{n}^{NC}\right)^{\gamma+3}\xi_{n}^{\frac{5}{2}},&\nu\in(0,2),\\ C_{Q}\Delta_{n+1}e^{-C\log\left[C_{L}\left(a^{NC}_{n}(\tau)\right)^{2}\left(r_{n}^{NC}\right)^{3+\gamma-\widetilde{\gamma}}\xi_{n}^{\frac{1}{2}}\right]\left(\sum\limits_{k\geq n+1}\Delta_{k}\right)\left(r_{n+1}^{NC}\right)^{\gamma^{+}}}\\ \times\left(a^{NC}_{n}(\tau)\right)^{2}\left(r_{n}^{NC}\right)^{\gamma+3}\xi_{n}^{\frac{5}{2}},&\nu=0.\end{cases}$
Finally, by proceeding as in [11], p.29–31, we can derive the ”weaker than Maxwellian” lower bound of Theorem 6.1.
A. Appendix: the characteristic line
In this section, we recall some useful properties about the bounce trajectory of $\Omega$ used in the definition of the characteristic line $X_{s,t}(x,v)$ and $V_{s,t}(x,v)$. To begin with, we introduce some notations from Definition A.1 in [2].
Definition A.1.
Given an open bounded domain with a $C^{1}$ boundary, we define a partition of $\partial\Omega\times\mathbb{R}^{3}$ as follows:
  • $\Omega_{rebounds}:=\{(x,v)\in\partial\Omega\times\mathbb{R}^{3}\mid v\cdot n(x)<0\}.$
  • $\Omega_{rolling}:=\{(x,v)\in\partial\Omega\times\mathbb{R}^{3}\mid v\cdot n(x)=0,\ \exists\,\delta>0\ s.t.\ x-vt\in\overline{\Omega},\ \forall\,t\in[0,\delta]\}.$
  • $\Omega_{stop}:=\{(x,v)\in\partial\Omega\times\mathbb{R}^{3}\mid v\cdot n(x)=0,\ \forall\,\delta>0\ \exists\,t\in(0,\delta)\ s.t.\ x-vt\notin\overline{\Omega}\}.$
  • $\Omega_{line}:=\{(x,v)\in\partial\Omega\times\mathbb{R}^{3}\mid v\cdot n(x)>0\}.$
Due to Proposition A.2 in [2], there is no specular reflection trajectory that leads to any point in $\Omega_{stop}$. Given $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$, we recall the definition of $t_{\partial}(x,v)$
$t_{\partial}(x,v)=\max\{t\geq 0\mid x-vs\in\overline{\Omega},\forall\,s\in[0,t]\}.$
The next proposition(Proposition A.3 in [2]) gives a picture of a backward trajectory touching the boundary:
Proposition A.1.
Given an open bounded domain in $\mathbb{R}^{3}$ with $C^{1}$ boundary. Then, we have
(1) If there exists $t\in(0,t_{\partial}(x,v))$ such that $x-vt\in\partial\Omega$, then $(x-vt,v)\in\Omega_{rolling}$.
(2) $t_{\partial}(x,v)=0$ if and only if $(x,v)\in\Omega_{stop}\cup\Omega_{rebound}$.
(3) $(x-vt_{\partial},v)\in\Omega_{stop}\cup\Omega_{rebound}$.
We always need to consider some strange trajectory with a strange path. Thankfully, the following proposition (Proposition A.4 in [2]) shows that we do not need to worry about the case that the backward trajectory reaching $\Omega_{stop}$ or bounce infinite times in a finite time since those set are measure zero:
Proposition A.2.
We consider an open bounded domain in $\mathbb{R}^{3}$ with $C^{1}$ boundary and $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$. Then, for any $t\geq 0$ the trajectory of $(x,v)$ with specular reflection after time $t$ has at most a countable number of rebound and rolling. Moreover, the set with infinite rebound and rolling in finite time is zero measure in $\overline{\Omega}\times\mathbb{R}^{3}$.
Remark A.1.
Similarly, with the above assumption the subset of $\overline{\Omega}\times\mathbb{R}^{3}$ with a backward trajectory leading to $\Omega_{stop}$ is also measure zero in $\overline{\Omega}\times\mathbb{R}^{3}$.
Now, we introduce the characteristic line by defining:
(A.1) $(t_{0},x_{0},v_{0}):=(0,x,v),$
(A.2) $\begin{split}&(t_{k+1},x_{k+1},v_{k+1})\\ :=&\begin{cases}(\infty,x_{k},v_{k}),\ &(x_{k},v_{k})\in\Omega_{stop},\\ (t_{k}+t_{\partial}(x_{k},v_{k}),x_{k}-v_{k}t_{\partial}(x_{k},v_{k}),R(x_{k}-v_{k}t_{\partial}(x_{k},v_{k}),v_{k})),\ &(x_{k},v_{k})\notin\Omega_{stop},\end{cases}\end{split}$
(A.3) $n(t,x,v):=\sup\{k\in\mathbb{N}:\ t_{k}(x,v)\leq t\}.$
Note that $n(t,x,v)<\infty$ in case $\{t_{k}\}$ is unbounded.
Next, we define the last rebound from the backward trajectory of $(x,v)$:
(A.4) $(t_{fin},x_{fin},v_{fin}):=\begin{cases}(t,x_{n},v_{n}),&n(t,x,v)<\infty,\ \,t_{n(t,x,v)+1}=\infty,\\ (t_{n},x_{n},v_{n}),&n(t,x,v)<\infty,\ \,t_{n(t,x,v)+1}<\infty,\\ \lim\limits_{n\rightarrow\infty}(t_{n},x_{n},v_{n}),&n(t,x,v)=\infty.\end{cases}$
Finally, for $0\leq s\leq t$, we define the characteristic line :
$\begin{cases}X_{t}(x,v)&:=x_{fin}(t,x,-v)-(t-t_{fin}(t,x,-v))v_{fin}(t,x,-v),\\ V_{t}(x,v)&:=-v_{fin}(t,x,-v).\end{cases}$
$\begin{cases}X_{s,t}(x,v)&:=X_{s}(X_{t}(x,-v),-V_{t}(x,-v)),\\ V_{s,t}(x,v)&:=V_{s}(X_{t}(x,-v),-V_{t}(x,-v)).\end{cases}$
Acknowledgement
The author would like to expresses sincere gratitude to Professor Laurent Desvillettes for his valuable suggestions during the preparation of this manuscript, in the framework of a research sojourn in Université Paris Cité, IMJ-PRG.
This work is supported by the NSTC Graduate Student Study Abroad Program with grant number 113-2917-I-002-054.
Data availability statement
Data sharing is not applicable to this article, as no data sets were generated or analyzed during the current study.